### v1

# High-dimensional dynamics in low-dimensional networks

Abstract
Many networks that arise in nature and applications are effectively low-dimensional in the sense that their connectivity structure is dominated by a few dimensions. It is natural to expect that dynamics on such networks might also be low-dimensional. Indeed, recent results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or noise. However, networks in nature are rarely isolated. We show that recurrent networks with low-rank structure often produce high-dimensional dynamics in the presence of high-dimensional perturbations. Counter to intuition, dynamics in these networks are suppressed in directions that are aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” Our results clarify important, but counterintuitive relationships between a network’s connectivity structure and the structure of the dynamics it generates.
Introduction
Recent work shows that many networks arising in nature and in applications are effectively low-rank in the sense that their connectivity matrices have a small number of larger singular values, while the remaining singular values are much smaller [1]. What are the implications of low-rank network structure on the dimensionality of network dynamics?
Several studies have shown that networks with low-rank structure tend to produce low-dimensional dynamics [1, 2, 3, 4]. However, these studies rely on specific assumptions about the network model. Some studies assume that the low-rank part of the network is weak [2, 4], others assume that the network is self-contained in the sense that it operates in the absence of internal noise or external perturbations or inputs [5, 1], while other work assumes that external perturbations are perfectly aligned with the network’s low-rank structure [3]. In reality, networks in nature are rarely isolated from perturbations, and these perturbations are not necessarily aligned with the network’s low-rank structure.
We use mathematical analysis and computer simulations to show that, in the presence of high-dimensional inputs or perturbations, many networks with low-dimensional structure produce high-dimensional dynamics. Moreover, perhaps counterintuitively, these networks suppress variability in directions that are aligned with their low-rank structure relative to random directions, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity structure promote low-rank suppression and high-dimensional dynamics. Finally, we demonstrate our conclusions in a real world epidemiological network with low-rank structure.
Our conclusions have important implications for the interpretation of low-dimensional structure in recurrent networks. In neuroscience, our results offer a mathematical explanation of the widely observed balance between excitatory and inhibitory synaptic currents [6, 7, 8, 9, 10, 11, 12] and also explain why neural populations generate high-dimensional responses to high-dimensional stimuli [13]. Beyond neuroscience, our results imply that randomly structured perturbations are often most effective at driving activity in networks with low-rank structure.
Results
High-dimensional dynamics and low-rank suppression in a network with rank-one structure
For illustrative purposes, we begin with a simple, linear model (Figure 1a)
$\tau\frac{d{\bf{z}}}{dt}$ $=-{\bf{z}}+W{\bf{z}}+{\bf{x}}.$ (1)
where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and we take $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
$W=W_{0}+W_{1}$
where
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
is a rank-one matrix and
$W_{1}=\frac{\rho}{\sqrt{N}}Z$
is a full rank random matrix. Specifically, ${\bf{u}}$ is a random vector with $\|{\bf{u}}\|=1$, $Z$ is an $N\times N$ matrix with entries drawn i.i.d. from a standard normal distribution, and $\rho>0$. If
$|c|\gg\rho$
then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are bounded by $2\rho$ (Figure 1b). Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane [14, 15]. Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [16]. Since low-rank structure requires $|c|\gg\rho$, stability requires $c<0$ when $\rho\sim\mathcal{O}(1)$. Here, we take $\rho=0.5$ and $c=-10$. We consider more general network structures later.
One might expect a network with connectivity $W$ to produce low-dimensional dynamics due to its low-rank structure. Indeed, a feedforward network with the same connectivity matrix produces approximately one-dimensional dynamics because it amplifies inputs aligned to ${\bf{u}}$ relative to other directions (Supplementary Figure S.1). We next demonstrate that this seemingly natural conclusion does not carry over to recurrent network dynamics.
Figure 1: Response properties of a recurrent network with rank-one structure. a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, ${\bf{z}}$, which receive external perturbations or input, ${\bf{x}}$. b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure. c) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$). d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue). e,f) The network response (e) and input (f) projected onto the plane determined by ${\bf{u}}$ and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along ${\bf{u}}$. g) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of ${\bf{u}}$.
We first simulated the network with two different input patterns. The “aligned” input pattern was exactly aligned with the low-rank component of $W$,
${\bf{x}}_{\textrm{aligned}}={\bf{u}},$
while the “random” input pattern is aligned with a random direction,
${\bf{x}}_{\textrm{random}}={\bf{u}}_{\textrm{rand}},$
which is generated identically to, but independently from ${\bf{u}}$. For large $N$, ${\bf{u}}$ and ${\bf{u}}_{\textrm{rand}}$ are approximately orthogonal. Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network. Perhaps surprisingly, we observed exactly the opposite (Figure 1c): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
To demonstrate some consequences of this effect, we next provided a high-dimensional input, ${\bf{x}}(t)$, that varied smoothly in time. Specifically, each ${\bf{x}}_{j}(t)$ was an i.i.d. realization of a smooth, unbiased, stationary Gaussian process. This can model internal noise or irregular external input to the network.
Conventional wisdom and intuition might lead us to expect the network response to be low-dimensional and dominated by variability in the direction of ${\bf{u}}$, owing to the low-rank structure of $W$. In contrast, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$. The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d). Perhaps surprisingly, this weakest principal component direction was closely aligned to ${\bf{u}}$ (the angle between ${\bf{u}}$ and the corresponding principal component direction was less than $8^{\circ}$). Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of ${\bf{u}}$ was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
We propose the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to dominant the low-rank structure of a network are suppressed by the network’s dynamics. In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional perturbations, as in Figure 1d. In this manuscript, we derive conditions under which low-rank suppression and high-dimensional dynamics arise in low-rank networks. We then show that these conditions are met by many natural low-rank network structures.
The results in Figure 1 contrast with previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1]. In some of these studies, external input or perturbations to the network are assumed to be absent [1, 5] or perfectly aligned with the low-rank component of the network connectivity [3]. In other studies [2, 4], the network is assumed to have only a weak low-rank component ($|c|\sim\rho$). None of these assumptions are consistent with the model considered in Figure 1. See Supplementary Text S.1 for a more detailed review of this previous work and its relation to ours.
A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady state solution to Eq. (1),
${\bf{z}}=[I-W]^{-1}{\bf{x}}.$ (2)
Convergence to the steady state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq. (2). Because of the matrix inverse in Eq. (2), the large singular value of $W$ near $|c|$ produces a small singular value of $[I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near ${\bf{u}}$. Hence, the recurrent network suppresses inputs in the direction of ${\bf{u}}$. A related explanation is that the Jacobian matrix $J=W-I$ of Eq. (1) has a large, negative eigenvalue near $c-1<0$ with associated eigenvector near ${\bf{u}}$ so the dynamics in Eq. (1) are highly compressive along ${\bf{u}}$. These explanations depend on specific properties of the model considered here, but we later describe more general conditions under which low-rank suppression and high-dimensional dynamics occur.
An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with network interactions in the direction of ${\bf{u}}$. More precisely, note that $W{\bf{z}}(t)$ in Eq. (1) can be interpreted as a vector of local inputs to each node whereas ${\bf{x}}(t)$ is external input, and ${\bf{z}}(t)$ is the network response. In the qausi-steady state, ${\bf{z}}\approx W{\bf{z}}+{\bf{x}}$. Under low-rank suppression, local input cancels nearly perfectly with external input ($W{\bf{z}}\approx-{\bf{x}}$) in the direction of ${\bf{u}}$ so that the response, ${\bf{z}}$, is weak in this direction (Figure 1g). This cancellation is a key feature of low-rank suppression.
Conditions for low-rank suppression and high-dimensional dynamics
Above, we considered a simple example of a network with rank-one structure. In Supplementary Text S.2, we derive conditions for low-rank suppression and high-dimensional dynamics in a general class of models that includes all dynamics of the form $\tau d{\bf{z}}/dt={\bf{F}}({\bf{z}},{\bf{x}})$. Here, we summarize and interpret those results.
Low-rank suppression (as in Figure 1c) is a more general phenomenon than high-dimensional dynamics (as in Figure 1d). In Supplementary Text S.2, we show that low-rank suppression arises under two assumptions:
  1. 1.
    Assumptions on singular values. First, we assume that the network is strongly low-rank in the sense that the connectivity matrix $W$ has a small number of large singular values and the rest are much smaller (Figure 2a). Specifically, $W=W_{0}+W_{1}$ where $W_{0}$ has rank $r\ll N$ with large singular values ($\sigma_{W_{0}}\gg 1$) while $W_{1}$ has rank $N$ with small or moderate singular values ($\sigma_{W_{1}}\leq\mathcal{O}(1)$). “Weakly low-rank” networks for which $\sigma_{k}\sim\mathcal{O}(1)$ are studied in other work [2, 17, 4, 18, 19, 20] and do not necessarily produce low-rank suppression (see Supplementary Text S.1).
  2. 2.
    Assumptions on dynamics. In addition, our analysis relies on a linearization of the dynamics around a stable equilibrium. This requires that ${\bf{z}}(t)$ settles to a stable equilibrium whenever ${\bf{x}}(t)=0$, and that ${\bf{x}}(t)$ is sufficiently weak or the dynamics of ${\bf{z}}(t)$ are sufficiently linear. Nonlinear dynamics away from a stable equilibrium can produce any dynamics in principle [21, 22, 23] (see Supplementary Figures S.7 and S.8 for examples of unstable network states).
Under these two assumptions, low-rank suppression emerges in the sense that there are inputs, ${\bf{x}}$, approximately parallel to the column space of $U$, that produce much weaker responses than inputs at random directions (as in Figure 1c). The emergence of high-dimensional dynamics (as in Figure 1d) requires additional assumptions:
  1. 3.
    Assumptions on input perturbations. We assume that ${\bf{x}}(t)$ is a smooth, high-dimensional stochastic input. In simulations, we take each ${\bf{x}}_{j}(t)$ to be an i.i.d. Gaussian process. Low-dimensional inputs, ${\bf{x}}(t)$, can produce low-dimensional dynamics in networks that would otherwise produce high-dimensional dynamics (Supplementary Figure S.9).
  2. 4.
    Conditions on the alignment of singular vectors. Under the assumptions above, the emergence of high-dimensional dynamics depends on the structure of the low-rank part, $W_{0}$, of the connectivity. Specifically, let
    $W_{0}=U\Sigma V^{T}$
    be the singular value decomposition of $W$ with $U$ and $V$ being $N\times r$ orthonormal matrices. We define the $r\times r$ recurrent alignment matrix [3, 24],
    $P=U^{T}V.$
    which measures the alignment between the left and right singular vectors of $W_{0}$, i.e., the columns of $U$ and $V$. High-dimensional dynamics can only emerge when no singular value of $P$ is small, i.e., when $\sigma_{P}=\mathcal{O}(1)$ for all singular values of $P$.
To better understand this condition on $P$, we consider simulations of a model that is widely used in computational neuroscience to model a recurrently interacting network of neurons [16, 25, 2, 26],
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (3)
We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 2a) and we fix
$U=\left[\begin{array}[]{cc}{\bf{u}}_{1}&{\bf{u}}_{2}\end{array}\right]$
while exploring different choices of $V$.
We first consider the special case that $W_{0}$ is normal, so ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$. In this case, $P$ is diagonal with $|P_{kk}|=1$ (Figure 2b) so all singular values of $P$ are $\sigma_{P}=1$ and the network exhibits high-dimensional dynamics (Figure 2c). Since $W_{0}$ is normal, its eigenvalues are real and equal in magnitude to the singular values, $|\lambda_{W_{0}}|=\sigma_{W_{0}}$. Since $\sigma_{W_{0}}\gg 1$, stability of the dynamics requires that $\lambda_{W_{0}}=-\sigma_{W_{0}}$ and therefore
$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&-{\bf{u}}_{2}\end{array}\right].$
Dynamics on the column space of $U$ are strongly suppressed due to the large, negative eigenvalues.
Figure 2: Conditions for high-dimensional dynamics in a network with rank-two structure. a) Singular values of $W$ demonstrate an effective low-rank structure. b) The alignment matrix, $P$, when $W_{0}$ is normal. c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics. d,e) Same as b,c except $W_{0}$ is EP, but not normal. f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP. h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq. (4) with $c=0.5$).
However, low-rank suppression does not require that $W_{0}$ is normal (${\bf{v}}_{k}=\pm{\bf{u}}_{k}$). Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$. Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [27]. All normal matrices are EP, but an EP matrix is not necessarily normal. An example of a non-normal, EP matrix with rank 2 is given by
$V$ $=\left[\begin{array}[]{cc}{\bf{u}}_{2}&-{\bf{u}}_{1}\end{array}\right].$
The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 2d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$. Regardless, all singular values of $P$ are $\sigma_{P}=1$, and therefore the network produces low-rank suppression and high-dimensional dynamics (Figure 2e).
To obtain an example of a network that produces low-dimensional dynamics, we take
$V$ $=\left[\begin{array}[]{cc}{\bf{v}}_{\perp}&-{\bf{u}}_{2}\end{array}\right].$
where ${\bf{v}}_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$. In this case, $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 2f). Because of the singular value at zero, the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 2g).
Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap, for example
$V$ $=\left[\begin{array}[]{cc}\sqrt{1-c}{\bf{v}}_{\perp}-\sqrt{c}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (4)
where $10<c<$ 1. In this case, $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and another at $\sigma_{P}=1$ (Figure 2h). As long as $c$ is not asymptotically close to zero, the network produces low-rank suppression and high-dimensional dynamics (Figure 2i).
In summary, the class of effectively low-rank matrices, $W$, that produce low-dimensional dynamics in the presence of high-dimensional inputs is the very specific class of matrices for which the alignment matrix, $P=U^{T}V$, of the low-rank part is singular or nearly singular in the sense that it has a small or zero singular value.
It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean. In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.2). However, many networks arising in nature and applications do not have purely random structure. In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
Biased weights and modular networks
So far, we considered example networks unbiased weights, $E[W_{jk}]=0$, which is common in many modeling studies, but many networks in nature have weights with non-zero mean. Biased weights can produce low-rank structure. As a simple example, consider a random network with independent weights satisfying
$E[W_{jk}]=m\;\textrm{ and }\;\textrm{std}(W_{jk})=s.$
If $m$ and $s$ scale similarly to each other, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}\sim\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure because $\sigma_{1}/\sigma_{2}\sim\mathcal{O}(\sqrt{N})$. The dominant rank-one part of such a matrix has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.3).
Figure 3: Low-rank suppression and excitatory-inhibitory balance in a modular network. a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure. c) The alignment matrix shows that the network is EP, but not normal. d) An input that is constant within each block (red) is aligned to the low-rank part, but a random input (gray) is not. e) Response magnitude is suppressed for the aligned input relative to the random input. f) The excitatory (positive; red) component of an input balances with the inhibitory (negative; blue) component to produce a much smaller total (gray) component, a widely observed phenomenon in neural circuits.
More generally, modular structure arises in many natural settings [1, 28]. Specifically, many networks in nature represent the interaction between $n$ populations, and mean connection weights between populations are often non-zero. The adjacency matrices of modular networks can be arranged to have a block structure
$W=\left[\begin{array}[]{ccc}W^{1,1}&\cdots&W^{1,n}\\ W^{2,1}&\cdots&W^{2,n}\\ \vdots&\cdots&\vdots\\ W^{n,1}&\cdots&W^{n,n}\end{array}\right]$
where $W^{a,b}$ is a sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
$E[W_{jk}]=m_{ab}\;\textrm{ and }\;\textrm{std}(W_{jk})=s_{ab}.$
If we assume that each population has $\mathcal{O}(N)$ members and that $m_{ab}$ scale similarly to $s_{ab}$ for large $N$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(\sigma_{ab}\sqrt{N})$. Hence, large modular networks with biased weights naturally produce low-rank structure [1].
Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is the element-wise expectation of $W_{0}$, which is constant within each block, so $W_{0}$ has rank at most $n$. In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the $n$ indicator vectors of the $n$ populations. Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics in response to high-dimensional perturbations.
As a specific example of a modular network, we consider a model of a local neuronal network in the cerebral cortex. Cortical neurons obey Dale’s Law: All outgoing connection weights from a particular neuron have the same sign, positive for excitatory neurons and negative for inhibitory neurons, and mean connection weights also depend on the postsynaptic neuron type [29, 30]. These properties produce a modular structure in which the columns of the adjacency matrix corresponding to excitatory neurons are non-negative, while the columns corresponding to inhibitory neurons are non-positive (Figure 3a). We consider the case in which $m_{ab}$ and $s_{ab}$ scale like $1/\sqrt{N}$, consistent with experiments [31] and theoretical work [32, 33, 34, 35, 3, 36].
This network has an effective rank-two structure (Figure 3b): Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$ and the low-rank part of the connectivity matrix is EP (Figure 3c), so our theory predicts that the network exhibits low-rank suppression and high-dimensional responses to high-dimensional inputs. The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
$\textrm{col}(U)=\textrm{col}(V)=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}.$
Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 3d,e) consistent with recordings from monkey motor cortex [36] and previous theoretical work on balanced network models [37, 38, 39, 36].
As predicted, simulations also show high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.4), consistent with observations that neural responses in primate visual cortex are high-dimensional when visual stimuli are high-dimensional [13].
Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)$. As a result, the cancellation mechanism illustrated in Figure 1g manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 3f), a phenomenon that is widely observed in neural recordings [6, 7, 8, 9, 10, 11, 12] and widely studied in computational models [32, 33, 39, 3]. Therefore, the widely studied phenomenon of excitatory-inhibitory balance emerges as a consequence of low-rank suppression.
Amplification of disordered perturbations in networks with spatial structure
Figure 4: Low-rank suppression in a network with spatial structure. a) Connectivity matrix, $W$. Connection strength is a Gaussian function of distance. b) Singular values of $W$ demonstrate effective low-rank structure. c) A spatially smooth perturbation (red) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (gray) is not. d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
Many networks in nature exhibit connection strength that depends on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 34]. As a simple example, we considered a model in which each node is assigned a location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are more strongly connected (Figure 4a). Connectivity was also perturbed by a random component, $W_{1}$, as above. This connectivity structure is effectively low-rank with the first several singular values proportional to the Fourier coefficients of the Gaussian connectivity kernel while the remaining singular values, inherited from $W_{1}$, are much smaller (Figure 4b). The low-rank part is normal, with left- and right singular vectors equal to the Fourier basis vectors (Supplementary Figure S.5), so the network satisfies our conditions for low-rank suppression and high-dimensional responses.
Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low frequency spatial Fourier modes. Therefore, due to low-rank suppression, network responses to spatially disordered perturbations are amplified relative to smooth perturbations (Figure 4c-e; compare tick labels on vertical axes). The network also exhibits high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.6), as predicted.
Low-rank suppression and high-dimensional dynamics in an epidemiological network.
Figure 5: Low-rank suppression and high-dimensional dynamics in a real epidemiological network. a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction. b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
We next consider a real epidemiological network, specifically a network of high school social contacts [40], which was used in a recent theoretical study of low-rank networks [1]. In that study, the authors considered quenched mean-field reduction of the susceptible-infected-susceptible epidemiological model,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (5)
where $g>0$, $\gamma>0$ and $\circ$ denotes element-wise multiplication. In this model, each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected. Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [40]. This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a). In [1], it was shown that the dynamics, ${\bf{z}}(t)$, generated by Eq. (5) on this network are effectively low-dimensional. However, Eq. (5) is completely self contained without any internal noise or external perturbations. Noise and perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the high school) and through the natural stochasticity of infection.
Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network will produce low-rank suppression and high-dimensional responses to high-dimensional perturbations. To test this prediction, we modified the model by adding an external forcing term,
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
Simulations of these dynamics indeed demonstrate low-rank suppression (Figure 5a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 5b) in contrast to the model from [1]. These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics. Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
Discussion
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. Moreover, in the presence of high-dimensional noise or perturbations, low-rank networks often produce high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional responses to high-dimensional perturbations.
The fact that stable, strongly low-rank networks produce low-rank suppression might seem obvious in hindsight: Whenever $W_{0}$ is normal, stability implies that $W$ has a large, negative eigenvalues, so dynamics are highly suppressive in the direction of the corresponding eigenvectors (as in Figures 1 and 2b,c). However, when $W_{0}$ is highly non-normal $W_{0}$, the network still produces low-rank suppression even when eigenvalues are small in magnitude (Supplementary Figure S.2), so this argument alone does not explain low-rank suppression.
The fact that low-rank networks produce high-dimensional responses to high-dimensional inputs might also seem obvious in hindsight: Inputs drive the network, so high-dimensional input should drive high-dimensional activity regardless the internal network structure. However, our condition on the alignment of singular vectors shows that this explanation is not sufficient. When left- and right-singular vectors are not aligned ($\sigma_{P}\ll 1$), high-dimensional inputs drive low-dimensional responses (Figure 2f,g and Supplementary Figure S.2).
Our results have implications for networks in nature and applications. For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 17, 4, 18, 19, 20, 3, 5, 24]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural dynamics [41, 42, 43]. However, many of these recordings are made during the presentation of low-dimensional stimuli or motor tasks. A growing number of more recent studies have shown that neural activity is high-dimensional during the presentation of high-dimensional stimuli or tasks [13, 43, 44, 45, 46, 47]. Our results are consistent with the hypothesis that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [13]. Our results are also consistent with observations that neural populations are more sensitive to random perturbations than to perturbations that are aligned with low-dimensional network structure [36].
Beyond neuroscience, our results imply that perturbations or inputs to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space. This result can be used to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equation simulations – except for Figure 5 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 5 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
https://github.com/RobertRosenbaum/HighDimLowDimCode
Acknowledgments
We thank Ashok Litwin-Kumar for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
Supplementary Materials
Figure S.1: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau{\bf{z}}^{\prime}=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure S.2: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 1 except $W_{0}=|c|{\bf{u}}{\bf{v}}^{T}$ where ${\bf{u}}$ and ${\bf{v}}$ are independent random unit vectors. Also, in panel c, we used a perturbation of ${\bf{x}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ as given in Eq. (15)) and in panel e-g, we projected onto ${\bf{v}}$. The network exhibits low-rank suppression in response to perturbations parallel to ${\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ (panel c) even though the connectivity matrix, $W$, does not have strongly negative eigenvalues (minimum real part of the eigenvalues of $W$ here is $\textrm{re}(\lambda)_{\textrm{min}}=-0.9$ compared to $\textrm{re}(\lambda)_{\textrm{min}}=-10$ in Figure 1). However, the network produces low-dimensional responses to high-dimensional inputs (panel d).
Figure S.3: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (3), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $-0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure S.4: Low-dimensional dynamics in a modular network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the modular network from Figure 3.
Figure S.5: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 4.
Figure S.6: Low-dimensional dynamics in a spatial network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the spatial network from Figure 4.
Figure S.7: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor ($c=-2$), and the dynamics obey Eq. (3). The instability produced by taking $\rho>1$ generates chaotic dynamics [16, 3].
Figure S.8: Dynamics of an unstable, non-chaotic network with rank-one structure. Same as Figure 1 except $c=10$ and the dynamics obey Eq. (3). The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
Figure S.9: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1d–g except ${\bf{x}}(t)$ was made low-dimensional by multiplying it by a matrix with rank-one structure, ${\bf{x}}\leftarrow W_{x}{\bf{x}}$ where $W_{x}$ is a matrix generated identically to, but independently from $W$. The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$.
Figure S.10: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of ${\bf{z}}$ for the model in Figure 5.
S.1 Comparison to previous work on dynamics of low-rank networks
In this section we review the relationship between our models and analysis and the models and analyses from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 17, 4, 18, 19, 20], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
All three studies quoted above consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model. We next describe the salient properties that distinguish these models from ours.
Thibeault et al. [1] assume that their networks are self-contained and do not receive any time varying external input. Specifically, they explicitly restrict their dynamics to models of the form
$\frac{d{\bf{z}}}{dt}=F({\bf{z}},W{\bf{z}})$
which excludes the possibility of modeling a time-varying external input, like ${\bf{x}}(t)$ in our model. Models of this form cannot describe networks that are part of a larger network, or networks that are modulated by time-varying, external factors. Similar assumptions were made in one study by Landau and Sompolinsky [5]. In other work by Landau and Sompolinsky [3], external input was included, but this input was assumed to be perfectly aligned to the low-rank structure of the connectivity matrix and therefore low-dimensional.
The salient differences between our model and the models considered by Ostojic and colleagues [2, 17, 4, 18, 19, 20] are more subtle. Like us, they consider external inputs that are not aligned to the low-rank part of $W$. Also, like us, they take $W=W_{0}+W_{1}$ where $W_{1}$ is full rank with random entries and the variance of the entries in $W_{1}$ scale like $\mathcal{O}(1/N)$ so that the maximum singular value of $W_{1}$ scales like $\mathcal{O}(1)$.
However, in contrast to our models, the low-rank components of the networks considered by Ostojic and colleagues take the form
$W_{0}=\sum_{\mu=1}^{r}\frac{{\bf{m}}_{\mu}{\bf{n}}_{\mu}^{T}}{N}$ (6)
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (6), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) scale the same in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
This difference in scaling is acknowledged by Ostojic et al. [2, 17, 4, 18, 19, 20] who refer to their networks as “weakly low-rank.” Keeping with this terminology, we will use “weakly low-rank” in this section to refer to networks in which the singular values of the low-rank components ($W_{0}$) and high-rank components ($W_{1}$) both scale like $\mathcal{O}(1)$. We additionally use “strongly low-rank” to refer to networks like ours in which the singular values of $W_{0}$ diverge to $\infty$ as $N\to\infty$ (typically, they scale like $\mathcal{O}(\sqrt{N})$) while the singular values of $W_{1}$ scale like $\mathcal{O}(1)$.
To compare weakly and strongly low-rank networks, we can consider the simple rank-one example from Figure 1. Specifically,
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
while $W_{1}$ has unbiased random entries with variance $\rho/\sqrt{N}$ and dynamics obey Eq. (1). Strongly low-rank networks have $c\gg 1$ while $c\sim\mathcal{O}(1)$ in weakly low-rank networks. In strongly low-rank networks, stability requires that $c<0$, which is critical for low-rank suppression. In weakly low-rank networks, however, it is possible to take $c>0$ without causing instability if $c<1$. When $c>0$, weakly low-rank networks can respond to perturbations aligned to ${\bf{u}}$ more strongly than random perturbations (Supplementary Figure S.11), reversing the trend of low-rank suppression. When $c<0$ in weakly low-rank networks, perturbations in the direction of ${\bf{u}}$ can be suppressed relative to random perturbations, but the effect is weaker than in strongly low-rank networks (Supplementary Figure S.12).
Landau and Sompolinsky [5, 3] consider strongly low-rank networks. However, as noted above, their external input, ${\bf{x}}(t)$, is aligned to the low-rank part of $W$ and is therefore low-dimensional itself. Thibeault et al. [1] consider several network models, but their “rank-perturbed Gaussian” model is strongly low-rank and equivalent to the network structure we study in Figure 1. Complicating matters, Thibeault et al. directly compare their rank-perturbed Gaussian model to the networks in Ostojic et al., despite the fact that the scaling of their low-rank parts differ by a magnitude of $\sqrt{N}$. Specifically, in Thibeault et al. [1], the low-rank part is defined by Eq. (6) where ${\bf{m}}$ and ${\bf{n}}$ are unbiased Gaussian random vectors. The entries of ${\bf{m}}$ have $\mathcal{O}(1/N)$ variance while the entries of ${\bf{n}}$ have $\mathcal{O}(1)$ variance, so the entries of $W_{0}$ have $\mathcal{O}(1/N)$ variance, in contrast to the $\mathcal{O}(1/N^{2})$ variance used by Ostojic et al. [2]. Importantly, this means that the singular values of $W_{0}$ are $\mathcal{O}(\sqrt{N})$ in the rank perturbed Gaussian model analyzed by Thibeault et al. [1], but $\mathcal{O}(1)$ in the models by Ostojic et al. Hence, the rank perturbed Gaussian models considered by Thibeault et al. are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except $c=0.75$.
Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=-0.75$.
S.2 Mathematical analysis of low-rank networks with external perturbations
In this section we present a mathematical analysis of low-rank suppression and high-dimensional dynamics in networks with effective low-rank connectivity and external perturbations. Our analysis applies to a broad class of dynamics. We begin by reviewing the precise assumptions made in our analysis.
S.2.1 Assumptions on the dynamics and definition of the effective connectivity matrix.
We start by assuming dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},{\bf{x}})$ (7)
for some smooth ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$ where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and ${\bf{x}}(t)$ is an $N\times 1$ vector of external forces, inputs, and/or noise. At the end of Section S.2.3, we show how our approach can be generalized to a larger class of causal, time invariant systems, but the dynamics in Eq. (7) are sufficient for all of the examples studied in this manuscript, and the majority of models studied in the literature on low-rank networks, so we focus on this class of dynamics for simplicity.
In general, it is not immediately clear how to define a network of interactions, $W$, from dynamics as general as those in Eq. (7). For example, in nonlinear systems, the effective connectivity between nodes can change with the state of the system. To circumvent this problem, we define the effective connectivity matrix, $W$, through a linearization around a stable equilibrium.
Specifically, we assume that there exists a time-constant forcing, ${\bf{x}}(t)={\bf{x}}_{0}$, that produces a stable steady state solution, ${\bf{z}}_{0}$. In other words,
${\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})={\bf{0}}.$
We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
$J_{{\bf{z}}}=\partial_{{\bf{z}}}F({\bf{z}}_{0},{\bf{x}}_{0})$
has eigenvalues with strictly negative real part. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
${\bf{x}}_{p}(t)={\bf{x}}_{0}+\epsilon{\bf{x}}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, ${\bf{x}}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}+\mathcal{O}(\epsilon^{2}).$ (8)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}(t)$ (9)
where $J_{{\bf{x}}}=\partial_{{\bf{x}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ is the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{x}}$ evaluated at the fixed point, and
$W=J_{{\bf{z}}}+I$ (10)
with $I$ the identity matrix and $J_{{\bf{z}}}=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{z}}$ evaluated at the fixed point. We interpret $W$ in Eq. (10) as the effective connectivity matrix.
It may seem counterintuitive at first to define $W=J_{{\bf{z}}}+I$ instead of $W=J_{{\bf{z}}}$. To understand our choice, note that when we choose $W=J_{{\bf{z}}}+I$, the perturbed system ${\bf{z}}_{p}(t)$ decays to the unperturbed fixed point, ${\bf{z}}_{0}(t)$, whenever there is no perturbation or connectivity (${\bf{x}}(t)=0$ and $W=0$). Moreover, when there is a perturbation without connectivity ($W=0$ and ${\bf{x}}\neq 0$), the perturbation converges to ${\bf{z}}=J_{{\bf{x}}}{\bf{x}}$. Hence, $W$ represents interactions that determine the dynamics under perturbations away from the fixed point. In any case, our overall results are not sensitive to this choice.
One caveat to our approach for defining effective connectivity, $W$, from the dynamics in Eq. (7) is that the input perturbation, ${\bf{x}}(t)$, must be weak enough and/or the dynamics must be sufficiently linear for the linearized dynamics to be accurate, i.e., for the $\mathcal{O}(\epsilon^{2})$ term in Eq. (8) to be small. Moreover, the dynamics of the unperturbed system must be stable near the steady state. Examples with unstable, nonlinear dynamics are considered in Supplementary Figures S.7 and S.8.
Another caveat is that multiplication of ${\bf{x}}$ by $J_{{\bf{x}}}$ in Eq. (9) could suppress dimensionality, for example if $J_{{\bf{x}}}$ is low-rank or effectively low-rank. This would correspond to an effectively feedforward “read-in” mechanism that lowers dimensionality. Specifically, the “effective input” $J_{{\bf{x}}}{\bf{x}}$ would be low-dimensional and suppressed in some directions. For example, consider dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+W_{x}{\bf{x}}$ (11)
where $W_{x}$ is some matrix. In this model, we have $J_{{\bf{x}}}=W_{x}$. Therefore, if $W_{x}$ were low-rank then the effective input, ${\bf{x}}_{eff}=W_{x}{\bf{x}}$, would be low-dimensional even when the raw input, ${\bf{x}}$, were high-dimensional. This scenario would promote low-dimensional dynamics in ${\bf{z}}(t)$. An example of this form is considered in Supplementary Figure S.9. In many models (including Eqs. (1) and (3)), the dynamics take the form ${\bf{F}}({\bf{z}},{\bf{x}})={\bf{F}}({\bf{z}})+{\bf{x}}$ so that $J_{{\bf{x}}}=I$ and this is not an issue. More generally, we assume that $J_{{\bf{x}}}$ is full rank in all of our analysis.
Before continuing, we compare our approach for relating dynamics to connectivity with the approach taken in Thibeault et al. [1]. They define dynamics that explicitly take the connectivity matrix into account,
$\frac{d{\bf{z}}}{dt}$ $={\bf{G}}({\bf{z}},{\bf{y}})$
${\bf{y}}$ $=U{\bf{z}}$
where $U$ is interpreted as the connectivity matrix (they use “$W$”, but we switched to “$U$” to avoid confusion with our definition). This formulation represents a self-contained network in the absence of external perturbations or noise, but it can easily be extended to allow external perturbations by defining
$\frac{d{\bf{z}}}{dt}={\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}}).$ (12)
However, Eq. (12) is not suitable for our purposes because $U$ does not necessarily capture all of the interactions between nodes. For example, the $j$th entry of ${\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}})$ could depend explicitly (through its first argument) on the $k$th entry of ${\bf{z}}$ for some $j\neq k$. This would represent an interaction between nodes $j$ and $k$, which are not accounted for in $U$. For example, we could have $d{\bf{z}}_{j}/dt={{\bf{G}}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{z}}_{k}+{\bf{y}}_{j}={\bf{z}}_{k}+U_{ji}{\bf{z}}_{i}$ for three distinct indices $i,j,k$ where we assume that $U_{jk}=0$. In this case, the connectivity matrix, $U$, correctly captures the influence of node ${\bf{z}}_{i}$ on node ${\bf{z}}_{j}$ through the entry $U_{ji}$. However, node ${\bf{z}}_{j}$ is also influenced by node ${\bf{z}}_{k}$ and this interaction is not captured by the matrix $U$ (as we take $U_{jk}=0$) because it is instead captured by the explicit dependence of ${\bf{G}}$ on ${\bf{z}}_{k}$ (the first argument of ${\bf{G}}$). This could, of course, be overcome by incorporating the dependence on ${\bf{z}}_{k}$ into $U$ (i.e., by setting $U_{jk}=1$), but Eq. (12) makes no assumption that all such interactions are included into $U$. More importantly, it might not be possible to account for higher order, nonlinear interactions using the formulation in Eq. (12). For example, we could have $d{\bf{z}}_{j}/dt={\bf{G}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{y}}_{k}{\bf{z}}_{i}=U_{jk}{\bf{z}}_{k}{\bf{z}}_{i}$ for three distinct indices $i,j,k$. In this case, the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ would not be accounted for in the matrix $U$, and it is not possible to account for both interactions ($i\to j$ and $k\to j$) in the matrix $U$ using the formulation in Eq. (12). Moreover, the strength of the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ depends on the magnitude of ${\bf{z}}_{k}$, which can change with the state of the network. These matters are handled in the mathematical analysis of Thibeault et al. [1] by bounding norms related to the Jacobian matrix of ${\bf{G}}$ with respect to ${\bf{z}}$ and ${\bf{y}}$, and they also separately consider higher order interactions. Our approach avoids these bounds on the Jacobian matrices at the expense of relying on a linearization near a steady state. Under the formulation from Eq. (12), if there is a fixed point at ${\bf{z}}_{0}$ in response to static input ${\bf{x}}_{0}$, then our interpretation of the connectivity matrix is given by
$W=\partial_{{\bf{z}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})+\partial_{{\bf{y}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})U+I$
where ${\bf{y}}_{0}=U{\bf{z}}_{0}$ and the two partial derivatives denote Jacobian matrices with respect to the first and second arguments, respectively. Hence, the two approaches can easily be directly related.
In summary, we consider the dynamics defined by Eq. (7), linearized around a stable steady state, which ultimately gives rise to the definition of network dynamics in Eq. (9). This approach has some caveats, which are discussed above.
S.2.2 Assumptions on the network structure.
The network represents $N$ interacting nodes, the interactions are defined by an $N\times N$ connectivity matrix, $W$, and we consider asymptotics in the $N\to\infty$ limit. Here, we specify the assumptions made about the structure of $W$ in the large $N$ limit.
Letting $\sigma_{1},\ldots,\sigma_{N}$ be the singular values of $W$, we assume that $\sigma_{k}\gg 1$ for $k=1,\ldots,r$ (in the sense that $\sigma_{k}\to\infty$ as $N\to\infty$) and that $\sigma_{k}\leq\mathcal{O}(1)$ for $k=r+1,\ldots,N$ (in the sense that $\sigma_{k}$ has a finite limit as $N\to\infty$). We also assume that $r\sim\mathcal{O}(1)$ so that $r\ll N$. In summary, $W$ has a small number of large singular values and the remaining singular values are much smaller.
Under these assumptions, $W$ can be decomposed as [1]
$W=W_{0}+W_{1}$
where $W_{0}$ has rank $r$ and we write its singular value decomposition as
$W_{0}=U\Sigma V^{T}.$
Here, $U$ and $V$ are $N\times r$ orthonormal matrices and $\Sigma$ is an $r\times r$ diagonal matrix with diagonal entries, $\Sigma_{kk}=\sigma_{k}$. The left and right singular matrices, $U$ and $V$, can be written as
$U=\left[{\bf{u}}_{1}\;\;{\bf{u}}_{2}\;\cdots\;{\bf{u}}_{r}\right]$
and
$V=\left[{\bf{v}}_{1}\;\;{\bf{v}}_{2}\;\cdots\;{\bf{v}}_{r}\right]$
where their columns, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$, are the left and right singular vectors, which satisfy $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$ and ${\bf{u}}_{k}\cdot{\bf{u}}_{j}={\bf{v}}_{k}\cdot{\bf{v}}_{j}=0$ for $j\neq k$. We further assume that $W_{1}$ is a random matrix, which is almost surely full rank and is statistically independent from $W_{0}$. Finally, we assume that the eigenvalues and singular values of $W_{1}$ are smaller than $1$ in magnitude and bounded away from $1$ ($\sigma_{W_{1}},|\lambda_{W_{1}}|<1-\delta$ for some fixed $\delta>0$) with probability $1$.
S.2.3 Analysis of low-rank suppression.
We now analyze the conditions under which low-rank suppression occurs. Our goal here is to understand the conditions under which the dynamics produce low-rank suppression like that observed in Figure 1c. In the next section, we will derive conditions on high-dimensional dynamics in the presence of stochastic perturbations such as those considered in Figure 1d–g.
We begin by considering the case in which ${\bf{x}}(t)$ is static or slow, so we can use a steady state or quasi-steady state approximation. Later in this section, we consider low-rank suppression with transient and fast perturbations.
If ${\bf{x}}(t)={\bf{x}}$ is static (such as during the duration of each stimulus in Figure 1c), then the network response converges to solutions of the perturbed steady state, which from Eq. (9) satisfies
${\bf{z}}=W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}$
and therefore
${\bf{z}}=[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}.$ (13)
If ${\bf{x}}(t)$ is not static, but changes much more slowly than $\tau$, then ${\bf{z}}(t)$ closely tracks the quasi-steady state given by Eq. (13).
Low-rank suppression, like that observed in Figure 1c, occurs whenever there is an input, ${\bf{x}}$, that is strongly suppressed by the recurrent dynamics in comparison to a random input (as in Figure 1c). In other words, we can say that low-rank suppression occurs whenever there is a raw input, ${\bf{x}}$, such that $\|{\bf{x}}\|=\mathcal{O}(1)$ and
$\|{\bf{z}}\|=\|[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}\|\ll\|{\bf{z}}_{rand}\|=\|[I-W]^{-1}{\bf{u}}_{rand}\|$
with high probability whenever ${\bf{u}}_{rand}$ is a random unit vector, independent from $W_{0}$, $W_{1}$, and $J_{{\bf{x}}}$. To simplify this statement, first note that $\|[I-W]^{-1}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ whenever ${\bf{u}}_{rand}$ is independent from $W_{1}$. Moreover, since we assume that $J_{{\bf{x}}}$ is full rank, we can write the condition in terms of the effective input,
${\bf{x}}_{eff}=J_{{\bf{x}}}{\bf{x}}$
instead of the raw input, ${\bf{x}}$.
In particular, low-rank suppression occurs whenever there is a vector ${\bf{x}}_{eff}$ such that
$\|{\bf{z}}\|=\|[I-W]^{-1}{\bf{x}}_{eff}\|\ll\|{\bf{x}}_{eff}\|.$ (14)
We take this property to be the definition of low-rank suppression.
For the dynamics that we use in our simulations, note that $J_{{\bf{x}}}=I$ so ${\bf{x}}_{eff}={\bf{x}}$. We discuss some caveats of the distinction between ${\bf{x}}$ and ${\bf{x}}_{eff}$ at the end of this section. We now prove that low-rank suppression arises under the assumptions from Sections S.2.2S.2.3.
Claim 1.
Under the model and assumptions of Sections S.2.2S.2.3, the network exhibits low-rank suppression.
Proof.
Let $k\in\{1,\ldots,r\}$ and take
${\bf{x}}_{eff}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (15)
We claim that the response to this input is given by
${\bf{z}}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
To see this, we only need to show that $[I-W]{\bf{z}}={\bf{x}}$. We have
${\bf{z}}$ $=\frac{W_{0}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$=\frac{U\Sigma V^{T}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{x}}_{eff}.$
Now note that $\|{\bf{x}}_{eff}\|=1+o(1)$ and $\|{\bf{z}}\|\ll 1$ because $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, $\|W_{1}\|_{2}=\mathcal{O}(1)$, and $\sigma_{k}\gg 1$. Therefore $\|{\bf{z}}\|\ll\|{\bf{x}}_{eff}\|$. ∎
Before continuing, we will make a few observations about this claim and its proof. First note that the definition of low-rank suppression in Eq. (14) is equivalent to the statement that the matrix
$A=[I-W]^{-1}$
has at least one asymptotically small singular value,
$\sigma_{A}\ll 1.$
Singular values commute with matrix inverses, so the singular values of $A$ are given by
$\sigma_{A}=\frac{1}{\sigma_{Q}}$
where $\sigma_{Q}$ are the singular values of the matrix
$Q=I-W.$
Therefore, low-rank suppression occurs whenever $Q$ has at least one asymptotically large singular value,
$\sigma_{Q}\gg 1.$
Since $W$ has an asymptotically large singular value, low-rank suppression occurs because this large singular value is conserved by adding the identity matrix. However, unlike eigenvalues, singular values are not transformed in a simple way by adding the identity matrix, in other words, we do not generally have $\sigma_{Q}=1-\sigma_{W}$. If $W$ were a normal matrix, then we would have $\sigma_{Q}=|1-\sigma_{W}|$, so the emergence of low-rank suppression would have been easier to prove, but this is not in general true of non-normal matrices. The potentially complicated relationship between the singular values of $W$ and those of $Q$ is more critical for the analysis of high-dimensional dynamics in the next section.
The proof of Claim 1 also provides insight into the input and response directions that are suppressed in low-rank suppression. Note that ${\bf{x}}_{eff}={\bf{u}}_{k}+o(1)$ and ${\bf{z}}\propto{\bf{v}}_{k}$, which implies that inputs approximately aligned to the column space of $U$ (equivalently, the column space of $W_{0}$) give rise to suppressed responses aligned to the column space of $V$ (equivalently, the row space of $W_{0}$).
Finally, it is worth discussing our use of ${\bf{x}}_{eff}$ instead of ${\bf{x}}$ in the definition of low-rank suppression. For our simulations ${\bf{x}}={\bf{x}}_{eff}$ so there is no issue. However, one could consider a model for which $J_{{\bf{x}}}$ is full rank, but amplifies some directions more than others. For example, consider the model in Eq. (11) with a $W_{x}$ chosen to amplify inputs in every direction parallel to the column space of $W$, i.e.., if $\|W_{x}{\bf{u}}_{k}\|\gg 1$ for $k=1,\ldots,r$ whereas $\|W_{x}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ for random directions. In this very specific case, the amplification by $W_{x}$ would compete with the suppression by $A=[I-W]^{-1}$, so the raw input, ${\bf{x}}$, would need to be smaller in the direction of ${\bf{u}}_{k}$ to obtain an ${\bf{x}}_{eff}$ that is $\mathcal{O}(1)$ in that direction. In this case, the suppression by recurrent dynamics would compete with the amplification by a feedforward read-in. Importantly, the recurrent dynamics would still be suppressive to effective inputs aligned to ${\bf{u}}_{k}$. We focus on effective inputs in our definition of low-rank suppression to avoid such an issue, but it would only arise in a very specific situation where $J_{{\bf{x}}}$ amplifies exactly the directions that $A=[I-W]^{-1}$ suppresses.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-static approximation from Eq. (13) in this case. Instead, we take the Laplace transform in Eq. (9) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (16)
where
$\hat{H}(s)$ $=(1+\tau s)^{-1}\left[J_{{\bf{z}}}+I\right]=(1+\tau s)^{-1}W$ (17)
$\hat{B}(s)$ $=(1+\tau s)^{-1}J_{{\bf{x}}}$
[…]
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$, i.e.,
$W=\hat{H}(s)$
because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. Solving the implicit Eq. (16) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-W]^{-1}\hat{B}\hat{{\bf{x}}}.$ (18)
Since $\hat{B}$ quantifies feedforward, read-in dynamics, we can interpret $\hat{B}\hat{{\bf{x}}}$ as effective input in the Laplace domain, ${\bf{\hat{x}}}_{eff}=\hat{B}\hat{{\bf{x}}}$. This is analogous to our treatment of $J_{{\bf{x}}}{\bf{x}}$ above.
Eq. (18) is identical to Eq. (13) except that Eq. (18) quantifies interactions at individual Laplace modes, $s$. Hence, Claim 1 applies to situations in which ${\bf{x}}(t)$ is transient and fast, so long as the effective connectivity, $W$, and effective input, ${\bf{x}}_{eff}$ are interpreted in the Laplace domain ($W=H(s)$ and ${\bf{x}}_{eff}=\hat{B}(s)\hat{{\bf{x}}}(s)$).
Moreover, the approach of working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (7). The dynamics in Eq. (7) capture a large class of models, but are not fully general. Specifically, in Eq. (7), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$ and the values of $x(s)$ for $s\leq t_{1}$. However, some systems like integro-differential equations and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (7). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (19)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivriant function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (19) encompasses an extremely broad class of systems, including those described by Eq. (7).
We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq. (19). As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$. However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$. We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
A first order Volterra expansion gives Eq. (16) as an implicit equation for ${\bf{z}}(t)$. The matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (9), we obtain the $\hat{H}$ and $\hat{B}$ defined in Eq. (17). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (19).
S.2.4 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–g. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
[…]
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. This is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (9) to get [48, 49, 34, 35, 50]
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}J_{{\bf{x}}}\widetilde{C}^{{\bf{x}}}(f)J_{{\bf{x}}}^{T}[\widetilde{r}(f)I-W]^{-*}$ (20)
where
$\widetilde{r}(f)=1-2\pi\tau fi$
is scalar.
Note that Eq. (20) can be derived from Eqs. (18) and (17) by first taking $s=-2\pi if$ in Eq. (18) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the results derived in this section apply to the general class of dynamics defined in Eq. (19).
Eq. (20) quantifies the covariance structure of ${\bf{z}}(t)$ at any given frequency mode, $f$, but we are often specifically interested in the zero-lag temporal covariance,
[…]
From Eq. (20), we therefore have
[…]
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (21) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}J_{{\bf{x}}}\overline{C}^{{\bf{x}}}J_{{\bf{x}}}^{T}[I-W]^{-T}.$ (22)
When principal component analysis is applied to ${\bf{z}}(t)$, the variance explained by each principal component is given by the ordered list of eigenvalues of the covariance matrix, $\overline{C}^{{\bf{z}}}$. From Eq. (22), we see that these eigenvalues are proportional to the squares of the singular values of the matrix
$R=[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}.$
where $\sqrt{\overline{C}^{{\bf{x}}}}$ is the matrix square root of the symmetric positive definite covariance matrix, $\overline{C}^{{\bf{x}}}$ (distinct from the entry-wise square root in general). Therefore, the decay of the variance explained by each principal component of ${\bf{z}}$ (as in the blue dots in Figure 1d) are described by the squared singular values of $R$.
In the models we consider, each ${\bf{x}}_{j}(t)$ is an i.i.d. stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$. Moreover, in the networks we consider, $J_{{\bf{x}}}=I$ is equal to the identity matrix. Therefore, for our models,
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}v$
where $v$ is a scalar.
In more general classes of models in which $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are not multiples of the identity matrix, dimensionality of the dynamics of ${\bf{z}}(t)$ could be reduced whenever $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) or a low-rank read-in mechanism ($J_{{\bf{x}}}$ low-dimensional; see Supplementary Figure S.9 and discussion in previous sections), and would not be related to the actual dynamics of interactions within the network. Therefore, our focus on the effective rank of $[I-W]^{-1}$ in place of $[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}$ is justified even in more general settings.
In conclusion, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
$A=[I-W]^{-1}.$
More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
[…]
where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
Therefore, ${\bf{z}}(t)$ is low-dimensional whenever $A$ has a small number of large singular values. As above, we can use the fact that singular values commute with matrix inverses to write our conclusions in terms of
$Q=I-W.$
Specifically,
[…]
where $\sigma_{Q,N-k}$ is the $(N-k)$th singular value of $Q$, i.e., the $k$th from the last singular value (assuming singular values are sorted in decreasing order).
Figure S.13: Comparison between theory and simulations for shorter and longer simulations. a) Same as Figure 1d except we added the theoretical values form Eq. (23) (red dots) and the values obtained from Eq. (24) (green dots). The network was simulated for $T=5\times 10^{3}$ time units. b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq. (23) as red dots (Supplementary Figure S.13a). Surprisingly, the theory did not closely match the simulations. We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$). We suspected that Eq. (22) would be accurate when $\overline{C}^{\bf{x}}$ is replace by the empirical covariance matrix of ${\bf{x}}(t)$. Under this substitution, Eq. (23) would be replaced by
[…]
where
$U=[I-W]\sqrt{{\overline{C}}^{\bf{x}}}^{-1}$
and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
[…]
where
$\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
is the sampled value of $R$ (assuming $J^{{\bf{x}}}=I$) and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used. Using Eq. (24) gives a much more accurate prediction (Supplementary Figure S.13a, green dots) This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$. We next increased the simulation time ten-fold to $T=5\times 10^{4}$. In this case, the original Eq. (23) was accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
As concluded at the end of the previous section, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values). Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values). Conversely, high-dimensional dynamics occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
Combining these conclusions with those reached in Section S.2.3, we can summarize as follows
Low-rank suppression (as in Figure 1c) occurs whenever $Q=I-W$ has at least one asymptotically large singular value.
High-dimensional dynamics (as in Figure 1d) occur whenever $Q=I-W$ lacks any asymptotically small singular values.
In the previous section, we showed that $Q=I-W$ has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank. Conditions under which $Q=I-W$ lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple. Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
$P=V^{T}U,$
which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$. Specifically, $P_{jk}={\bf{v}}_{k}\cdot{\bf{u}}_{k}$ so that $P_{jk}=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$. Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$. We next show that if all singular values of $P$ are $\mathcal{O}(1)$ then the network exhibits high-dimensional responses to high-dimensional inputs.
Claim 2.
Under the model and assumptions of Sections S.2.2 and S.2.4, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
Proof.
We will prove this claim by proving its contrapositive:
If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d). From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever $Q=I-W$ has an asymptotically small singular value. Therefore, our original claim is equivalent to the following:
If $Q=I-W$ has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
We will prove this version of the claim directly. Assume that $Q$ has an asymptotically small singular value. Then there is a ${\bf{z}}$ satisfying $\|{\bf{z}}\|=1$ and
$(I-W){\bf{z}}=o(1)$
where the notation $o(1)$ means that $\|(I-W){\bf{z}}\|\to 0$ as $N\to\infty$. It is sufficient to show that there is a ${\bf{y}}$ with $\|{\bf{y}}\|=1+o(1)$ satisfying
$\|P{\bf{y}}\|=o(1).$
We have that
$W_{0}{\bf{z}}+W_{1}{\bf{z}}={\bf{z}}+o(1)$
Multiplying both sides on the left by $U^{T}$ gives
$\Sigma V^{T}{\bf{z}}+U^{T}W_{1}{\bf{z}}=U^{T}{\bf{z}}+o(1)$
Since $W_{1}$ is random and independent from $W_{0}$ and $U$, and since $U^{T}$ projects from $N$ to $r\ll N$ dimensions, the term $U^{T}W_{1}{\bf{z}}$ represents the projection of a random vector onto the low-dimensional column space of $U$, so $\|U^{T}W_{1}{\bf{z}}\|=o(1)$. We therefore have
$\Sigma V^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
We next claim that ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. To see why this is true, note that $UU^{T}{\bf{z}}$ is the orthogonal projection of ${\bf{z}}$ onto the column space of $W_{0}$. Since $W_{0}$ dominates $W=W_{0}+W_{1}$ (and therefore the the column space of $W$ is dominated by that of $W_{0}$) and $W{\bf{z}}\approx{\bf{z}}$, we may conclude that ${\bf{z}}$ lies predominantly in the column space of $W$ and therefore of $W_{0}$. In other words, ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. Hence, we can rewrite the equality above as
$\Sigma V^{T}UU^{T}{\bf{z}}=U^{T}{\bf{z}}+o(1).$
which reduces to
$\Sigma P{\bf{y}}={\bf{y}}+o(1).$
where $P=V^{T}U$ and ${\bf{y}}=U^{T}{\bf{z}}$. Note again that $\|{\bf{y}}\|=\|U^{T}{\bf{z}}\|=1+o(1)$ since $\|{\bf{z}}\|=1$ and ${\bf{z}}=UU^{T}{\bf{z}}+o(1)$. We then have that
$\|P{\bf{y}}\|=\|\Sigma^{-1}{\bf{y}}\|+o(1)=o(1)$
because $\Sigma^{-1}$ is a diagonal matrix with $o(1)$ terms on the diagonal. ∎
The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2g. Since ${\bf{z}}$ is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$. Indeed, in Figure 2g, the angle between ${\bf{u}}_{2}$ and the dominant principle component direction of ${\bf{z}}$ is $18^{\circ}$.
The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$. However, note that Eq. (20) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies. Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq. (22). The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq. (20) instead of the identity itself, as in Eq. (22). This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).

### v2

@@
 Abstract
-Many networks that arise in nature and applications are effectively low-dimensional in the sense that their connectivity structure is dominated by a few dimensions.
-It is natural to expect that dynamics on such networks might also be low-dimensional.
-Indeed, recent results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or noise.
+Many networks in nature and applications have an approximate low-rank structure in the sense that their connectivity structure is dominated by a few dimensions.
+It is natural to expect that dynamics on such networks would also be low-dimensional.
+Indeed, theoretical results show that low-rank networks produce low-dimensional dynamics whenever the network is isolated from external perturbations or input.
 However, networks in nature are rarely isolated.
-We show that recurrent networks with low-rank structure often produce high-dimensional dynamics in the presence of high-dimensional perturbations.
-Counter to intuition, dynamics in these networks are suppressed in directions that are aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” Our results clarify important, but counterintuitive relationships between a network’s connectivity structure and the structure of the dynamics it generates.
+Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations.
+We find that dynamics in such networks can be high- or low-dimensional and we derive precise conditions on the network structure under which dynamics are high-dimensional.
+In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression.
+Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
 Introduction
-Recent work shows that many networks arising in nature and in applications are effectively low-rank in the sense that their connectivity matrices have a small number of larger singular values, while the remaining singular values are much smaller [1].
-What are the implications of low-rank network structure on the dimensionality of network dynamics?
-Several studies have shown that networks with low-rank structure tend to produce low-dimensional dynamics [1, 2, 3, 4].
-However, these studies rely on specific assumptions about the network model.
-Some studies assume that the low-rank part of the network is weak [2, 4], others assume that the network is self-contained in the sense that it operates in the absence of internal noise or external perturbations or inputs [5, 1], while other work assumes that external perturbations are perfectly aligned with the network’s low-rank structure [3].
-In reality, networks in nature are rarely isolated from perturbations, and these perturbations are not necessarily aligned with the network’s low-rank structure.
-We use mathematical analysis and computer simulations to show that, in the presence of high-dimensional inputs or perturbations, many networks with low-dimensional structure produce high-dimensional dynamics.
-Moreover, perhaps counterintuitively, these networks suppress variability in directions that are aligned with their low-rank structure relative to random directions, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity structure promote low-rank suppression and high-dimensional dynamics.
-Finally, we demonstrate our conclusions in a real world epidemiological network with low-rank structure.
-Our conclusions have important implications for the interpretation of low-dimensional structure in recurrent networks.
-In neuroscience, our results offer a mathematical explanation of the widely observed balance between excitatory and inhibitory synaptic currents [6, 7, 8, 9, 10, 11, 12] and also explain why neural populations generate high-dimensional responses to high-dimensional stimuli [13].
-Beyond neuroscience, our results imply that randomly structured perturbations are often most effective at driving activity in networks with low-rank structure.
+Recent work shows that many networks arising in nature and applications have low-dimensional structure [1].
+This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
+In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15].
+However, many of these recordings are made in the context of low-dimensional stimuli or behavior.
+A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20].
+These observations raise the important point that the dimensionality of dynamics on a network likely depends on the structure of the network’s external input, in addition to the network’s internal structure.
+Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [1, 5, 12].
+Other work considers external input, but assumes that the input is perfectly aligned with the network’s low-dimensional structure [3].
+In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure.
+Other work [2, 4] considers the case of general external input, but the recurrent network is assumed to be weakly low-rank in the sense that the largest singular values are not asymptotically larger than the bulk of the singular values.
+Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk.
+We use mathematical analysis to derive precise conditions under which linearized dynamics in these networks produce high-dimensional dynamics.
+Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity promote strongly low-rank structure, high-dimensional dynamics, and low-rank suppression.
+We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3].
+Finally, we demonstrate our conclusions in dynamics on a real epidemiological network.
+Our conclusions have important implications for the interpretation of low-dimensional network structure.
+In neuroscience, our results can explain why neural populations generate high-dimensional responses to high-dimensional stimuli and tasks [16].
+Our results also generalize and extend the theory of excitatory-inhibitory balance [21, 3], and amplification arising from breaks in this balance [22, 23, 24].
+Beyond neuroscience, our results imply that perturbations misaligned to low-rank network structure are most effective at driving responses.
+This counterintuitive observation could be used to design more effective interventions to epidemiological, biological, social, and other networks.
 Results
@@
 $\tau\frac{d{\bf{z}}}{dt}$ $=-{\bf{z}}+W{\bf{z}}+{\bf{x}}.$ (1)
-where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and we take $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
+where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
 $W=W_{0}+W_{1}$
@@
 $|c|\gg\rho$
-then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are bounded by $2\rho$ (Figure 1b).
-Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane [14, 15].
+then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (bounded by $2\rho$; Figure 1b).
+Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [25, 26].
 Importantly, stability of the network dynamics requires that $\rho<1$ and $c<1$ [27].
-Since low-rank structure requires $|c|\gg\rho$, stability requires $c<0$ when $\rho\sim\mathcal{O}(1)$.
-Here, we take $\rho=0.5$ and $c=-10$.
+If the dynamics were unstable, then $\|{\bf{z}}(t)\|$ would grow exponentially toward $\infty$, so we only consider parameter regimes with stable dynamics.
+Since strong low-rank structure also requires $|c|\gg\rho$, stability requires $c<0$ when $\rho=\mathcal{O}(1)$.
+In simulations here, we take $\rho=0.5$ and c=10c=-10.
 We consider more general network structures later.
-One might expect a network with connectivity $W$ to produce low-dimensional dynamics due to its low-rank structure.
-Indeed, a feedforward network with the same connectivity matrix produces approximately one-dimensional dynamics because it amplifies inputs aligned to ${\bf{u}}$ relative to other directions (Supplementary Figure S.1).
-We next demonstrate that this seemingly natural conclusion does not carry over to recurrent network dynamics.
+We first provided a high-dimensional input, ${\bf{x}}(t)$.
+Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input.
+Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}.
+Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Supplementary Figure S.1).
+However, this conclusion does not necessarily carry over to recurrent networks.
+In the recurrent network, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$.
+The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d).
+Perhaps surprisingly, this weakest principal component direction was closely aligned to $\bf{u}$} (the angle was less than $8^{\circ}$).
+Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of $\bf{u}$} was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
 Figure 1: Response properties of a recurrent network with rank-one structure.
@@
 b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure.
-c) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$).
+c) The eigenvalues of $W$ have a dominant, negative term.
 d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue).
 e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}.
-g) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of ${\bf{u}}$.
-We first simulated the network with two different input patterns. The “aligned” input pattern was exactly aligned with the low-rank component of $W$,
+g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$).
+h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
+To better understand these results, we next simulated the network with two different external input patterns. One pattern was aligned with the low-dimensional structure of $W$,
 ${\bf{x}}_{\textrm{aligned}}={\bf{u}},$
-while the “random” input pattern is aligned with a random direction,
+while the other had a random direction,
 ${\bf{x}}_{\textrm{random}}={\bf{u}}_{\textrm{rand}},$
-which is generated identically to, but independently from ${\bf{u}}$.
-For large $N$, ${\bf{u}}$ and ${\bf{u}}_{\textrm{rand}}$ are approximately orthogonal.
-Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network.
-Perhaps surprisingly, we observed exactly the opposite (Figure 1c): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
-To demonstrate some consequences of this effect, we next provided a high-dimensional input, ${\bf{x}}(t)$, that varied smoothly in time.
-Specifically, each ${\bf{x}}_{j}(t)$ was an i.i.d.
-realization of a smooth, unbiased, stationary Gaussian process.
-This can model internal noise or irregular external input to the network.
-Conventional wisdom and intuition might lead us to expect the network response to be low-dimensional and dominated by variability in the direction of ${\bf{u}}$, owing to the low-rank structure of $W$.
-In contrast, the variance explained by the principal components of ${\bf{z}}(t)$ decayed similarly to those of ${\bf{x}}(t)$ (Figure 1d), indicating that ${\bf{z}}(t)$ was high-dimensional like ${\bf{x}}(t)$.
-The only exception was the principal component that explained the least variance in ${\bf{z}}(t)$, which was much weaker than the other principal components (last blue dot in Figure 1d).
-Perhaps surprisingly, this weakest principal component direction was closely aligned to ${\bf{u}}$ (the angle between ${\bf{u}}$ and the corresponding principal component direction was less than $8^{\circ}$).
-Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of ${\bf{u}}$ was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
-We propose the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to dominant the low-rank structure of a network are suppressed by the network’s dynamics.
-In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional perturbations, as in Figure 1d.
-In this manuscript, we derive conditions under which low-rank suppression and high-dimensional dynamics arise in low-rank networks.
-We then show that these conditions are met by many natural low-rank network structures.
-The results in Figure 1 contrast with previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1].
-In some of these studies, external input or perturbations to the network are assumed to be absent [1, 5] or perfectly aligned with the low-rank component of the network connectivity [3].
-In other studies [2, 4], the network is assumed to have only a weak low-rank component ($|c|\sim\rho$).
-None of these assumptions are consistent with the model considered in Figure 1.
-See Supplementary Text S.1 for a more detailed review of this previous work and its relation to ours.
-A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady state solution to Eq. (1),
+which was generated identically to, but independently from $\bf{u}$}.
+Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Supplementary Figure S.1).
+In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
+We use the term “low-rank suppression” to refer to this phenomenon in which inputs aligned to the low-rank structure of a network are suppressed by the network’s dynamics.
+In the absence of other directions that are amplified by the network, low-rank suppression leads to high-dimensional responses to high-dimensional inputs (as in Figure 1d).
+The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1].
+In Supplementary Text S.1, we provide a detailed review of this previous work and its relation to ours.
+A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady-state-solution to Eq. (1),
 ${\bf{z}}=[I-W]^{-1}{\bf{x}}.$ (2)
-Convergence to the steady state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq.
+Convergence to the steady-state requires that ${\bf{x}}(t)={\bf{x}}$ is static, but as long as ${\bf{x}}(t)$ varies more slowly than $\tau=1$, solutions approximately track the quasi-steady state given by Eq.
 (2).
 Because of the matrix inverse in Eq.
-(2), the large singular value of $W$ near $|c|$ produces a small singular value of $[I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near ${\bf{u}}$.
+(2), the large singular value of $W$ near $|c|$ produces a small singular value of $I-W]^{-1}$ near $1/(1-|c|)\approx 1/|c|$ with left and right singular vectors near $\bf{u}$}.
 Hence, the recurrent network suppresses inputs in the direction of $\bf{u}$}.
-A related explanation is that the Jacobian matrix $J=W-I$ of Eq.
-(1) has a large, negative eigenvalue near $c-1<0$ with associated eigenvector near ${\bf{u}}$ so the dynamics in Eq.
+A related explanation is that the Jacobian matrix J=WIJ=W-I of Eq.
+(1) has a large, negative eigenvalue near c1<0c-1<0 with associated eigenvector near $\bf{u}$} so the dynamics in Eq.
 (1) are highly compressive along $\bf{u}$}.
-These explanations depend on specific properties of the model considered here, but we later describe more general conditions under which low-rank suppression and high-dimensional dynamics occur.
-An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with network interactions in the direction of ${\bf{u}}$.
+This second explanation relies on the assumption that $W_{0}$ is a symmetric matrix (or at least a normal matrix) but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
+An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}.
 More precisely, note that $W{\bf{z}}(t)$ in Eq.
-(1) can be interpreted as a vector of local inputs to each node whereas ${\bf{x}}(t)$ is external input, and ${\bf{z}}(t)$ is the network response.
-In the qausi-steady state, ${\bf{z}}\approx W{\bf{z}}+{\bf{x}}$.
-Under low-rank suppression, local input cancels nearly perfectly with external input ($W{\bf{z}}\approx-{\bf{x}}$) in the direction of ${\bf{u}}$ so that the response, ${\bf{z}}$, is weak in this direction (Figure 1g).
-This cancellation is a key feature of low-rank suppression.
-Conditions for low-rank suppression and high-dimensional dynamics
-Above, we considered a simple example of a network with rank-one structure.
-In Supplementary Text S.2, we derive conditions for low-rank suppression and high-dimensional dynamics in a general class of models that includes all dynamics of the form $\tau d{\bf{z}}/dt={\bf{F}}({\bf{z}},{\bf{x}})$.
-Here, we summarize and interpret those results.
-Low-rank suppression (as in Figure 1c) is a more general phenomenon than high-dimensional dynamics (as in Figure 1d).
-In Supplementary Text S.2, we show that low-rank suppression arises under two assumptions:
-  1. 1.
-    Assumptions on singular values.
-First, we assume that the network is strongly low-rank in the sense that the connectivity matrix $W$ has a small number of large singular values and the rest are much smaller (Figure 2a).
-Specifically, $W=W_{0}+W_{1}$ where $W_{0}$ has rank $r\ll N$ with large singular values ($\sigma_{W_{0}}\gg 1$) while $W_{1}$ has rank $N$ with small or moderate singular values ($\sigma_{W_{1}}\leq\mathcal{O}(1)$).
-“Weakly low-rank” networks for which $\sigma_{k}\sim\mathcal{O}(1)$ are studied in other work [2, 17, 4, 18, 19, 20] and do not necessarily produce low-rank suppression (see Supplementary Text S.1).
-  2. 2.
-    Assumptions on dynamics.
-In addition, our analysis relies on a linearization of the dynamics around a stable equilibrium.
-This requires that ${\bf{z}}(t)$ settles to a stable equilibrium whenever ${\bf{x}}(t)=0$, and that ${\bf{x}}(t)$ is sufficiently weak or the dynamics of ${\bf{z}}(t)$ are sufficiently linear.
-Nonlinear dynamics away from a stable equilibrium can produce any dynamics in principle [21, 22, 23] (see Supplementary Figures S.7 and S.8 for examples of unstable network states).
-Under these two assumptions, low-rank suppression emerges in the sense that there are inputs, ${\bf{x}}$, approximately parallel to the column space of $U$, that produce much weaker responses than inputs at random directions (as in Figure 1c).
-The emergence of high-dimensional dynamics (as in Figure 1d) requires additional assumptions:
-  1. 3.
-    Assumptions on input perturbations.
-We assume that ${\bf{x}}(t)$ is a smooth, high-dimensional stochastic input.
-In simulations, we take each ${\bf{x}}_{j}(t)$ to be an i.i.d.
-Gaussian process.
-Low-dimensional inputs, ${\bf{x}}(t)$, can produce low-dimensional dynamics in networks that would otherwise produce high-dimensional dynamics (Supplementary Figure S.9).
-  2. 4.
-    Conditions on the alignment of singular vectors.
-Under the assumptions above, the emergence of high-dimensional dynamics depends on the structure of the low-rank part, $W_{0}$, of the connectivity.
-Specifically, let
+(1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response.
+In the qausi-steady state,
+${\bf{z}}\approx W{\bf{z}}+{\bf{x}}.$ (3)
+In other words, $\bf{z}$} tracks the sum of internal and external inputs in the quasi-steady state.
+Since $W$ has a large singular value with left- and right singular values aligned to $\bf{u}$}, multiplication by $W$ amplifies the direction $\bf{u}$}.
+Therefore, the internal input, $W{\bf{z}}$, is large in the direction of $\bf{u}$} whenever $\bf{z}$} is moderate in the direction of $\bf{u}$}.
+In other words, $|{\bf{u}}\cdot W{\bf{z}}|\gg|{\bf{u}}\cdot{\bf{z}}|$.
+This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq.
+(3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}.
+This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}.
+Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h).
+We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [21, 28] as we will show later.
+Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
+Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue.
+Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input.
+We next address these counterpoints with an example in which $W_{0}$ is highly non-normal.
+In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
+Figure 2: Response properties of a non-normal recurrent network with rank-one structure.
+Same as Figure 1 except the low-rank part of $W$ is non-normal.
+The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c).
+The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
+We specifically consider an example in which the model and all parameters are the same as in Figure 1 except that
+$W_{0}=c{\bf{u}}{\bf{v}}^{T}$
+where $\bf{u}$} and $\bf{v}$} are random, orthogonal vectors with $\|{\bf{u}}\|=\|{\bf{v}}\|=1$.
+This matrix is highly non-normal in the sense that its left- and right-singular vectors are orthogonal, and its non-zero eigenvalue does not have a simple relationship to its singular values.
+The distribution of singular values of $W=W_{0}+W_{1}$ are identical to those from our previous example (Figure 2b; compare to Figure 1b), but the eigenvalues are different (Figure 2c; compare to Figure 1c).
+Specifically, the eigenvalues no longer have a dominant negative outlier.
+They are all $\mathcal{O}(1)$ in magnitude and clustered in a single bulk.
+Repeating the simulations from above for this network, we found that the response to high-dimensional input was low-dimensional in the sense that variability was much greater in one direction than all others: Over 40% of the variance in ${\bf{z}}(t)$ was captured by a single principle component (first blue dot in Figure 2d), compared to less than 2% for any orthogonal direction (the remaining blue dots).
+This dominant direction was partially aligned to $\bf{u}$} (the angle between the dominant direction and $\bf{u}$} was $31^{\circ}$).
+Consistent with this observation, variability was $100$ times larger in the direction of $\bf{u}$} than in a random direction (Figure 2e).
+Hence, low-dimensional dynamics emerged even in the presence of high-dimensional input.
+In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example.
+This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$).
+While the variability along $\bf{v}$} itself was not much smaller than a random directions (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f).
+Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
+Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g).
+The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in a later section, but note that it is closely aligned to $\bf{u}$} whenever $c$ is large.
+In other words, inputs in the direction of ${\bf{x}}_{\textrm{aligned}}$, which is close to $\bf{u}$}, are suppressed by the network.
+Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
+To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics.
+We next derive precise conditions for low-rank suppression and low-dimensional dynamics.
+General conditions for low-rank suppression and high-dimensional dynamics
+Above, we considered a simple network with rank-one structure.
+Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and derive precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
+Assumption 1: Dynamics are linear or near a stable equilibrium.
+We consider a general class of dynamics of the form
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},\bm{\eta})$ (4)
+for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics.
+Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and $\bm{\eta}(t)$ is an $N\times 1$ vector of external perturbations, inputs, and/or noise.
+In general, the dependence of ${\bf{F}}({\bf{z}},\bm{\eta})$ on $\bf{z}$} defines interactions between nodes, but it is not immediately clear how to define a fixed network of interactions, $W$, from the dynamics in Eq.
+(4).
+When dynamics are linear, $\bf{F}$} can be written as ${\bf{F}}({\bf{z}},\bm{\eta})=W{\bf{z}}+{\bf{G}}(\bm{\eta})$, and we can take $W$ to represent the effective connectivity.
+In nonlinear systems, the effective connectivity between nodes can change with the state of the system.
+To resolve this difficulty in defining network structure when dynamics are nonlinear, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
+Specifically, we assume that there exists a time-constant forcing, $\bm{\eta}(t)=\bm{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$.
+In other words, there exist $\bm{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
+${\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})={\bf{z}}_{0}.$
+We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
+$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$
+has eigenvalues with strictly negative real part.
+Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point.
+We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
+$\bm{\eta}_{p}(t)=\bm{\eta}_{0}+\epsilon\bm{\eta}(t).$
+The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\bm{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
+${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$ (5)
+The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (6)
+where
+$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ (7)
+is interpreted as the effective connectivity matrix and
+${\bf{x}}(t)=\partial_{\bm{\eta}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})\bm{\eta}(t)$
+is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
+In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$.
+In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq.
+(4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq.
+(6), so care must be taken in interpreting our results.
+For illustrative purposes, we consider a contrived example of such a model in Supplementary Figure S.2.
+Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix.
+For example, the system might be written as $d{\bf{z}}/dt={\bf{H}}(W{\bf{z}})+{\bf{x}}$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{H}}(W{\bf{z}}_{0})\neq W$.
+In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality.
+We consider one such example from epidemiology at the end of the Results.
+The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive.
+However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Supplementary Figures S.4 and S.5 for two examples of nonlinear, unstable dynamics).
+Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state.
+Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input).
+This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$.
+We discuss the relationships between their analysis and ours in more depth in Supplementary Text S.1.
+If a model is fully linear, i.e., if a model is already written in the form of Eq.
+(6), then we do not need to assume that dynamics remain close to the equilibrium, so inputs and perturbations need not be weak when dynamics are linear.
+In summary, we assume that the network dynamics are either fully linear or, if they are nonlinear, they represent a small perturbation away from a stable equilibrium.
+This assumption allows us to focus our analysis on the linearized dynamics in Eq.
+(6).
+Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
+We now formalize an asymptotic notion of a strongly low-rank network.
+Specifically, we consider the limit of large network size, $N\to\infty$, and assume that $W$ has $r\sim\mathcal{O}(1)$ asymptotically large singular values,
+$\sigma_{k}\gg 1,\;\;k=1,\ldots,r$
+where $\sigma_{k}$ denotes the $k$th largest singular value of $W$. We assume that the remaining singular values are small or moderate in magnitude,
+$\sigma_{j}\leq\mathcal{O}(1),\;\;j=r+1,\ldots,N.$
+Later, we consider some examples in which these scaling laws arise naturally.
+Matrices that satisfy the assumptions above can be decomposed as [1]
+$W=W_{0}+W_{1}$
+where $W_{0}$ has rank $r$ and asymptotically large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank.
+Since $W_{0}$ has rank $r$, we can write its singular value decomposition as
 $W_{0}=U\Sigma V^{T}$
-    be the singular value decomposition of $W$ with $U$ and $V$ being $N\times r$ orthonormal matrices. We define the $r\times r$ recurrent alignment matrix [3, 24],
-    $P=U^{T}V.$
-    which measures the alignment between the left and right singular vectors of $W_{0}$, i.e., the columns of $U$ and $V$.
-High-dimensional dynamics can only emerge when no singular value of $P$ is small, i.e., when $\sigma_{P}=\mathcal{O}(1)$ for all singular values of $P$.
-To better understand this condition on $P$, we consider simulations of a model that is widely used in computational neuroscience to model a recurrently interacting network of neurons [16, 25, 2, 26],
+where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values ($\Sigma_{kk}=\sigma_{k}\gg 1$) while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
+We additionally assume that $W_{1}$ is a random matrix that is independent of $U$ and $V$.
+In other words, $W$ is a low-rank matrix perturbed by noise.
+This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
+[…]
+Low-rank suppression occurs under Assumptions 1 and 2.
+Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically,
+Claim 1.
+Consider the dynamics defined by Eq.
+(6) under Assumptions 1 and 2.
+There exists an input ${\bf{x}}_{a}$ with $\|{\bf{x}}_{a}\|=\mathcal{O}(1)$ such that the network’s steady-state response to ${\bf{x}}_{a}$ is asymptotically larger than the $\mathcal{O}(1)$ response to a random input, ${\bf{x}}_{\textrm{r}}$ (where $\|{\bf{x}}_{\textrm{r}}\|=\mathcal{O}(1)$).
+To show why Claim 1 is true, first choose a $k\in\{1,\ldots,r\}$ and define
+${\bf{x}}_{a}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (8)
+Note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$.
+Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
+${\bf{z}}_{a}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
+This can be checked by substituting the equations for ${\bf{x}}_{a}$ and ${\bf{z}}_{a}$ into the steady-state equation, ${\bf{z}}=W{\bf{z}}+{\bf{x}}$.
+Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$.
+Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq.
+(8).
+In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is symmetric or normal.
+This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
+Note from Eq.
+(8) that ${\bf{x}}_{a}$ is closely aligned to ${\bf{u}}_{k}$ since $\sigma_{k}\gg 1$.
+Hence, the suppressed input, ${\bf{x}}_{a}$, is nearly (but not exactly) parallel to the left singular vector, ${\bf{u}}_{k}$.
+Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$.
+Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
+Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Supplementary Figure S.1 for which ${\bf{z}}=W{\bf{x}}$ in the steady-state).
+More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$).
+This is a left-right reversal and a suppression-amplification reversal of the behavior for recurrent networks, for which inputs aligned to ${\bf{u}}_{k}$ produce suppressed responses in the direction of ${\bf{v}}_{k}$.
+In summary, recurrent networks with strongly low-rank structure satisfying Assumptions 1 and 2 admit a small number of input directions along which inputs are strongly suppressed.
+We call this form of suppression low-rank suppression.
+In later sections, we will explore consequences of low-rank suppression in common low-rank network structures.
+Before discussing the dimensionality of dynamics, we first need to make an additional assumption about the dimensionality of the input.
+Assumption 3: External inputs are high-dimensional and slowly varying.
+To begin discussing the dimensionality of ${\bf{z}}(t)$, we need a model of ${\bf{x}}(t)$ that allows us to quantify dimensionality in a concrete way.
+To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process in which case ${\bf{z}}(t)$, as defined in Eq.
+(6), is also a stationary ergodic process whenever dynamics are stable [31].
+This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
+We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Supplementary Text S.2 for more mathematical details).
+The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics.
+When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Supplementary Figure S.2 for an example).
+We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions.
+To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$.
+We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
+${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
+In Supplementary Text S.2, we give a more mathematically detailed definition of this assumption.
+In Supplementary Text S.3, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
+In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d.
+Gaussian process with correlation time $\tau_{x}=10$.
+Our assumption of high-dimensionality is satisfied because each ${\bf{x}}_{j}(t)$ is independent.
+Our assumption that ${\bf{x}}(t)$ varies slowly is satisfied because $\tau_{x}$ is much larger than the network’s intrinsic timescale, $\tau=1$.
+Conditions for low-dimensional dynamics.
+In Supplementary Text S.2, we prove that (under Assumptions 1, 2, and 3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
+$P=V^{T}U$
+This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$.
+Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$.
+This matrix is also sometimes referred to as the “overlap matrix” [11].
+Our main result from Supplementary Text S.2 can be stated as follows:
+Claim 2.
+Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
+Here, “low-dimensional dynamics” is interpreted to mean that the covariance matrix of ${\bf{z}}(t)$ has an $\mathcal{O}(1)$ number of dominant eigenvalues, equivalently, ${\bf{z}}(t)$ has a small number of dominant principal component directions.
+Note that this result is equivalent to the following:
+Claim 2.
+Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If all singular values of $P=V^{T}U$ are $\mathcal{O}(1)$ then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
+Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
+Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure.
+a) Singular values of $W$ demonstrate an effective low-rank structure.
+b) The alignment matrix, $P$, when $W_{0}$ is normal.
+c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics.
+d,e) Same as b,c except $W_{0}$ is EP, but non-normal.
+f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP.
+h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq.
+(10) with $c=0.5$).
+To better understand the conditions on $P$ in Claim 2, we consider simulations of a widely used model from computational neuroscience [27, 32, 2, 33],
 $\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (9)
@@
 while exploring different choices of $V$.
-We first consider the special case that $W_{0}$ is normal, so ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$.
-In this case, $P$ is diagonal with $|P_{kk}|=1$ (Figure 2b) so all singular values of $P$ are $\sigma_{P}=1$ and the network exhibits high-dimensional dynamics (Figure 2c).
-Since $W_{0}$ is normal, its eigenvalues are real and equal in magnitude to the singular values, $|\lambda_{W_{0}}|=\sigma_{W_{0}}$.
-Since $\sigma_{W_{0}}\gg 1$, stability of the dynamics requires that $\lambda_{W_{0}}=-\sigma_{W_{0}}$ and therefore
-$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&-{\bf{u}}_{2}\end{array}\right].$
-Dynamics on the column space of $U$ are strongly suppressed due to the large, negative eigenvalues.
-Figure 2: Conditions for high-dimensional dynamics in a network with rank-two structure.
-a) Singular values of $W$ demonstrate an effective low-rank structure.
-b) The alignment matrix, $P$, when $W_{0}$ is normal.
-c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics.
-d,e) Same as b,c except $W_{0}$ is EP, but not normal.
-f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP.
-h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq.
-(4) with $c=0.5$).
-However, low-rank suppression does not require that $W_{0}$ is normal (${\bf{v}}_{k}=\pm{\bf{u}}_{k}$).
+We first consider a case in which $W_{0}$ is normal by setting
+$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&-{\bf{u}}_{2}\end{array}\right].$
+This is an extension of the example from Figure 1 to rank $r=2$.
+In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c).
+In this example, $W_{0}$ is both normal and symmetric.
+Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
+However, high-dimensional dynamics do not require that $W_{0}$ is normal.
 Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$.
@@
 All normal matrices are EP, but an EP matrix is not necessarily normal.
-An example of a non-normal, EP matrix with rank 2 is given by
+If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied.
+An example of a non-normal, EP matrix $W_{0}$ with rank 2 is given by taking
 $V$ $=\left[\begin{array}[]{cc}{\bf{u}}_{2}&-{\bf{u}}_{1}\end{array}\right].$
 The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 3d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$.
-Regardless, all singular values of $P$ are $\sigma_{P}=1$, and therefore the network produces low-rank suppression and high-dimensional dynamics (Figure 2e).
-To obtain an example of a network that produces low-dimensional dynamics, we take
+Regardless, all singular values of $P$ are $\sigma_{P}=1$, so the network produces high-dimensional dynamics (Figure 3e).
+To obtain an example of a network that produces low-dimensional dynamics (see also the example in Figure 2), we set
 $V$ $=\left[\begin{array}[]{cc}{\bf{v}}_{\perp}&-{\bf{u}}_{2}\end{array}\right].$
-where ${\bf{v}}_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$.
-In this case, $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 2f).
-Because of the singular value at zero, the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 2g).
-Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap, for example
+where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 3f).
+Because of the singular value at zero, the condition in Claim 3 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3g).
+Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
 $V$ $=\left[\begin{array}[]{cc}\sqrt{1-c}{\bf{v}}_{\perp}-\sqrt{c}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
-where $10<c<$ 1.
-In this case, $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and another at $\sigma_{P}=1$ (Figure 2h).
-As long as $c$ is not asymptotically close to zero, the network produces low-rank suppression and high-dimensional dynamics (Figure 2i).
-In summary, the class of effectively low-rank matrices, $W$, that produce low-dimensional dynamics in the presence of high-dimensional inputs is the very specific class of matrices for which the alignment matrix, $P=U^{T}V$, of the low-rank part is singular or nearly singular in the sense that it has a small or zero singular value.
+where $0<c<1$.
+In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3h).
+As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3i).
 It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean.
-In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.2).
-However, many networks arising in nature and applications do not have purely random structure.
+In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.3; compare to Figure 2).
+However, many networks arising in nature do not have purely random structure.
 In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
 Biased weights and modular networks
-So far, we considered example networks unbiased weights, $E[W_{jk}]=0$, which is common in many modeling studies, but many networks in nature have weights with non-zero mean.
+So far, we considered networks with unbiased weights, $E[W_{jk}]=0$, which is common in modeling studies, but many networks in nature have weights with non-zero mean.
 Biased weights can produce low-rank structure.
@@
 $E[W_{jk}]=m\;\textrm{ and }\;\textrm{std}(W_{jk})=s.$
-If $m$ and $s$ scale similarly to each other, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}\sim\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure because $\sigma_{1}/\sigma_{2}\sim\mathcal{O}(\sqrt{N})$.
-The dominant rank-one part of such a matrix has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.3).
+If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large.
+The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Supplementary Figure S.6; see Appendix B in [35] for related results).
 Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network.
@@
 $W=\left[\begin{array}[]{ccc}W^{1,1}&\cdots&W^{1,n}\\ W^{2,1}&\cdots&W^{2,n}\\ \vdots&\cdots&\vdots\\ W^{n,1}&\cdots&W^{n,n}\end{array}\right]$
-where $W^{a,b}$ is a sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
-$E[W_{jk}]=m_{ab}\;\textrm{ and }\;\textrm{std}(W_{jk})=s_{ab}.$
-If we assume that each population has $\mathcal{O}(N)$ members and that $m_{ab}$ scale similarly to $s_{ab}$ for large $N$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(\sigma_{ab}\sqrt{N})$.
-Hence, large modular networks with biased weights naturally produce low-rank structure [1].
-Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is the element-wise expectation of $W_{0}$, which is constant within each block, so $W_{0}$ has rank at most $n$.
-In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the $n$ indicator vectors of the $n$ populations.
-Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics in response to high-dimensional perturbations.
-As a specific example of a modular network, we consider a model of a local neuronal network in the cerebral cortex.
-Cortical neurons obey Dale’s Law: All outgoing connection weights from a particular neuron have the same sign, positive for excitatory neurons and negative for inhibitory neurons, and mean connection weights also depend on the postsynaptic neuron type [29, 30].
-These properties produce a modular structure in which the columns of the adjacency matrix corresponding to excitatory neurons are non-negative, while the columns corresponding to inhibitory neurons are non-positive (Figure 3a).
-We consider the case in which $m_{ab}$ and $s_{ab}$ scale like $1/\sqrt{N}$, consistent with experiments [31] and theoretical work [32, 33, 34, 35, 3, 36].
-This network has an effective rank-two structure (Figure 3b): Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$ and the low-rank part of the connectivity matrix is EP (Figure 3c), so our theory predicts that the network exhibits low-rank suppression and high-dimensional responses to high-dimensional inputs.
+where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
+$E\left[W^{a,b}_{jk}\right]=m_{ab}\;\textrm{ and }\;\textrm{std}\left(W^{a,b}_{jk}\right)=s_{ab}.$
+If we assume that each population has $N_{a}\sim\mathcal{O}(N)$ members and that the $m_{ab}$ scale similarly to the $s_{ab}$, then $W$ has up to $n$ dominant singular values that scale like $\mathcal{O}(m_{ab}N)$ and the remaining singular values scale like $\mathcal{O}(s_{ab}\sqrt{N})$.
+Hence, when $s_{ab}\sim m_{ab}$, large modular networks with biased weights naturally produce a low-rank structure in the sense that a small number of singular values are asymptotically larger than the rest [1].
+If additionally $m_{ab}\gg\mathcal{O}(1/N)$, then the dominant singular values are asymptotically large, so the network is strongly low-rank in the sense defined in Assumption 2.
+Networks of this form can be decomposed as $W=W_{0}+W_{1}$ where $W_{0}=E[W]$ is constant within each block, so $W_{0}$ has rank at most $n$.
+In general, $W_{0}$ is not a normal matrix, but it is an EP matrix because $\textrm{col}(U)$ and $\textrm{col}(V)$ are each spanned by the indicator vectors of the $n$ populations,
+$\textrm{col}(U)=\textrm{col}(V)=\textrm{span}\left\{{\bf{1}}_{1},{\bf{1}}_{2},\ldots,{\bf{1}}_{n}\right\}$
+where each ${\bf{1}}_{k}$ is an $N$-dimensional indicator vector with entries defined by
+${\bf{1}}_{k}]_{j}=\begin{cases}1&\textrm{if index $j$ is in population $k$}\\ 0&\textrm{otherwise}\end{cases}$
+Therefore, modular networks with biased weights exhibit low-rank suppression and high-dimensional dynamics.
+[…]
+Together, these biologically justified assumptions imply that the network has a strongly low rank structure, as defined in Assumption 2, with rank $r=2$ (Figure 4b).
+Specifically, Two singular values scale like $\sqrt{N}$ while the others are $\mathcal{O}(1)$.
+Moreover, the low-rank part of the connectivity matrix is EP (Figure 4c).
+Hence, the network produces high-dimensional dynamics and low-rank suppression despite the fact that $W$ is low-dimensional and non-normal.
 The column space of $U$ and $V$ consist of all vectors that are uniform within each population,
-$\textrm{col}(U)=\textrm{col}(V)=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}.$
-Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 3d,e) consistent with recordings from monkey motor cortex [36] and previous theoretical work on balanced network models [37, 38, 39, 36].
-As predicted, simulations also show high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.4), consistent with observations that neural responses in primate visual cortex are high-dimensional when visual stimuli are high-dimensional [13].
-Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)$.
-As a result, the cancellation mechanism illustrated in Figure 1g manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 3f), a phenomenon that is widely observed in neural recordings [6, 7, 8, 9, 10, 11, 12] and widely studied in computational models [32, 33, 39, 3].
-Therefore, the widely studied phenomenon of excitatory-inhibitory balance emerges as a consequence of low-rank suppression.
+$\textrm{col}(U)=\textrm{col}(V)$ $=\textrm{span}\{{\bf{1}}_{e},{\bf{1}}_{i}\}$
+$=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}$
+where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$.
+Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e).
+In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24].
+As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Supplementary Figure S.7), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
+Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$.
+As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [44, 45, 46, 47, 48, 49, 50] and widely studied in computational models [21, 40, 24, 3].
+Hence, the widely studied theory of balanced network models can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to a particular modular network structure.
 Amplification of disordered perturbations in networks with spatial structure
@@
 d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
-Many networks in nature exhibit connection strength that depends on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 34].
-As a simple example, we considered a model in which each node is assigned a location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are more strongly connected (Figure 4a).
-Connectivity was also perturbed by a random component, $W_{1}$, as above.
-This connectivity structure is effectively low-rank with the first several singular values proportional to the Fourier coefficients of the Gaussian connectivity kernel while the remaining singular values, inherited from $W_{1}$, are much smaller (Figure 4b).
-The low-rank part is normal, with left- and right singular vectors equal to the Fourier basis vectors (Supplementary Figure S.5), so the network satisfies our conditions for low-rank suppression and high-dimensional responses.
-Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low frequency spatial Fourier modes.
-Therefore, due to low-rank suppression, network responses to spatially disordered perturbations are amplified relative to smooth perturbations (Figure 4c-e; compare tick labels on vertical axes).
-The network also exhibits high-dimensional responses to high-dimensional perturbations (Supplementary Figure S.6), as predicted.
+Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 41].
+As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a).
+Connectivity is also perturbed by a random component, $W_{1}$, as above.
+This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Supplementary Figure S.8).
+Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Supplementary Figure S.9).
+Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes.
+Hence, perhaps surprisingly, strongly low-rank networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
 Low-rank suppression and high-dimensional dynamics in an epidemiological network.
+We next consider a real epidemiological network, specifically a network of high school social contacts [51], which was used in recent theoretical work on low-rank network dynamics [1].
+In that work, the authors considered a quenched mean-field reduction of the susceptible-infected-susceptible model,
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (11)
+where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication.
+Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected.
+Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51].
+This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a).
+In [1], it was shown that the dynamics generated by Eq.
+(11) on this network are effectively low-dimensional.
 Figure 6: Low-rank suppression and high-dimensional dynamics in a real epidemiological network.
@@
 b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
-We next consider a real epidemiological network, specifically a network of high school social contacts [40], which was used in a recent theoretical study of low-rank networks [1].
-In that study, the authors considered quenched mean-field reduction of the susceptible-infected-susceptible epidemiological model,
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (5)
-where $g>0$, $\gamma>0$ and $\circ$ denotes element-wise multiplication.
-In this model, each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected.
-Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [40].
-This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a).
-In [1], it was shown that the dynamics, ${\bf{z}}(t)$, generated by Eq.
-(5) on this network are effectively low-dimensional.
 However, Eq.
-(5) is completely self contained without any internal noise or external perturbations.
-Noise and perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the high school) and through the natural stochasticity of infection.
-Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network will produce low-rank suppression and high-dimensional responses to high-dimensional perturbations.
+(11) is completely self-contained without any perturbations.
+Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the school) and through the natural stochasticity of infection.
+Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network produces low-rank suppression and high-dimensional responses to high-dimensional perturbations.
 To test this prediction, we modified the model by adding an external forcing term,
 $\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
-Simulations of these dynamics indeed demonstrate low-rank suppression (Figure 5a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 5b) in contrast to the model from [1].
+Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1].
 These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics.
@@
 We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices.
-Moreover, in the presence of high-dimensional noise or perturbations, low-rank networks often produce high-dimensional dynamics.
-We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional responses to high-dimensional perturbations.
-The fact that stable, strongly low-rank networks produce low-rank suppression might seem obvious in hindsight: Whenever $W_{0}$ is normal, stability implies that $W$ has a large, negative eigenvalues, so dynamics are highly suppressive in the direction of the corresponding eigenvectors (as in Figures 1 and 2b,c).
-However, when $W_{0}$ is highly non-normal $W_{0}$, the network still produces low-rank suppression even when eigenvalues are small in magnitude (Supplementary Figure S.2), so this argument alone does not explain low-rank suppression.
-The fact that low-rank networks produce high-dimensional responses to high-dimensional inputs might also seem obvious in hindsight: Inputs drive the network, so high-dimensional input should drive high-dimensional activity regardless the internal network structure.
-However, our condition on the alignment of singular vectors shows that this explanation is not sufficient.
-When left- and right-singular vectors are not aligned ($\sigma_{P}\ll 1$), high-dimensional inputs drive low-dimensional responses (Figure 2f,g and Supplementary Figure S.2).
+We also derived conditions under which low-rank networks generate high-dimensional dynamics.
+We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
+Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics.
+In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Supplementary Text S.1 for a more complete comparison).
+Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
+Low-rank suppression might seem obvious in hindsight: When $W_{0}$ is normal and symmetric, the combination of stability and large singular values implies eigenvalues with large, negative real part, so dynamics are highly suppressive in one direction.
+Similarly, high-dimensional dynamics might seem like an obvious consequence of high-dimensional inputs.
+However, the example in Figure 2 coupled with the results in Claims 1 and 2 show that reality is more nuanced: Large negative eigenvalues are not necessary for low-rank suppression, and high-dimensional input is not sufficient for high-dimensional responses.
 Our results have implications for networks in nature and applications.
-For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 17, 4, 18, 19, 20, 3, 5, 24].
-Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural dynamics [41, 42, 43].
-However, many of these recordings are made during the presentation of low-dimensional stimuli or motor tasks.
-A growing number of more recent studies have shown that neural activity is high-dimensional during the presentation of high-dimensional stimuli or tasks [13, 43, 44, 45, 46, 47].
-Our results are consistent with the hypothesis that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [13].
-Our results are also consistent with observations that neural populations are more sensitive to random perturbations than to perturbations that are aligned with low-dimensional network structure [36].
-Beyond neuroscience, our results imply that perturbations or inputs to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space.
-This result can be used to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
+For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11].
+In neuroscience, our results are consistent with the observation that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [16].
+Beyond neuroscience, our results imply that perturbations to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space.
+More generally, perturbations to a low-rank network are more effective when they are not aligned to the network’s low rank structure.
+This result can be leveraged to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
 Methods
 All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code.
-All differential equation simulations – except for Figure 5 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
+All differential equations – except for Figure 6 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
 The simulations in Figure 6 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1].
@@
 Acknowledgments
-We thank Ashok Litwin-Kumar for helpful conversations and comments on drafts of the manuscript.
+We thank Ashok Litwin-Kumar and David G.
+Clark for helpful conversations and comments on drafts of the manuscript.
 This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
@@
 Figure S.1: Response properties of a feedforward network with rank-one structure.
-Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau{\bf{z}}^{\prime}=-{\bf{z}}^{\prime}+W{\bf{x}}$.
+Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$.
 Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
+Figure S.2: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure.
+Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$.
+The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a).
+The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
 Figure S.3: Response properties of a network with independent singular left- and right-singular vectors.
-Same as Figure 1 except $W_{0}=|c|{\bf{u}}{\bf{v}}^{T}$ where ${\bf{u}}$ and ${\bf{v}}$ are independent random unit vectors.
-Also, in panel c, we used a perturbation of ${\bf{x}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ as given in Eq.
-(15)) and in panel e-g, we projected onto ${\bf{v}}$.
-The network exhibits low-rank suppression in response to perturbations parallel to ${\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ (panel c) even though the connectivity matrix, $W$, does not have strongly negative eigenvalues (minimum real part of the eigenvalues of $W$ here is $\textrm{re}(\lambda)_{\textrm{min}}=-0.9$ compared to $\textrm{re}(\lambda)_{\textrm{min}}=-10$ in Figure 1).
-However, the network produces low-dimensional responses to high-dimensional inputs (panel d).
-Figure S.3: Response properties of a network with biased weights.
-Same as Figure 1 except we used the dynamics in Eq.
-(3), and $W_{jk}$ are drawn i.i.d.
-from a normal distribution with mean $-0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
-Figure S.4: Low-dimensional dynamics in a modular network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the modular network from Figure 3.
-Figure S.5: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 4.
-Figure S.6: Low-dimensional dynamics in a spatial network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the spatial network from Figure 4.
+Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors.
+The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
 Figure S.4: Dynamics of an unstable, chaotic network with rank-one structure.
-Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor ($c=-2$), and the dynamics obey Eq.
+Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor (c=2c=-2), and the dynamics obey Eq.
 (9).
@@
 The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
-Figure S.9: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure.
-Same as Figure 1d–g except ${\bf{x}}(t)$ was made low-dimensional by multiplying it by a matrix with rank-one structure, ${\bf{x}}\leftarrow W_{x}{\bf{x}}$ where $W_{x}$ is a matrix generated identically to, but independently from $W$.
-The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$.
+Figure S.6: Response properties of a network with biased weights.
+Same as Figure 1 except we used the dynamics in Eq.
+(9), and $W_{jk}$ are drawn i.i.d.
+from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
+Figure S.7: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
+Figure S.8: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
+Figure S.9: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
 Figure S.10: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.
 S.1 Comparison to previous work on dynamics of low-rank networks
-In this section we review the relationship between our models and analysis and the models and analyses from previous theoretical work on low-rank networks, specifically work by Ostojic et al.
+In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al.
 [2, 6, 4, 7, 8, 9], work by Thibeault et al.
 [1], and work by Landau and Sompolinsky [5, 3]
-All three studies quoted above consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model.
+All of these studies consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model.
 We next describe the salient properties that distinguish these models from ours.
@@
 Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2].
-Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) scale the same in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
+Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
 This difference in scaling is acknowledged by Ostojic et al.
-[2, 17, 4, 18, 19, 20] who refer to their networks as “weakly low-rank.” Keeping with this terminology, we will use “weakly low-rank” in this section to refer to networks in which the singular values of the low-rank components ($W_{0}$) and high-rank components ($W_{1}$) both scale like $\mathcal{O}(1)$.
-We additionally use “strongly low-rank” to refer to networks like ours in which the singular values of $W_{0}$ diverge to $\infty$ as $N\to\infty$ (typically, they scale like $\mathcal{O}(\sqrt{N})$) while the singular values of $W_{1}$ scale like $\mathcal{O}(1)$.
-To compare weakly and strongly low-rank networks, we can consider the simple rank-one example from Figure 1. Specifically,
-$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
-while $W_{1}$ has unbiased random entries with variance $\rho/\sqrt{N}$ and dynamics obey Eq.
-(1).
-Strongly low-rank networks have $c\gg 1$ while $c\sim\mathcal{O}(1)$ in weakly low-rank networks.
-In strongly low-rank networks, stability requires that $c<0$, which is critical for low-rank suppression.
-In weakly low-rank networks, however, it is possible to take $c>0$ without causing instability if $c<1$.
-When $c>0$, weakly low-rank networks can respond to perturbations aligned to ${\bf{u}}$ more strongly than random perturbations (Supplementary Figure S.11), reversing the trend of low-rank suppression.
-When $c<0$ in weakly low-rank networks, perturbations in the direction of ${\bf{u}}$ can be suppressed relative to random perturbations, but the effect is weaker than in strongly low-rank networks (Supplementary Figure S.12).
+[2, 6, 4, 7, 8, 9] who refer to their networks as “weakly low-rank.”
+Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
+Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
+The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent.
+For example, recent work by Mastrogiuseppe et al.
+[10] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values.
+While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [2], the absence of a random part means that their connectivity matrices are truly low-rank.
+In that work [10], Mastrogiuseppe et al.
+showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
+For the sake of comparison to the models in [10], let us consider a rank-one network structure like the one in Figure 1.
+Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
+$W_{0}=c{\bf{u}}{\bf{u}}^{T}.$
+In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$.
+In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$.
+Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2 and Supplementary Figure S.3 which are strongly low-rank without large eigenvalues).
+We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]).
+In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Supplementary Figure S.11 and S.12).
+When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Supplementary Figure S.11).
+Similar examples and results are considered in [10].
+In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
+Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$).
+When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Supplementary Figure S.12), reversing the trend of low-rank suppression.
+However, the amplification is weak.
+In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Supplementary Figure S.12d; compare to Figure 2 and Supplementary Figure S.3).
 Landau and Sompolinsky [5, 3] consider strongly low-rank networks.
@@
 are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
-Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except $c=0.75$.
-Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=-0.75$.
-S.2 Mathematical analysis of low-rank networks with external perturbations
-In this section we present a mathematical analysis of low-rank suppression and high-dimensional dynamics in networks with effective low-rank connectivity and external perturbations.
-Our analysis applies to a broad class of dynamics.
-We begin by reviewing the precise assumptions made in our analysis.
-S.2.1 Assumptions on the dynamics and definition of the effective connectivity matrix.
-We start by assuming dynamics of the form
-$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},{\bf{x}})$ (7)
-for some smooth ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$ where $\tau>0$ sets the timescale of the dynamics.
-Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and ${\bf{x}}(t)$ is an $N\times 1$ vector of external forces, inputs, and/or noise.
-At the end of Section S.2.3, we show how our approach can be generalized to a larger class of causal, time invariant systems, but the dynamics in Eq.
-(7) are sufficient for all of the examples studied in this manuscript, and the majority of models studied in the literature on low-rank networks, so we focus on this class of dynamics for simplicity.
-In general, it is not immediately clear how to define a network of interactions, $W$, from dynamics as general as those in Eq.
-(7).
-For example, in nonlinear systems, the effective connectivity between nodes can change with the state of the system.
-To circumvent this problem, we define the effective connectivity matrix, $W$, through a linearization around a stable equilibrium.
-Specifically, we assume that there exists a time-constant forcing, ${\bf{x}}(t)={\bf{x}}_{0}$, that produces a stable steady state solution, ${\bf{z}}_{0}$. In other words,
-${\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})={\bf{0}}.$
-We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
-$J_{{\bf{z}}}=\partial_{{\bf{z}}}F({\bf{z}}_{0},{\bf{x}}_{0})$
-has eigenvalues with strictly negative real part. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
-${\bf{x}}_{p}(t)={\bf{x}}_{0}+\epsilon{\bf{x}}(t).$
-The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, ${\bf{x}}_{p}(t)$, can be written to linear order in $\epsilon$ as
-${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}+\mathcal{O}(\epsilon^{2}).$ (8)
-The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}(t)$ (9)
-where $J_{{\bf{x}}}=\partial_{{\bf{x}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ is the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{x}}$ evaluated at the fixed point, and
-$W=J_{{\bf{z}}}+I$ (10)
-with $I$ the identity matrix and $J_{{\bf{z}}}=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{z}}$ evaluated at the fixed point.
-We interpret $W$ in Eq.
-(10) as the effective connectivity matrix.
-It may seem counterintuitive at first to define $W=J_{{\bf{z}}}+I$ instead of $W=J_{{\bf{z}}}$.
-To understand our choice, note that when we choose $W=J_{{\bf{z}}}+I$, the perturbed system ${\bf{z}}_{p}(t)$ decays to the unperturbed fixed point, ${\bf{z}}_{0}(t)$, whenever there is no perturbation or connectivity (${\bf{x}}(t)=0$ and $W=0$).
-Moreover, when there is a perturbation without connectivity ($W=0$ and ${\bf{x}}\neq 0$), the perturbation converges to ${\bf{z}}=J_{{\bf{x}}}{\bf{x}}$.
-Hence, $W$ represents interactions that determine the dynamics under perturbations away from the fixed point.
-In any case, our overall results are not sensitive to this choice.
-One caveat to our approach for defining effective connectivity, $W$, from the dynamics in Eq.
-(7) is that the input perturbation, ${\bf{x}}(t)$, must be weak enough and/or the dynamics must be sufficiently linear for the linearized dynamics to be accurate, i.e., for the $\mathcal{O}(\epsilon^{2})$ term in Eq.
-(8) to be small.
-Moreover, the dynamics of the unperturbed system must be stable near the steady state.
-Examples with unstable, nonlinear dynamics are considered in Supplementary Figures S.7 and S.8.
-Another caveat is that multiplication of ${\bf{x}}$ by $J_{{\bf{x}}}$ in Eq.
-(9) could suppress dimensionality, for example if $J_{{\bf{x}}}$ is low-rank or effectively low-rank.
-This would correspond to an effectively feedforward “read-in” mechanism that lowers dimensionality.
-Specifically, the “effective input” $J_{{\bf{x}}}{\bf{x}}$ would be low-dimensional and suppressed in some directions.
-For example, consider dynamics of the form
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+W_{x}{\bf{x}}$ (11)
-where $W_{x}$ is some matrix.
-In this model, we have $J_{{\bf{x}}}=W_{x}$.
-Therefore, if $W_{x}$ were low-rank then the effective input, ${\bf{x}}_{eff}=W_{x}{\bf{x}}$, would be low-dimensional even when the raw input, ${\bf{x}}$, were high-dimensional.
-This scenario would promote low-dimensional dynamics in ${\bf{z}}(t)$.
-An example of this form is considered in Supplementary Figure S.9.
-In many models (including Eqs.
-(1) and (3)), the dynamics take the form ${\bf{F}}({\bf{z}},{\bf{x}})={\bf{F}}({\bf{z}})+{\bf{x}}$ so that $J_{{\bf{x}}}=I$ and this is not an issue.
-More generally, we assume that $J_{{\bf{x}}}$ is full rank in all of our analysis.
-Before continuing, we compare our approach for relating dynamics to connectivity with the approach taken in Thibeault et al.
-[1].
-They define dynamics that explicitly take the connectivity matrix into account,
-$\frac{d{\bf{z}}}{dt}$ $={\bf{G}}({\bf{z}},{\bf{y}})$
-${\bf{y}}$ $=U{\bf{z}}$
-where $U$ is interpreted as the connectivity matrix (they use “$W$”, but we switched to “$U$” to avoid confusion with our definition).
-This formulation represents a self-contained network in the absence of external perturbations or noise, but it can easily be extended to allow external perturbations by defining
-$\frac{d{\bf{z}}}{dt}={\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}}).$ (12)
-However, Eq.
-(12) is not suitable for our purposes because $U$ does not necessarily capture all of the interactions between nodes.
-For example, the $j$th entry of ${\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}})$ could depend explicitly (through its first argument) on the $k$th entry of ${\bf{z}}$ for some $j\neq k$.
-This would represent an interaction between nodes $j$ and $k$, which are not accounted for in $U$.
-For example, we could have $d{\bf{z}}_{j}/dt={{\bf{G}}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{z}}_{k}+{\bf{y}}_{j}={\bf{z}}_{k}+U_{ji}{\bf{z}}_{i}$ for three distinct indices $i,j,k$ where we assume that $U_{jk}=0$.
-In this case, the connectivity matrix, $U$, correctly captures the influence of node ${\bf{z}}_{i}$ on node ${\bf{z}}_{j}$ through the entry $U_{ji}$.
-However, node ${\bf{z}}_{j}$ is also influenced by node ${\bf{z}}_{k}$ and this interaction is not captured by the matrix $U$ (as we take $U_{jk}=0$) because it is instead captured by the explicit dependence of ${\bf{G}}$ on ${\bf{z}}_{k}$ (the first argument of ${\bf{G}}$).
-This could, of course, be overcome by incorporating the dependence on ${\bf{z}}_{k}$ into $U$ (i.e., by setting $U_{jk}=1$), but Eq.
-(12) makes no assumption that all such interactions are included into $U$.
-More importantly, it might not be possible to account for higher order, nonlinear interactions using the formulation in Eq.
-(12).
-For example, we could have $d{\bf{z}}_{j}/dt={\bf{G}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{y}}_{k}{\bf{z}}_{i}=U_{jk}{\bf{z}}_{k}{\bf{z}}_{i}$ for three distinct indices $i,j,k$.
-In this case, the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ would not be accounted for in the matrix $U$, and it is not possible to account for both interactions ($i\to j$ and $k\to j$) in the matrix $U$ using the formulation in Eq.
-(12).
-Moreover, the strength of the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ depends on the magnitude of ${\bf{z}}_{k}$, which can change with the state of the network.
-These matters are handled in the mathematical analysis of Thibeault et al.
-[1] by bounding norms related to the Jacobian matrix of ${\bf{G}}$ with respect to ${\bf{z}}$ and ${\bf{y}}$, and they also separately consider higher order interactions.
-Our approach avoids these bounds on the Jacobian matrices at the expense of relying on a linearization near a steady state.
-Under the formulation from Eq.
-(12), if there is a fixed point at ${\bf{z}}_{0}$ in response to static input ${\bf{x}}_{0}$, then our interpretation of the connectivity matrix is given by
-$W=\partial_{{\bf{z}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})+\partial_{{\bf{y}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})U+I$
-where ${\bf{y}}_{0}=U{\bf{z}}_{0}$ and the two partial derivatives denote Jacobian matrices with respect to the first and second arguments, respectively.
-Hence, the two approaches can easily be directly related.
-In summary, we consider the dynamics defined by Eq.
-(7), linearized around a stable steady state, which ultimately gives rise to the definition of network dynamics in Eq.
-(9).
-This approach has some caveats, which are discussed above.
-S.2.2 Assumptions on the network structure.
-The network represents $N$ interacting nodes, the interactions are defined by an $N\times N$ connectivity matrix, $W$, and we consider asymptotics in the $N\to\infty$ limit.
-Here, we specify the assumptions made about the structure of $W$ in the large $N$ limit.
-Letting $\sigma_{1},\ldots,\sigma_{N}$ be the singular values of $W$, we assume that $\sigma_{k}\gg 1$ for $k=1,\ldots,r$ (in the sense that $\sigma_{k}\to\infty$ as $N\to\infty$) and that $\sigma_{k}\leq\mathcal{O}(1)$ for $k=r+1,\ldots,N$ (in the sense that $\sigma_{k}$ has a finite limit as $N\to\infty$).
-We also assume that $r\sim\mathcal{O}(1)$ so that $r\ll N$.
-In summary, $W$ has a small number of large singular values and the remaining singular values are much smaller.
-Under these assumptions, $W$ can be decomposed as [1]
-$W=W_{0}+W_{1}$
-where $W_{0}$ has rank $r$ and we write its singular value decomposition as
-$W_{0}=U\Sigma V^{T}.$
-Here, $U$ and $V$ are $N\times r$ orthonormal matrices and $\Sigma$ is an $r\times r$ diagonal matrix with diagonal entries, $\Sigma_{kk}=\sigma_{k}$.
-The left and right singular matrices, $U$ and $V$, can be written as
-$U=\left[{\bf{u}}_{1}\;\;{\bf{u}}_{2}\;\cdots\;{\bf{u}}_{r}\right]$
-and
-$V=\left[{\bf{v}}_{1}\;\;{\bf{v}}_{2}\;\cdots\;{\bf{v}}_{r}\right]$
-where their columns, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$, are the left and right singular vectors, which satisfy $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$ and ${\bf{u}}_{k}\cdot{\bf{u}}_{j}={\bf{v}}_{k}\cdot{\bf{v}}_{j}=0$ for $j\neq k$.
-We further assume that $W_{1}$ is a random matrix, which is almost surely full rank and is statistically independent from $W_{0}$.
-Finally, we assume that the eigenvalues and singular values of $W_{1}$ are smaller than $1$ in magnitude and bounded away from $1$ ($\sigma_{W_{1}},|\lambda_{W_{1}}|<1-\delta$ for some fixed $\delta>0$) with probability $1$.
-S.2.3 Analysis of low-rank suppression.
-We now analyze the conditions under which low-rank suppression occurs.
-Our goal here is to understand the conditions under which the dynamics produce low-rank suppression like that observed in Figure 1c.
-In the next section, we will derive conditions on high-dimensional dynamics in the presence of stochastic perturbations such as those considered in Figure 1d–g.
-We begin by considering the case in which ${\bf{x}}(t)$ is static or slow, so we can use a steady state or quasi-steady state approximation.
-Later in this section, we consider low-rank suppression with transient and fast perturbations.
-If ${\bf{x}}(t)={\bf{x}}$ is static (such as during the duration of each stimulus in Figure 1c), then the network response converges to solutions of the perturbed steady state, which from Eq.
-(9) satisfies
-${\bf{z}}=W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}$
-and therefore
-${\bf{z}}=[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}.$ (13)
-If ${\bf{x}}(t)$ is not static, but changes much more slowly than $\tau$, then ${\bf{z}}(t)$ closely tracks the quasi-steady state given by Eq. (13).
-Low-rank suppression, like that observed in Figure 1c, occurs whenever there is an input, ${\bf{x}}$, that is strongly suppressed by the recurrent dynamics in comparison to a random input (as in Figure 1c).
-In other words, we can say that low-rank suppression occurs whenever there is a raw input, ${\bf{x}}$, such that $\|{\bf{x}}\|=\mathcal{O}(1)$ and
-$\|{\bf{z}}\|=\|[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}\|\ll\|{\bf{z}}_{rand}\|=\|[I-W]^{-1}{\bf{u}}_{rand}\|$
-with high probability whenever ${\bf{u}}_{rand}$ is a random unit vector, independent from $W_{0}$, $W_{1}$, and $J_{{\bf{x}}}$.
-To simplify this statement, first note that $\|[I-W]^{-1}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ whenever ${\bf{u}}_{rand}$ is independent from $W_{1}$.
-Moreover, since we assume that $J_{{\bf{x}}}$ is full rank, we can write the condition in terms of the effective input,
-${\bf{x}}_{eff}=J_{{\bf{x}}}{\bf{x}}$
-instead of the raw input, ${\bf{x}}$.
-In particular, low-rank suppression occurs whenever there is a vector ${\bf{x}}_{eff}$ such that
-$\|{\bf{z}}\|=\|[I-W]^{-1}{\bf{x}}_{eff}\|\ll\|{\bf{x}}_{eff}\|.$ (14)
-We take this property to be the definition of low-rank suppression.
-For the dynamics that we use in our simulations, note that $J_{{\bf{x}}}=I$ so ${\bf{x}}_{eff}={\bf{x}}$.
-We discuss some caveats of the distinction between ${\bf{x}}$ and ${\bf{x}}_{eff}$ at the end of this section.
-We now prove that low-rank suppression arises under the assumptions from Sections S.2.2S.2.3.
-Claim 1.
-Under the model and assumptions of Sections S.2.2S.2.3, the network exhibits low-rank suppression.
-Proof.
-Let $k\in\{1,\ldots,r\}$ and take
-${\bf{x}}_{eff}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (15)
-We claim that the response to this input is given by
-${\bf{z}}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
-To see this, we only need to show that $[I-W]{\bf{z}}={\bf{x}}$. We have
-${\bf{z}}$ $=\frac{W_{0}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
-$=\frac{U\Sigma V^{T}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
-$={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
-$={\bf{x}}_{eff}.$
-Now note that $\|{\bf{x}}_{eff}\|=1+o(1)$ and $\|{\bf{z}}\|\ll 1$ because $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, $\|W_{1}\|_{2}=\mathcal{O}(1)$, and $\sigma_{k}\gg 1$.
-Therefore $\|{\bf{z}}\|\ll\|{\bf{x}}_{eff}\|$.
-∎
-Before continuing, we will make a few observations about this claim and its proof. First note that the definition of low-rank suppression in Eq. (14) is equivalent to the statement that the matrix
-$A=[I-W]^{-1}$
-has at least one asymptotically small singular value,
-$\sigma_{A}\ll 1.$
-Singular values commute with matrix inverses, so the singular values of $A$ are given by
-$\sigma_{A}=\frac{1}{\sigma_{Q}}$
-where $\sigma_{Q}$ are the singular values of the matrix
-$Q=I-W.$
-Therefore, low-rank suppression occurs whenever $Q$ has at least one asymptotically large singular value,
-$\sigma_{Q}\gg 1.$
-Since $W$ has an asymptotically large singular value, low-rank suppression occurs because this large singular value is conserved by adding the identity matrix.
-However, unlike eigenvalues, singular values are not transformed in a simple way by adding the identity matrix, in other words, we do not generally have $\sigma_{Q}=1-\sigma_{W}$.
-If $W$ were a normal matrix, then we would have $\sigma_{Q}=|1-\sigma_{W}|$, so the emergence of low-rank suppression would have been easier to prove, but this is not in general true of non-normal matrices.
-The potentially complicated relationship between the singular values of $W$ and those of $Q$ is more critical for the analysis of high-dimensional dynamics in the next section.
-The proof of Claim 1 also provides insight into the input and response directions that are suppressed in low-rank suppression.
-Note that ${\bf{x}}_{eff}={\bf{u}}_{k}+o(1)$ and ${\bf{z}}\propto{\bf{v}}_{k}$, which implies that inputs approximately aligned to the column space of $U$ (equivalently, the column space of $W_{0}$) give rise to suppressed responses aligned to the column space of $V$ (equivalently, the row space of $W_{0}$).
-Finally, it is worth discussing our use of ${\bf{x}}_{eff}$ instead of ${\bf{x}}$ in the definition of low-rank suppression.
-For our simulations ${\bf{x}}={\bf{x}}_{eff}$ so there is no issue.
-However, one could consider a model for which $J_{{\bf{x}}}$ is full rank, but amplifies some directions more than others.
-For example, consider the model in Eq.
-(11) with a $W_{x}$ chosen to amplify inputs in every direction parallel to the column space of $W$, i.e.., if $\|W_{x}{\bf{u}}_{k}\|\gg 1$ for $k=1,\ldots,r$ whereas $\|W_{x}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ for random directions.
-In this very specific case, the amplification by $W_{x}$ would compete with the suppression by $A=[I-W]^{-1}$, so the raw input, ${\bf{x}}$, would need to be smaller in the direction of ${\bf{u}}_{k}$ to obtain an ${\bf{x}}_{eff}$ that is $\mathcal{O}(1)$ in that direction.
-In this case, the suppression by recurrent dynamics would compete with the amplification by a feedforward read-in.
-Importantly, the recurrent dynamics would still be suppressive to effective inputs aligned to ${\bf{u}}_{k}$.
-We focus on effective inputs in our definition of low-rank suppression to avoid such an issue, but it would only arise in a very specific situation where $J_{{\bf{x}}}$ amplifies exactly the directions that $A=[I-W]^{-1}$ suppresses.
-Above, we considered slow or static perturbations, ${\bf{x}}(t)$.
-Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models).
-This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics.
-We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales.
-We can no longer rely on the quasi-static approximation from Eq.
-(13) in this case.
-Instead, we take the Laplace transform in Eq.
-(9) to obtain
-$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (16)
-where
-$\hat{H}(s)$ $=(1+\tau s)^{-1}\left[J_{{\bf{z}}}+I\right]=(1+\tau s)^{-1}W$ (17)
-$\hat{B}(s)$ $=(1+\tau s)^{-1}J_{{\bf{x}}}$
-[…]
-The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$, i.e.,
-$W=\hat{H}(s)$
-because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$.
-Solving the implicit Eq.
-(16) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
-$\hat{{\bf{z}}}=[I-W]^{-1}\hat{B}\hat{{\bf{x}}}.$ (18)
-Since $\hat{B}$ quantifies feedforward, read-in dynamics, we can interpret $\hat{B}\hat{{\bf{x}}}$ as effective input in the Laplace domain, ${\bf{\hat{x}}}_{eff}=\hat{B}\hat{{\bf{x}}}$.
-This is analogous to our treatment of $J_{{\bf{x}}}{\bf{x}}$ above.
-Eq.
-(18) is identical to Eq.
-(13) except that Eq.
-(18) quantifies interactions at individual Laplace modes, $s$.
-Hence, Claim 1 applies to situations in which ${\bf{x}}(t)$ is transient and fast, so long as the effective connectivity, $W$, and effective input, ${\bf{x}}_{eff}$ are interpreted in the Laplace domain ($W=H(s)$ and ${\bf{x}}_{eff}=\hat{B}(s)\hat{{\bf{x}}}(s)$).
-Moreover, the approach of working in the Laplace domain allows us to expand our class of models.
-In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq.
-(7).
-The dynamics in Eq.
-(7) capture a large class of models, but are not fully general.
-Specifically, in Eq.
-(7), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$ and the values of $x(s)$ for $s\leq t_{1}$.
-However, some systems like integro-differential equations and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq.
-(7).
-In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
-${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (19)
-In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivriant function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$.
-Eq.
-(19) encompasses an extremely broad class of systems, including those described by Eq.
-(7).
-We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq.
-(19).
-As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$.
-However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$.
-We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
-${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
-A first order Volterra expansion gives Eq.
-(16) as an implicit equation for ${\bf{z}}(t)$.
-The matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline.
-Under the dynamics in Eq.
-(9), we obtain the $\hat{H}$ and $\hat{B}$ defined in Eq.
-(17).
-Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq.
-(19).
-S.2.4 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
+S.2 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
 We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d).
-We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–g.
+We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–f.
 Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
-[…]
+$\widetilde{C}^{{\bf{x}}}_{jk}(f)=\int_{-\infty}^{\infty}C_{jk}^{{\bf{x}}}(\tau)e^{-2\pi if\tau}d\tau$
 where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance.
 When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$.
-This is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
+The assumption of i.i.d.
+elements is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
 The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (6) to get [31, 53, 41, 42, 54]
-$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}J_{{\bf{x}}}\widetilde{C}^{{\bf{x}}}(f)J_{{\bf{x}}}^{T}[\widetilde{r}(f)I-W]^{-*}$ (20)
+$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}$ (13)
 where
@@
 is scalar.
-Note that Eq.
-(20) can be derived from Eqs.
-(18) and (17) by first taking $s=-2\pi if$ in Eq.
-(18) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$.
-Hence, the results derived in this section apply to the general class of dynamics defined in Eq.
-(19).
 Eq. (13) quantifies the covariance structure of ${\bf{z}}(t)$ at any given frequency mode, $f$, but we are often specifically interested in the zero-lag temporal covariance,
-[…]
+$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$
 From Eq. (13), we therefore have
-[…]
+$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}df$ (14)
 If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$.
-Indeed, we take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$.
+Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$.
 Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero.
@@
 In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
-$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}J_{{\bf{x}}}\overline{C}^{{\bf{x}}}J_{{\bf{x}}}^{T}[I-W]^{-T}.$ (22)
+$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}\overline{C}^{{\bf{x}}}[I-W]^{-T}.$ (15)
+Note that Eq.
+(15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$.
+Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate, but our approach here is more mathematically precise.
 When principal component analysis is applied to ${\bf{z}}(t)$, the variance explained by each principal component is given by the ordered list of eigenvalues of the covariance matrix, $\overline{C}^{{\bf{z}}}$.
@@
 (15), we see that these eigenvalues are proportional to the squares of the singular values of the matrix
-$R=[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}.$
+$R=[I-W]^{-1}\sqrt{\overline{C}^{{\bf{x}}}}.$
 where $\sqrt{\overline{C}^{{\bf{x}}}}$ is the matrix square root of the symmetric positive definite covariance matrix, $\overline{C}^{{\bf{x}}}$ (distinct from the entry-wise square root in general).
@@
 stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$.
-Moreover, in the networks we consider, $J_{{\bf{x}}}=I$ is equal to the identity matrix.
 Therefore, for our models,
-$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}v$
-where $v$ is a scalar.
-In more general classes of models in which $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are not multiples of the identity matrix, dimensionality of the dynamics of ${\bf{z}}(t)$ could be reduced whenever $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are effectively low-rank.
-However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) or a low-rank read-in mechanism ($J_{{\bf{x}}}$ low-dimensional; see Supplementary Figure S.9 and discussion in previous sections), and would not be related to the actual dynamics of interactions within the network.
-Therefore, our focus on the effective rank of $[I-W]^{-1}$ in place of $[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}$ is justified even in more general settings.
-In conclusion, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
+$R=[I-W]^{-1}\sqrt{v}$
+and therefore
+$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}[I-W]^{-T}v$
+where recall that $v>0$ is a scalar.
+In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank.
+However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Supplementary Figure S.2), and would not be related to the intrinsic dynamics of interactions within the network.
+In conclusion, when ${\bf{x}}(t)$ is high-dimensional, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
 $A=[I-W]^{-1}.$
 More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
-[…]
-where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
+$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{A,k}\,v$
+where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values and principal components are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
 Therefore, ${\bf{z}}(t)$ is low-dimensional whenever $A$ has a small number of large singular values.
 As above, we can use the fact that singular values commute with matrix inverses to write our conclusions in terms of
-$Q=I-W.$
+Q=IW.Q=I-W.
 Specifically,
-[…]
-where $\sigma_{Q,N-k}$ is the $(N-k)$th singular value of $Q$, i.e., the $k$th from the last singular value (assuming singular values are sorted in decreasing order).
+$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{v}{\sigma^{2}_{Q,N-k}}$ (16)
+where $\sigma_{Q,N-k}$ is the (Nk)(N-k)th singular value of $Q$, i.e., the $k$th from the last singular value.
 Figure S.13: Comparison between theory and simulations for shorter and longer simulations.
@@
 (16) would be replaced by
-[…]
+$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{U,N-k}}$
 where
@@
 and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
-[…]
+$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{\hat{R},k}$ (17)
 where
 $\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
-is the sampled value of $R$ (assuming $J^{{\bf{x}}}=I$) and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used.
+is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used.
 Using Eq.
-(24) gives a much more accurate prediction (Supplementary Figure S.13a, green dots) This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$.
+(17) gives a much more accurate prediction (Supplementary Figure S.13a, green dots).
+This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$.
 We next increased the simulation time ten-fold to $T=5\times 10^{4}$.
 In this case, the original Eq.
-(23) was accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
-As concluded at the end of the previous section, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values).
+(16) was more accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
+As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values).
 Now, we may conclude that ${\bf{z}}(t)$ has low-dimensional dynamics whenever $Q$ has some asymptotically small singular values (equivalently, whenever $A$ has a small number of asymptotically large singular values).
-Conversely, high-dimensional dynamics occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
-Combining these conclusions with those reached in Section S.2.3, we can summarize as follows
-Low-rank suppression (as in Figure 1c) occurs whenever $Q=I-W$ has at least one asymptotically large singular value.
-High-dimensional dynamics (as in Figure 1d) occur whenever $Q=I-W$ lacks any asymptotically small singular values.
-In the previous section, we showed that $Q=I-W$ has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank.
-Conditions under which $Q=I-W$ lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple.
+Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
+Combining these conclusions with those in Claim 1, we can summarize as follows
+Low-rank suppression (as in Figure 1g) occurs whenever Q=IWQ=I-W has at least one asymptotically large singular value.
+High-dimensional dynamics (as in Figure 1d) can only occur whenever Q=IWQ=I-W lacks any asymptotically small singular values.
+Previously, in proving Claim 1, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank.
+Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple.
 Specifically, the lack or presence of small singular values depends on the recurrent alignment matrix,
@@
 which measures the alignment between the left and right singular vectors, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$.
-Specifically, $P_{jk}={\bf{v}}_{k}\cdot{\bf{u}}_{k}$ so that $P_{jk}=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$.
+Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ so that $|P_{jk}|=1$ whenever ${\bf{v}}_{k}=\pm{\bf{u}}_{k}$ and $P_{jk}=0$ whenever ${\bf{v}}_{k}$ is orthogonal to ${\bf{u}}_{k}$.
 Note that singular values of $P$ are bounded by unity, $\sigma_{P}\leq 1$.
@@
 Claim 2.
-Under the model and assumptions of Sections S.2.2 and S.2.4, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
+Under the model and assumptions 1–3 in the main text, if $P=V^{T}U$ does not have any asymptotically small singular values then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
 Proof.
@@
 If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
-The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d).
-From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever $Q=I-W$ has an asymptotically small singular value.
+The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2 and Figure 3g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d).
+From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value.
 Therefore, our original claim is equivalent to the following:
-If $Q=I-W$ has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
+If Q=IWQ=I-W has at least one asymptotically small singular value then $P=V^{T}U$ also has at least one asymptotically small singular value.
 We will prove this version of the claim directly. Assume that $Q$ has an asymptotically small singular value. Then there is a $\bf{z}$} satisfying $\|{\bf{z}}\|=1$ and
@@
 because $\Sigma^{-1}$ is a diagonal matrix with $o(1)$ terms on the diagonal. ∎
-The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2g.
+The proof of Claim 2 also tells us the dominant directions of variability whenever dynamics are low-dimensional as in Figure 2 and Figure 3g.
 Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$.
-Indeed, in Figure 2g, the angle between ${\bf{u}}_{2}$ and the dominant principle component direction of ${\bf{z}}$ is $18^{\circ}$.
 The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$.
@@
 This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).
+Next, in Supplementary Text S.3, we generalize away from the assumption of slow inputs using a Laplace transform.
+S.3 The analysis of fast and transient dynamics.
+Above, we considered slow or static perturbations, ${\bf{x}}(t)$.
+Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models).
+This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics.
+We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales.
+We can no longer rely on the quasi-steady state approximation from Eq.
+(2) in this case.
+Instead, we take the Laplace transform in Eq.
+(6) to obtain
+$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (18)
+where $\hat{\psi}(s)=\int\psi(t)e^{-st}dt$ is the Laplace transform of $\psi(t)$ and
+$\hat{H}(s)$ $=(1+\tau s)^{-1}W$ (19)
+$\hat{B}(s)$ $=(1+\tau s)^{-1}.$
+Solving the implicit Eq. (18) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
+$\hat{{\bf{z}}}=[I-\hat{H}]^{-1}\hat{B}\hat{{\bf{x}}}$. (20)
+The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$ because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$.
+At mode $s=0$ (the static or “DC” mode), we recover the connectivity from the steady-state or quasi-steady state analysis, $\hat{H}(0)=W$.
+Similarly, $\hat{B}(0)=1$.
+Hence, the quasi-steady state analysis from above evaluated how the DC component of the input, $\hat{{\bf{x}}}(0)$, affects the DC component of the dynamics, $\hat{{\bf{z}}}(0)$, even when we do not assume that ${\bf{x}}(t)$ changes slowly.
+The analysis here is a generalization of the quasi-steady state analysis to non-static modes, $s\neq 0$.
+Note also that Eq.
+(13) from Supplementary Text S.2 can be derived from Eqs.
+(20) and (19) above by first taking $s=-2\pi if$ in Eq.
+(20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$.
+Hence, the approach in this section is a generalization of the approach in Supplementary Text S.2.
+To interpret Eq.
+(20), let us consider two separate limits.
+At modes that are much slower than the intrinsic timescale of the dynamics, $|s|\ll 1/\tau$, we have that $\hat{H}(s)\approx W$ and $\hat{B}(s)\approx 1$, so
+$\hat{{\bf{z}}}(s)\approx[I-W]^{-1}\hat{{\bf{x}}}(s)$
+and we recover the results from our quasi-steady state analysis at these modes.
+In particular, Claim 1 holds and we expect low-rank suppression at mode $s$ whenever $W$ is strongly low rank.
+While this may at first seem like a repetition of our results from the quasi-steady state analysis, note that we do not need to assume that ${\bf{x}}(t)$ varies slowly here.
+The Laplace analysis at slow modes ($|s|\ll 1/\tau$) applies even when ${\bf{x}}(t)$ has power at slow and fast modes.
+At fast modes, $|s|\gg\tau$, the identity term in Eq. (20) dominates the $\hat{H}$ and $\hat{B}$ terms so that
+$\hat{{\bf{z}}}(s)\approx\hat{{\bf{x}}}(s)$
+at fast modes.
+Therefore, fast modes (representing transient responses to transient input dynamics) do not exhibit low-rank suppression, but they do produce high-dimensional responses to high-dimensional stimuli.
+Modes between these two extremes interpolate between them.
+In summary, when our previous assumption that ${\bf{x}}(t)$ varies more slowly than $\tau$ is violated, we expect to see low-rank suppression at slow temporal modes, but not at fast temporal modes.
+Additionally, working in the Laplace domain allows us to expand our class of models.
+In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq.
+(4), which we later linearized.
+The dynamics in Eq.
+(4) capture a large class of models, but are not fully general.
+Specifically, in Eq.
+(4), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$.
+However, some systems like integral equations, integro-differential equations, operator equations, and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq.
+(4).
+In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
+${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (21)
+In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivariant, causal function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$.
+Eq.
+(21) encompasses an extremely broad class of systems, including those described by Eq.
+(4).
+We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq.
+(21).
+As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$.
+However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$.
+We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
+${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
+A first order Volterra expansion gives Eq.
+(18) as an implicit equation for ${\bf{z}}(t)$, but the matrices $\hat{H}$ and $\hat{B}$ are no longer defined as above.
+Instead, the matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline.
+Under the dynamics in Eq.
+(6), we recover the $\hat{H}$ and $\hat{B}$ defined in Eq.
+(19).
+Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq.
+(21).

### v4

@@
 Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations.
-We find that dynamics in such networks can be high- or low-dimensional and we derive precise conditions on the network structure under which dynamics are high-dimensional.
+We find that dynamics in such networks can be high- or low-dimensional and we derive novel, precise conditions on the network structure under which dynamics are high-dimensional.
 In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression.
@@
 Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk.
-We use mathematical analysis to derive precise conditions under which linearized dynamics in these networks produce high-dimensional dynamics.
-Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.” We show that many common structural features such as biased weights, modularity, and spatial connectivity promote strongly low-rank structure, high-dimensional dynamics, and low-rank suppression.
+We show that such networks can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the detailed structure of the network connectivity.
+Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.”
+We use mathematical analysis to derive precise conditions under which networks with low-rank structure produce high-dimensional dynamics and low-rank suppression.
+Our analysis shows that low-rank suppression is a general feature of networks with linearly stable dynamics and “strong” low-rank structure in the sense that the dominant singular values of the network are large.
+On the other hand, high-dimensional dynamics are less general, and only arise when the network’s connectivity matrix has a non-vanishing component that satisfies the equal projector (EP) property.
+To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been pointed out in any previous work.
+We show that many common structural features such as biased weights, modularity, and spatial connectivity naturally produce connectivity matrices with the EP property, implying that these networks produce high-dimensional dynamics in addition to low-rank suppression.
 We also draw connections between low-rank suppression and the mathematical theory of balanced networks [21], extending previous work in this direction [3].
@@
 Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}.
-Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Supplementary Figure S.1).
+Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 10 in Appendix D).
 However, this conclusion does not necessarily carry over to recurrent networks.
@@
 which was generated identically to, but independently from $\bf{u}$}.
-Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Supplementary Figure S.1).
+Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 10 in in Appendix D).
 In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
 The results in Figure 1 contrast with some previous modeling work demonstrating low-dimensional dynamics in low-rank recurrent networks [2, 5, 3, 4, 1].
-In Supplementary Text S.1, we provide a detailed review of this previous work and its relation to ours.
+In Appendix A, we provide a detailed review of this previous work and its relation to ours.
 A simple explanation for the low-rank suppression observed in Figure 1 is given by considering the steady-state-solution to Eq. (1),
@@
 Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue.
+It is widely known that an eigenvalue with large, negative real part produces a suppressive dynamic.
 Likewise, our observation of high-dimensional dynamics might seem like a simple consequence of high-dimensional input.
@@
 Hence, low-dimensional dynamics emerged even in the presence of high-dimensional input.
+It might be tempting to conclude that non-normal structure implies low-dimensional dynamics, but in the next section we will show that EP matrices, not normal matrices, are the correct class of matrices to constrain the dimensionality of network dynamics.
 In addition, there was a direction in which variability was especially weak (last blue dot in Figure 2d), similar to our previous example.
@@
 To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics.
-We next derive precise conditions for low-rank suppression and low-dimensional dynamics.
+In the next section, we derive precise conditions for low-rank suppression and low-dimensional dynamics.
 General conditions for low-rank suppression and high-dimensional dynamics
@@
 (6), so care must be taken in interpreting our results.
-For illustrative purposes, we consider a contrived example of such a model in Supplementary Figure S.2.
+For illustrative purposes, we consider a contrived example of such a model in Figure 11 in Appendix D.
 Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix.
@@
 The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive.
-However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Supplementary Figures S.4 and S.5 for two examples of nonlinear, unstable dynamics).
+However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Figures 13 and 14 in Appendix D for two examples of nonlinear, unstable dynamics).
 Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state.
@@
 This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$.
-We discuss the relationships between their analysis and ours in more depth in Supplementary Text S.1.
+We discuss the relationships between their analysis and ours in more depth in Appendix A.
 If a model is fully linear, i.e., if a model is already written in the form of Eq.
@@
 This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
-[…]
+Strongly low rank structure is a special case of low-rank structure.
+For example, a network could have a small number of $\mathcal{O}(1)$ singular values with the remaining singular values $o(1)$.
+Such a network would still be effectively low-rank in an asymptotic sense, but it would not be strongly low-rank according to our definition because the dominant singular values are not asymptotically large.
+These “weakly low-rank” networks [2] can exhibit a weaker form of low-rank suppression and high-dimensional dynamics, depending on the details of their structure (see Appendix A for further discussion and examples).
 Low-rank suppression occurs under Assumptions 1 and 2.
@@
 (8).
-In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is symmetric or normal.
+In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is normal.
 This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
@@
 Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
-Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Supplementary Figure S.1 for which ${\bf{z}}=W{\bf{x}}$ in the steady-state).
+Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 10 in Appendix D for which ${\bf{z}}=W{\bf{x}}$ in the steady-state).
 More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$).
@@
 This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
-We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Supplementary Text S.2 for more mathematical details).
+We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Appendix B for more mathematical details).
 The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics.
-When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Supplementary Figure S.2 for an example).
+When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 11 in Appendix D for an example).
 We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions.
@@
 ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
-In Supplementary Text S.2, we give a more mathematically detailed definition of this assumption.
-In Supplementary Text S.3, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
+In Appendix B, we give a more mathematically detailed definition of this assumption.
+In Appendix C, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
 In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d.
@@
 Conditions for low-dimensional dynamics.
-In Supplementary Text S.2, we prove that (under Assumptions 1, 2, and 3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
+In Appendix B, we prove that (under Assumptions 1–3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
 $P=V^{T}U$
@@
 This matrix is also sometimes referred to as the “overlap matrix” [11].
-Our main result from Supplementary Text S.2 can be stated as follows:
+Our main result from Appendix B can be stated as follows:
 Claim 2.
@@
 a) Singular values of $W$ demonstrate an effective low-rank structure.
-b) The alignment matrix, $P$, when $W_{0}$ is normal.
-c) The variance explained by each principal component of the inputs, ${\bf{x}}(t)$, and network response, ${\bf{z}}(t)$, demonstrates high-dimensional dynamics.
-d,e) Same as b,c except $W_{0}$ is EP, but non-normal.
-f,g) Same as a,b except the first left singular vector is orthogonal to all right singular vectors, so $W_{0}$ is not EP.
-h,i) Same as a,b except the left and right singular vectors of $W_{0}$ are only partially aligned (Eq.
-(10) with $c=0.5$).
+b) The alignment matrix, $P$, when $W_{0}$ is normal and symmetric.
+c) The variance explained by each principal component of the network dynamics, ${\bf{z}}(t)$.
+d) The eigenvalues of $W$.
+e–g) Same as b–d except $W_{0}$ is EP, but non-normal.
+h–j) Same as b–d except $W_{0}$ is not EP.
+k–m) Same as h–j but using a different non-EP structure.
+n–p) Same as b–d except $W_{0}$ is non-EP with a non-vanishing EP component.
+See the text for the precise definition of $W_{0}$ in each case.
 To better understand the conditions on $P$ in Claim 2, we consider simulations of a widely used model from computational neuroscience [27, 32, 2, 33],
@@
 while exploring different choices of $V$.
-We first consider a case in which $W_{0}$ is normal by setting
+We first consider a case in which $W_{0}$ is normal and symmetric by setting
 $V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&-{\bf{u}}_{2}\end{array}\right].$
@@
 In this case, $P$ is diagonal with $P_{kk}=-1$ (Figure 3b) so all singular values of $P$ are $\sigma_{P}=1$ and the network produces high-dimensional dynamics (Figure 3c).
+Like the example from Figure 1, $W_{0}$ has eigenvalues with large, negative real part (Figure 3d).
 In this example, $W_{0}$ is both normal and symmetric.
 Results are similar when $W_{0}$ is normal and asymmetric (obtained, for example, by applying a rotation to $V$ in $r=2$ dimensions) except the conclusion $|P_{kk}|=1$ only holds when the non-zero eigenvalues of $W_{0}$ are real, which is necessarily the case when the $\sigma_{k}$ are distinct.
-However, high-dimensional dynamics do not require that $W_{0}$ is normal.
+High-dimensional dynamics do not require that $W_{0}$ is normal.
 Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$.
@@
 If $W_{0}$ is an EP matrix, then $\sigma_{P}=1$ for all singular values of $P$, so the condition in Claim 2 is satisfied.
-An example of a non-normal, EP matrix $W_{0}$ with rank 2 is given by taking
+An example of a non-normal, EP matrix, $W_{0}$, is given by taking
 $V$ $=\left[\begin{array}[]{cc}{\bf{u}}_{2}&-{\bf{u}}_{1}\end{array}\right].$
-The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$, $P$ is zero along the diagonal (Figure 3d), and $|\lambda_{W_{0}}|\neq\sigma_{W_{0}}$.
-Regardless, all singular values of $P$ are $\sigma_{P}=1$, so the network produces high-dimensional dynamics (Figure 3e).
+The resulting matrix $W_{0}=U\Sigma V^{T}$ is highly non-normal because ${\bf{u}}_{1}\cdot{\bf{v}}_{1}={\bf{u}}_{2}\cdot{\bf{v}}_{2}=0$.
+For this example, $P$ is zero along the diagonal (Figure 3e), but all singular values of $P$ are $\sigma_{P}=1$, so the network still produces high-dimensional dynamics (Figure 3f).
+Note that $W$ does not have any eigenvalues with large, negative real part (Figure 3g).
+Hence, this example combined with the previous example (Figure 3b–d) show that high-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
 To obtain an example of a network that produces low-dimensional dynamics (see also the example in Figure 2), we set
-$V$ $=\left[\begin{array}[]{cc}{\bf{v}}_{\perp}&-{\bf{u}}_{2}\end{array}\right].$
-where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=0$ and another at $\sigma_{P}=1$ (Figure 3f).
-Because of the singular value at zero, the condition in Claim 3 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3g).
-Finally, we consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
-$V$ $=\left[\begin{array}[]{cc}\sqrt{1-c}{\bf{v}}_{\perp}-\sqrt{c}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
+$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{1}&{\bf{v}}_{\perp}\end{array}\right].$
+where $\bf{v}$$_{\perp}$ is orthogonal to ${\bf{u}}_{1}$ and ${\bf{u}}_{2}$, so $P$ has one singular value at $\sigma_{P}=1$ and another at $\sigma_{P}=0$ (Figure 3h).
+Because of the singular value at zero, the condition in Claim 2 is not met, and the network produces low-dimensional dynamics in which one principal component captures an outsized proportion of the variability (Figure 3i).
+For this example, $W$ has an eigenvalue with large, negative real part (Figure 3j).
+However, low-dimensional dynamics also arise in the case (Figure 3k,l)
+$V$ $=\left[\begin{array}[]{cc}-{\bf{u}}_{2}&{\bf{v}}_{\perp}\end{array}\right].$
+in which case $W$ lacks eigenvalues with large, negative real part (Figure 3m).
+Hence, these two examples demonstrate that low-dimensional dynamics can arise in the presence or absence of eigenvalues with large, negative real part.
+So far, in all of the examples with high-dimensional dynamics, $W_{0}$ is an EP matrix.
+However, $W_{0}$ need not be fully EP for dynamics to be high-dimensional.
+We now consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
+$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
 where $0<c<1$.
-In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3h).
-As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3i).
+This is a linear combination of the EP and non-EP matrices from the examples above.
+In this case, $W_{0}$ is not EP, but $P$ has a singular value at $\sigma_{P}=\sqrt{c}$ and the other at $\sigma_{P}=1$ (Figure 3n).
+As long as $c$ is not close to zero, the condition in Claim 2 is satisfied, and the network produces high-dimensional dynamics (Figure 3o).
+Also, $W$ has eigenvalues with large, negative real part (Figure 3p).
+From this example, we may conclude that the low-rank part of the connectivity matrix only needs to have a non-vanishing EP component to produce high-dimensional dynamics.
+The numerical results in Figure 3 provide an interpretation of Claim 2: If $W_{0}$ has a non-vanishing component with the EP property, then dynamics are high-dimensional.
+To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been made before.
+The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks with high-dimensional input.
+As indicated above by Claim 1, the eigenvalue spectrum is also not useful for determining whether strongly low-rank networks produce low-rank suppression.
 It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean.
-In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Supplementary Figure S.3; compare to Figure 2).
+In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 12 in Appendix D; compare to Figure 2).
 However, many networks arising in nature do not have purely random structure.
@@
 If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large.
-The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Supplementary Figure S.6; see Appendix B in [35] for related results).
+The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix D; see Appendix B in [35] for related results).
 Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network.
@@
 In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24].
-As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Supplementary Figure S.7), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
+As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Figure 16 in Appendix D), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
 Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$.
@@
 Connectivity is also perturbed by a random component, $W_{1}$, as above.
-This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Supplementary Figure S.8).
-Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Supplementary Figure S.9).
+This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix D).
+Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix D).
 Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes.
@@
 Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51].
-This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Supplementary Figure S.10a).
+This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix D).
 In [1], it was shown that the dynamics generated by Eq.
@@
 $\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
-Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Supplementary Figure S.10b) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1].
+Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Figure 19b in Appendix D) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1].
 These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics.
@@
 We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
-Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics.
-In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Supplementary Text S.1 for a more complete comparison).
+Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics under some assumptions.
+In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Appendix A for a more complete comparison).
 Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
-Low-rank suppression might seem obvious in hindsight: When $W_{0}$ is normal and symmetric, the combination of stability and large singular values implies eigenvalues with large, negative real part, so dynamics are highly suppressive in one direction.
-Similarly, high-dimensional dynamics might seem like an obvious consequence of high-dimensional inputs.
-However, the example in Figure 2 coupled with the results in Claims 1 and 2 show that reality is more nuanced: Large negative eigenvalues are not necessary for low-rank suppression, and high-dimensional input is not sufficient for high-dimensional responses.
+Our results also rely on a linearization of the dynamics around a stable fixed point.
+The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question.
+However, nonlinear networks can produce any dynamics [29, 30] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks.
+Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then using bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations.
+Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
+It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d).
+A large, negative eigenvalue is well-known to produce a suppressive dynamic.
+However, the results in Figure 2, Figure 12 in Appendix D, and Claim 1 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression.
+Low-rank suppression occurs in strongly low-rank networks regardless of their eigenvalue spectrum.
 Our results have implications for networks in nature and applications.
@@
 All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code.
-All differential equations – except for Figure 6 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
-The simulations in Figure 6 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1].
+All differential equations – except for Figure 6 and Figure 19 in Appendix D – were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
+The simulations in Figure 6 and Figure 19 in Appendix D were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1].
 Code to produce all figures is available at
@@
 This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
-Supplementary Materials
-Figure S.1: Response properties of a feedforward network with rank-one structure.
-Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$.
-Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
-Figure S.2: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure.
-Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$.
-The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a).
-The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
-Figure S.3: Response properties of a network with independent singular left- and right-singular vectors.
-Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors.
-The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
-Figure S.4: Dynamics of an unstable, chaotic network with rank-one structure.
-Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor (c=2c=-2), and the dynamics obey Eq.
-(9).
-The instability produced by taking $\rho>1$ generates chaotic dynamics [27, 3].
-Figure S.5: Dynamics of an unstable, non-chaotic network with rank-one structure.
-Same as Figure 1 except $c=10$ and the dynamics obey Eq.
-(9).
-The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
-Figure S.6: Response properties of a network with biased weights.
-Same as Figure 1 except we used the dynamics in Eq.
-(9), and $W_{jk}$ are drawn i.i.d.
-from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
-Figure S.7: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
-Figure S.8: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
-Figure S.9: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
-Figure S.10: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.
-S.1 Comparison to previous work on dynamics of low-rank networks
+Appendix
+Appendix A Comparison to previous work on dynamics of low-rank networks
 In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al.
@@
 [2, 6, 4, 7, 8, 9] who refer to their networks as “weakly low-rank.”
-Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
-Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
 The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent.
@@
 In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$.
-Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2 and Supplementary Figure S.3 which are strongly low-rank without large eigenvalues).
+Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 12 in Appendix D which are strongly low-rank without large eigenvalues).
 We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]).
-In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Supplementary Figure S.11 and S.12).
-When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Supplementary Figure S.11).
+In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Figures 7 and 8 in Appendix D).
+Figure 7: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
+Figure 8: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
+When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 7).
 Similar examples and results are considered in [10].
@@
 Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$).
-When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Supplementary Figure S.12), reversing the trend of low-rank suppression.
+When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 8), reversing the trend of low-rank suppression.
 However, the amplification is weak.
-In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Supplementary Figure S.12d; compare to Figure 2 and Supplementary Figure S.3).
+In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 8d; compare to Figure 2 and Figure 12).
 Landau and Sompolinsky [5, 3] consider strongly low-rank networks.
@@
 are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
-S.2 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
+Appendix B Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
 We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d).
@@
 In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank.
-However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Supplementary Figure S.2), and would not be related to the intrinsic dynamics of interactions within the network.
+However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Figure 11), and would not be related to the intrinsic dynamics of interactions within the network.
 In conclusion, when ${\bf{x}}(t)$ is high-dimensional, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
@@
 where $\sigma_{Q,N-k}$ is the (Nk)(N-k)th singular value of $Q$, i.e., the $k$th from the last singular value.
-Figure S.13: Comparison between theory and simulations for shorter and longer simulations.
+Figure 9: Comparison between theory and simulations for shorter and longer simulations.
 a) Same as Figure 1d except we added the theoretical values form Eq.
@@
 To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq.
-(16) as red dots (Supplementary Figure S.13a).
+(16) as red dots (Figure 9a).
 Surprisingly, the theory did not closely match the simulations.
@@
 Using Eq.
-(17) gives a much more accurate prediction (Supplementary Figure S.13a, green dots).
-This confirms that the error in the red dots from Supplementary Figure S.13a is due largely to under-sampling of ${\bf{x}}(t)$.
+(17) gives a much more accurate prediction (Figure 9a, green dots).
+This confirms that the error in the red dots from Figure 9a is due largely to under-sampling of ${\bf{x}}(t)$.
 We next increased the simulation time ten-fold to $T=5\times 10^{4}$.
 In this case, the original Eq.
-(16) was more accurate (Supplementary Figure S.13b, red dots), further confirming that the errors in Supplementary Figure S.13a are due largely to sampling error.
+(16) was more accurate (Figure 9b, red dots), further confirming that the errors in Figure 9a are due largely to sampling error.
 As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values).
@@
 This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).
-Next, in Supplementary Text S.3, we generalize away from the assumption of slow inputs using a Laplace transform.
-S.3 The analysis of fast and transient dynamics.
+Next, in Appendix C, we generalize away from the assumption of slow inputs using a Laplace transform.
+Appendix C The analysis of fast and transient dynamics.
 Above, we considered slow or static perturbations, ${\bf{x}}(t)$.
@@
 Note also that Eq.
-(13) from Supplementary Text S.2 can be derived from Eqs.
+(13) from Appendix B can be derived from Eqs.
 (20) and (19) above by first taking $s=-2\pi if$ in Eq.
 (20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$.
-Hence, the approach in this section is a generalization of the approach in Supplementary Text S.2.
+Hence, the approach in this section is a generalization of the approach in Appendix B.
 To interpret Eq.
@@
 (21).
+Appendix D Additional Simulations and Data
+This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text.
+Figure 10: Response properties of a feedforward network with rank-one structure.
+Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$.
+Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
+Figure 11: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure.
+Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$.
+The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a).
+The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
+Figure 12: Response properties of a network with independent singular left- and right-singular vectors.
+Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors.
+The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
+Figure 13: Dynamics of an unstable, chaotic network with rank-one structure.
+Same as Figure 1 except $\rho=2$ and the magnitude of $c$ was increased by the same factor (c=40c=-40) to maintain the same degree of low-rank structure (the same ratio between the spectral radius of the random part, $\rho$, and the magnitude of the low-rank part, $|c|$), and the dynamics obey Eq.
+(9).
+The instability produced by taking $\rho>1$ generates chaotic dynamics.
+Figure 14: Dynamics of an unstable, non-chaotic network with rank-one structure.
+Same as Figure 1 except $c=10$ and the dynamics obey Eq.
+(9).
+The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
+Figure 15: Response properties of a network with biased weights.
+Same as Figure 1 except we used the dynamics in Eq.
+(9), and $W_{jk}$ are drawn i.i.d.
+from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
+Figure 16: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
+Figure 17: Dominant singular vectors of a spatial network are Fourier modes. The singular vectors corresponding to the ten largest singular values of the network in Figure 5.
+Figure 18: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
+Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.

### v5

@@
 Here, we study the dimensionality of dynamics in recurrent networks with low-dimensional structure driven by high-dimensional inputs or perturbations.
-We find that dynamics in such networks can be high- or low-dimensional and we derive novel, precise conditions on the network structure under which dynamics are high-dimensional.
+We find that dynamics in such networks can be high- or low-dimensional and we derive mathematical conditions on the network structure under which dynamics are high-dimensional.
 In many low-rank networks, dynamics are suppressed in directions aligned with the network’s low-rank structure, a phenomenon we term “low-rank suppression.” We show that several low-rank network structures arising in nature satisfy the conditions for generating high-dimensional dynamics and low-rank suppression.
-Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of the dynamics it generates.
-Introduction
-Recent work shows that many networks arising in nature and applications have low-dimensional structure [1].
+Our results clarify important, but counterintuitive relationships between a recurrent network’s connectivity structure and the structure of its response to external input.
+preprint: APS/123-QED
+I Introduction
+Recent work shows that many networks arising in nature and applications have low-rank structure [1].
 This observation raises an important question: What is the relationship between the dimensionality of a network’s structure and the dimensionality of the dynamics on the network?
-In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
+In Neuroscience, for example, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 2, 12].
 Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural activity [13, 14, 15].
@@
 A growing number of more recent studies have shown that neural activity can be high-dimensional, particularly in response to high-dimensional stimuli or behavior [16, 15, 17, 18, 19, 20].
-These observations raise the important point that the dimensionality of dynamics on a network likely depends on the structure of the network’s external input, in addition to the network’s internal structure.
+These observations raise the important point that the dimensionality of dynamics on a network depends on the structure of the network’s external input, in addition to the network’s internal structure.
+There is a large body of theoretical work on the dimensionality of dynamics on low-rank networks [21, 22, 2, 6, 3, 7, 5, 8, 9, 10, 1, 6, 23, 4, 24, 10, 9, 25, 11, 12, 2].
 Some theoretical and computational studies demonstrating low-dimensional dynamics in low-rank networks focus on networks that are self-contained in the sense that they operate in the absence of noise or external inputs or perturbations [2, 6, 1].
@@
 In reality, networks in nature are rarely isolated from internal or external perturbations or inputs, and these perturbations are not necessarily aligned with the network’s low-dimensional structure.
-Other work [2, 4] considers the case of general external input, but the recurrent network is assumed to be weakly low-rank in the sense that the largest singular values are not asymptotically larger than the bulk of the singular values.
-Here, we consider networks with high-dimensional external input and strongly low-rank structure, in which the largest few singular values are much larger than the bulk.
-We show that such networks can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the detailed structure of the network connectivity.
-Perhaps counterintuitively, we find that these networks suppress dynamics in directions aligned to their low-dimensional structure, an effect we call “low-rank suppression.”
-We use mathematical analysis to derive precise conditions under which networks with low-rank structure produce high-dimensional dynamics and low-rank suppression.
-Our analysis shows that low-rank suppression is a general feature of networks with linearly stable dynamics and “strong” low-rank structure in the sense that the dominant singular values of the network are large.
-On the other hand, high-dimensional dynamics are less general, and only arise when the network’s connectivity matrix has a non-vanishing component that satisfies the equal projector (EP) property.
-To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been pointed out in any previous work.
-We show that many common structural features such as biased weights, modularity, and spatial connectivity naturally produce connectivity matrices with the EP property, implying that these networks produce high-dimensional dynamics in addition to low-rank suppression.
+Other work [3, 5] considers the case of general external input to networks that are weakly low-rank in the sense that the largest singular values are $\mathcal{O}(1)$ in magnitude, consistent with connection weights that scale like $1/N$ where network size, $N$, is assumed to be large.
+In this study, we consider networks with strongly low-rank structure in the sense that they have a small number of asymptotically large singular values, while the rest are small or moderate in magnitude.
+This is consistent with connection weights that are asymptotically larger than $1/N$ in magnitude, often taken to be $\mathcal{O}(1/\sqrt{N})$ in theoretical work.
+Importantly, we study the dimensionality of the dynamics on these networks when they are driven by external input.
+The dimensionality of activity driven by intrinsic dynamics is a fundamentally different question treated in previous work [1, 6, 2].
+We also focus on dynamics near linearly stable equilibrium.
+We show by example that networks under these conditions can produce high- or low-dimensional dynamics in response to high-dimensional inputs, depending on the network’s structure.
+Perhaps counterintuitively, we also find that networks under these conditions can suppress inputs aligned to their low-dimensional structure relative to randomly aligned inputs, an effect we call “low-rank suppression.”
+We use mathematical analysis to derive conditions under which low-rank suppression and high-dimensional dynamics can occur in strongly low-rank networks close to a stable equilibrium.
+Our analysis shows that low-rank suppression arises generally in these networks.
+High-dimensional responses to high-dimensional external inputs, on the other hand, arise when the dominant low-rank part of the network’s connectivity matrix has a non-vanishing component with the EP property (meaning that its column space is equal to its row space).
+Low-dimensional dynamics can arise in these networks through a form of non-normal amplification [26, 27, 28], but our mathematical analysis shows that this form of non-normal amplification can only occur when the network is strongly non-EP.
+Strong non-normality is not sufficient.
+To our knowledge, this connection between the EP property of matrices, non-normal amplification, and the dimensionality of recurrent network dynamics has not been described in previous work.
+We show that many common structural features such as biased weights, modularity, and spatial connectivity structure naturally produce strongly low-rank connectivity matrices with low-rank parts satisfying the EP property, implying that these networks produce low-rank suppression and high-dimensional responses to high-dimensional inputs.
 We also draw connections between low-rank suppression and the mathematical theory of balanced networks [29], extending previous work in this direction [4].
@@
 This counterintuitive observation could be used to design more effective interventions to epidemiological, biological, social, and other networks.
-Results
-High-dimensional dynamics and low-rank suppression in a network with rank-one structure
-For illustrative purposes, we begin with a simple, linear model (Figure 1a)
+II High-dimensional dynamics and low-rank suppression in a network with rank-one structure
+Figure 1: Response properties of a recurrent network with rank-one structure.
+a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}.
+b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure.
+c) The eigenvalues of $W$ have a dominant, negative term.
+d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue).
+e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}.
+g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$).
+The norm of $\bf{z}$} is more than 11 times smaller in response to the aligned versus random perturbation.
+h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
+For illustrative purposes, we begin with a simple, linear recurrent network model (Figure 1a)
 $\tau\frac{d{\bf{z}}}{dt}$ $=-{\bf{z}}+W{\bf{z}}+{\bf{x}}.$ (1)
-where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ connectivity matrix, $W$, takes the form
+where ${\bf{x}}(t)$ is an external perturbation or input, ${\bf{z}}(t)$ is the network response, and $\tau=1$. The $N\times N$ recurrent connectivity matrix, $W$, takes the form
 $W=W_{0}+W_{1}$
@@
 $|c|\gg\rho$
-then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (bounded by $2\rho$; Figure 1b).
+then $W$ is “effectively low-rank” [1] in the sense that it has one large singular value near $|c|$ and the remaining singular values are much smaller (approximately bounded by $2\rho$; Figure 1b).
 Similarly, $W$ has one large eigenvalue near $c$ and the remaining eigenvalues lie approximately within a circle of radius $\rho$ in the complex plane (Figure 1c) [33, 34].
@@
 In simulations here, we take $\rho=0.5$ and c=10c=-10.
-We consider more general network structures later.
+We consider more general models and network structures later.
 We first provided a high-dimensional input, ${\bf{x}}(t)$.
-Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input.
+Specifically, for our first simulation, each ${\bf{x}}_{j}(t)$ was an i.i.d., smooth, stationary Gaussian process, which models internal noise, external perturbations, or external input to the network.
 Conventional wisdom might lead us to expect a low-dimensional network response dominated by variability in the direction of $\bf{u}$}.
-Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 10 in Appendix D).
+Indeed, a feedforward network with the same connectivity matrix and same input produces approximately one-dimensional dynamics because it amplifies inputs aligned to $\bf{u}$} (Figure 7 in Appendix F).
 However, this conclusion does not necessarily carry over to recurrent networks.
@@
 Consistent with this finding, the variance of ${\bf{z}}(t)$ in the direction of $\bf{u}$} was more than 132 times smaller than the variance of ${\bf{z}}(t)$ along a random direction (Figure 1e) even though the variability of ${\bf{x}}(t)$ was similar in each direction (Figure 1f).
-Figure 1: Response properties of a recurrent network with rank-one structure.
-a) Schematic of model: The connectivity matrix, $W$, quantifies connections between nodes, $\bf{z}$}, which receive external perturbations or input, $\bf{x}$}.
-b) The singular values of $W$ have one dominant term, indicating approximate rank-one structure.
-c) The eigenvalues of $W$ have a dominant, negative term.
-d) The distribution of variance across principal components of a Gaussian stochastic input (${\bf{x}}(t)$; green) and the response (${\bf{z}}(t)$; blue).
-e,f) The network response (e) and input (f) projected onto the plane determined by $\bf{u}$} and a random vector, ${\bf{u}}_{\textrm{rand}}$, demonstrates low-rank suppression along $\bf{u}$}.
-g) The network response (${\bf{z}}(t)$, top) and its norm ($\|{\bf{z}}(t)\|$, bottom) given an input aligned to the low-rank structure of the network (${\bf{x}}_{\textrm{aligned}}$) and a random input (${\bf{x}}_{\textrm{rand}}$).
-h) Local network input (purple) cancels with external input (green) to produce suppressed network responses (blue) in the direction of $\bf{u}$}.
 To better understand these results, we next simulated the network with two different external input patterns. One pattern was aligned with the low-dimensional structure of $W$,
@@
 which was generated identically to, but independently from $\bf{u}$}.
-Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 10 in in Appendix D).
+Intuitively, we might expect the network to respond more strongly to the aligned input than to the random input, as in a feedforward network (Figure 7 in in Appendix F).
 In reality, we observed exactly the opposite (Figure 1g): The response to ${\bf{x}}_{\textrm{random}}$ was more than 11 times larger than the response to ${\bf{x}}_{\textrm{aligned}}$.
@@
 (1) are highly compressive along $\bf{u}$}.
-This second explanation relies on the assumption that $W_{0}$ is a symmetric matrix (or at least a normal matrix) but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
+These explanations rely on the assumption that $W_{0}$ is a symmetric matrix, or at least a normal matrix, but we will next consider a case in which $W_{0}$ is non-normal and therefore non-symmetric.
 An interesting consequence of low-rank suppression is that external perturbations cancel nearly perfectly with recurrent inputs in the direction of $\bf{u}$}.
@@
 (1) can be interpreted as a vector of internal input to each node, whereas ${\bf{x}}(t)$ is external input and ${\bf{z}}(t)$ is the network response.
-In the qausi-steady state,
+In the quasi-steady state, we have
 ${\bf{z}}\approx W{\bf{z}}+{\bf{x}}.$ (3)
@@
 This fact might appear to present a paradox because the direction $\bf{u}$} is amplified in the product $W{\bf{z}}$, but $W{\bf{z}}$ is one component of $\bf{z}$} (Eq.
-(3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}.
+(3)) and the direction $\bf{u}$} is suppressed in $\bf{z}$}, as we saw in Figure 1e,g.
 This apparent paradox is resolved by a cancellation between ${\bf{W}}{\bf{z}}$ and $\bf{x}$} in the direction of $\bf{z}$}.
-Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h).
+Specifically, under low-rank suppression, internal input cancels nearly perfectly with external input in the direction of $\bf{u}$} (i.e., ${\bf{u}}\cdot W{\bf{z}}\approx-{\bf{u}}\cdot{\bf{x}}$) so that the response, $\bf{z}$}, is weak in the direction of $\bf{u}$} (Figure 1h).
 We refer to this effect as “low-rank cancellation,” which is closely related to the theory of excitatory-inhibitory balance in neural circuits [29, 36] as we will show later.
-Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
+III Low-dimensional dynamics and low-rank suppression in a non-normal network with rank-one structure
+Figure 2: Response properties of a non-normal recurrent network with rank-one structure.
+Same as Figure 1 except the low-rank part of $W$ is non-normal.
+The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c).
+The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
+The norm of $\bf{z}$} is more than 13 times smaller in response to the aligned versus random perturbation.
 Our observation of low-rank suppression in the example above might seem unsurprising due to the presence of a large, negative eigenvalue.
@@
 In this example, low-rank suppression arises in the absence of a large negative eigenvalue and the network’s response is low-dimensional even though the input is high-dimensional.
-Figure 2: Response properties of a non-normal recurrent network with rank-one structure.
-Same as Figure 1 except the low-rank part of $W$ is non-normal.
-The network still exhibits low-rank suppression and cancellation (f, g, h) even in the absence of strong negative eigenvalues (c).
-The network exhibits low-dimensional dynamics (d), in contrast to Figure 1.
 We specifically consider an example in which the model and all parameters are the same as in Figure 1 except that
@@
 This last principal component direction, ${\bf{u}}_{N}$, was partially aligned to $\bf{v}$} (the angle between ${\bf{u}}_{N}$ and $\bf{v}$} was $13^{\circ}$).
-While the variability along $\bf{v}$} itself was not much smaller than a random directions (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f).
+While the variability along $\bf{v}$} itself was not much smaller than a random direction (due to imperfect alignment between ${\bf{u}}_{N}$ and $\bf{v}$}), the variability along ${\bf{u}}_{N}$ was 190 times smaller than the variability in a random direction (Figure 2f).
 Also, low-rank cancellation was observed in the direction of ${\bf{u}}_{N}$ (Figure 2h).
 Low-rank suppression was observed in the sense that an input in the direction of ${\bf{x}}_{\textrm{aligned}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/c$ produced a much weaker response than input in a random direction (Figure 2g).
-The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in a later section, but note that it is closely aligned to $\bf{u}$} whenever $c$ is large.
-In other words, inputs in the direction of ${\bf{x}}_{\textrm{aligned}}$, which is close to $\bf{u}$}, are suppressed by the network.
+The choice of ${\bf{x}}_{\textrm{aligned}}$ will be discussed in Section IV.5 and Appendix D, but note that ${\bf{x}}_{\textrm{aligned}}$ is closely aligned to $\bf{u}$} whenever $c$ is large.
 Hence, low-rank suppression and low-rank cancellation arise even in the absence of a dominant negative eigenvalue.
 To summarize our observations from the two examples above (Figures 1 and 2), both networks produced low-rank suppression and low-rank cancellation, but only the network with a non-normal connectivity matrix (Figure 2) produced low-dimensional dynamics.
-In the next section, we derive precise conditions for low-rank suppression and low-dimensional dynamics.
-General conditions for low-rank suppression and high-dimensional dynamics
+In the next section, we derive more general conditions for low-rank suppression and low-dimensional dynamics.
+IV General conditions for low-rank suppression and high-dimensional dynamics
 Above, we considered a simple network with rank-one structure.
-Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and derive precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
-Assumption 1: Dynamics are linear or near a stable equilibrium.
+Here, we consider more general classes of networks, specify a list of assumptions that we make about the networks, and precise conditions for low-rank suppression and low-dimensional dynamics under these assumptions.
+We first state two assumptions that we make throughout this section.
+Then, we show that low-rank suppression occurs under these assumptions alone.
+Finally, we analyze the conditions for high-dimensional dynamics under various assumptions on ${\bf{x}}(t)$.
+IV.1 Assumption 1: Dynamics are linear or near a stable equilibrium.
 We consider a general class of dynamics of the form
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},\bm{\eta})$ (4)
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+{\bf{F}}({\bf{z}},{\bf{x}})$ (4)
 for some smooth vector field, ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$, where $\tau>0$ sets the timescale of the dynamics.
-Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and $\bm{\eta}(t)$ is an $N\times 1$ vector of external perturbations, inputs, and/or noise.
-In general, the dependence of ${\bf{F}}({\bf{z}},\bm{\eta})$ on $\bf{z}$} defines interactions between nodes, but it is not immediately clear how to define a fixed network of interactions, $W$, from the dynamics in Eq.
+The dependence of ${\bf{F}}({\bf{z}},\boldsymbol{x})$ on $\bf{z}$} defines interactions between nodes, but it is difficult to define a fixed network of interactions, $W$, from the general dynamics in Eq.
 (4).
-When dynamics are linear, $\bf{F}$} can be written as ${\bf{F}}({\bf{z}},\bm{\eta})=W{\bf{z}}+{\bf{G}}(\bm{\eta})$, and we can take $W$ to represent the effective connectivity.
 In nonlinear systems, the effective connectivity between nodes can change with the state of the system.
-To resolve this difficulty in defining network structure when dynamics are nonlinear, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
-Specifically, we assume that there exists a time-constant forcing, $\bm{\eta}(t)=\bm{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$.
-In other words, there exist $\bm{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
-${\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})={\bf{z}}_{0}.$
-We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
-$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$
-has eigenvalues with strictly negative real part.
-Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point.
-We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
-$\bm{\eta}_{p}(t)=\bm{\eta}_{0}+\epsilon\bm{\eta}(t).$
-The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\bm{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
-${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$ (5)
-The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (6)
-where
-$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})$ (7)
-is interpreted as the effective connectivity matrix and
-${\bf{x}}(t)=\partial_{\bm{\eta}}{\bf{F}}({\bf{z}}_{0},\bm{\eta}_{0})\bm{\eta}(t)$
-is interpreted as the effective input. Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
-In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\bm{\eta})={\bf{H}}({\bf{z}})+\bm{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\bm{\eta}$.
-In other models, the raw input, $\bm{\eta}$, to the nonlinear system in Eq.
-(4) can have a different structure and dimensionality to the effective input, $\bf{x}$}, to the linearized system in Eq.
-(6), so care must be taken in interpreting our results.
-For illustrative purposes, we consider a contrived example of such a model in Figure 11 in Appendix D.
-Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix.
-For example, the system might be written as $d{\bf{z}}/dt={\bf{H}}(W{\bf{z}})+{\bf{x}}$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{H}}(W{\bf{z}}_{0})\neq W$.
-In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality.
-We consider one such example from epidemiology at the end of the Results.
-The assumption of linearized dynamics around a stable equilibrium is somewhat restrictive.
-However, nonlinear and unstable dynamics can produce any dynamics [29, 30], so it is difficult to make any general statements about the dimensionality and structure of dynamics in networks far from a stable equilibrium (see Figures 13 and 14 in Appendix D for two examples of nonlinear, unstable dynamics).
-Moreover, the interpretation of network connectivity in nonlinear networks far from equilibrium is not clear, since effective interactions can change with network state.
-Previous work [1] considered models of the form $d{\bf{z}}/dt={\bf{G}}({\bf{y}},{\bf{z}})$ where ${\bf{y}}=W{\bf{z}}$ (note that this model lacks external input).
-This parameterization gives an explicit representation of the network structure, $W$, in nonlinear systems, and it allowed the authors to make statements about the dimensionality of dynamics away from equilibrium using global bounds on the Jacobian matrix, $\partial_{{\bf{y}}}{\bf{G}}({\bf{y}},{\bf{z}})$.
-We discuss the relationships between their analysis and ours in more depth in Appendix A.
-If a model is fully linear, i.e., if a model is already written in the form of Eq.
-(6), then we do not need to assume that dynamics remain close to the equilibrium, so inputs and perturbations need not be weak when dynamics are linear.
-In summary, we assume that the network dynamics are either fully linear or, if they are nonlinear, they represent a small perturbation away from a stable equilibrium.
-This assumption allows us to focus our analysis on the linearized dynamics in Eq.
-(6).
-Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
-We now formalize an asymptotic notion of a strongly low-rank network.
-Specifically, we consider the limit of large network size, $N\to\infty$, and assume that $W$ has $r\sim\mathcal{O}(1)$ asymptotically large singular values,
-$\sigma_{k}\gg 1,\;\;k=1,\ldots,r$
-where $\sigma_{k}$ denotes the $k$th largest singular value of $W$. We assume that the remaining singular values are small or moderate in magnitude,
-$\sigma_{j}\leq\mathcal{O}(1),\;\;j=r+1,\ldots,N.$
-Later, we consider some examples in which these scaling laws arise naturally.
-Matrices that satisfy the assumptions above can be decomposed as [1]
+Moreover, nonlinear recurrent networks can produce any dynamics (high- or low-dimensional) in response to any input [37, 38], so it is difficult to make any general statements about the dimensionality and structure of dynamics in recurrent networks far from a stable equilibrium.
+To resolve this difficulty, we assume that dynamics represent a small deviation away from a stable equilibrium so that we can linearize them.
+A change of coordinates under this linearization gives Eq.
+(1) where $W$ is defined as the Jacobian of $\bf{F}$} at the fixed point ($W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$; see Appendix B for details).
+Our assumption of stability guarantees that all eigenvalues of $W$ have real part less than 1.
+To avoid systems close to instability, we additionally assume that the eigenvalues are bounded away from 1, in other words $W$ does not have any eigenvalues close to $1$.
+For our examples, this implies that the spectral radius, $\rho$, of the random part, $W_{1}$, is not close to $1$.
+Eigenvalues near $1$ or eigenvalues larger than 11 can produce low-dimensional dynamics through a different mechanism than the mechanisms that we are studying [22, 21, 2, 23, 1] (see Figures 1012 in Appendix F for examples).
+IV.2 Assumption 2: The connectivity matrix is strongly low rank in an asymptotic sense.
+We additionally assume that the network, $W$, has a strongly low-rank structure.
+To make this assumption precise, we consider networks in the limit of large $N$ and assume that a small number of singular values diverge with $N$ while the rest remain $\mathcal{O}(1)$ or smaller (see Appendix C for a more precise statement).
+This assumption has two components.
+First, that the connectivity matrix has a small number of dominant singular values.
+In later sections, we show that this property arises naturally in several natural network structures such as networks with biased weights, modular structure, or smooth spatial structure.
+Secondly, this assumption implies that the dominant singular values grow large with $N$.
+This contrasts to other work on “weakly” low-rank networks in which the largest singular values are $\mathcal{O}(1)$ [3, 7, 5, 8, 9, 10].
+Weakly low-rank structure arises naturally when mean weights scale like $\mathcal{O}(1/N)$ whereas strongly low-rank structure arises when mean weights are asymptotically larger than $1/N$, for example when they scale like $\mathcal{O}(1/\sqrt{N})$ as in balanced networks [29, 39, 40, 41, 4].
+See Appendix A for a more in-depth comparison between weakly and strongly low-rank networks.
+Strongly low-rank networks can be decomposed as [1]
 $W=W_{0}+W_{1}$
-where $W_{0}$ has rank $r$ and asymptotically large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank.
+where $W_{0}$ has rank $r\sim\mathcal{O}(1)$ and large singular values, while the singular values of $W_{1}$ are $\mathcal{O}(1)$ at most and $W_{1}$ can be full rank.
+We specifically assume that $W_{1}$ is a random matrix and independent from $W_{0}$.
 Since $W_{0}$ has rank $r$, we can write its singular value decomposition as
 $W_{0}=U\Sigma V^{T}$
-where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values ($\Sigma_{kk}=\sigma_{k}\gg 1$) while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
-We additionally assume that $W_{1}$ is a random matrix that is independent of $U$ and $V$.
-In other words, $W$ is a low-rank matrix perturbed by noise.
-This is a stronger assumption than necessary ($W_{1}$ need not be fully independent from $W_{0}$ for our results to hold) but it simplifies our analysis.
-Strongly low rank structure is a special case of low-rank structure.
-For example, a network could have a small number of $\mathcal{O}(1)$ singular values with the remaining singular values $o(1)$.
-Such a network would still be effectively low-rank in an asymptotic sense, but it would not be strongly low-rank according to our definition because the dominant singular values are not asymptotically large.
-These “weakly low-rank” networks [2] can exhibit a weaker form of low-rank suppression and high-dimensional dynamics, depending on the details of their structure (see Appendix A for further discussion and examples).
-Low-rank suppression occurs under Assumptions 1 and 2.
-Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d). Specifically,
-Claim 1.
-Consider the dynamics defined by Eq.
-(6) under Assumptions 1 and 2.
-There exists an input ${\bf{x}}_{a}$ with $\|{\bf{x}}_{a}\|=\mathcal{O}(1)$ such that the network’s steady-state response to ${\bf{x}}_{a}$ is asymptotically larger than the $\mathcal{O}(1)$ response to a random input, ${\bf{x}}_{\textrm{r}}$ (where $\|{\bf{x}}_{\textrm{r}}\|=\mathcal{O}(1)$).
-To show why Claim 1 is true, first choose a $k\in\{1,\ldots,r\}$ and define
+where $\Sigma$ is a diagonal, $r\times r$ matrix of singular values
+$\sigma_{k}=\Sigma_{kk}\gg 1$
+while $U=[{\bf{u}}_{1}\cdots{\bf{u}}_{r}]$ and $V=[{\bf{v}}_{1}\cdots{\bf{v}}_{r}]$ are $N\times r$ orthonormal matrices with columns ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$ that define the left- and right-singular vectors of $W_{0}$.
+IV.3 Low-rank suppression occurs under Assumptions 1 and 2.
+Our first result is that Assumptions 1 and 2 are sufficient for low-rank suppression (as observed in Figures 1d and 2d).
+Specifically, we claim that low-rank suppression is realized under a static input of the form
 ${\bf{x}}_{a}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (5)
-Note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$.
+where $k\leq r$ so that $\sigma_{k}\gg 1$ is one of the dominant singular values.
+We claim that inputs of this form are strongly suppressed by the network.
+To see why this is true, first note that $\|{\bf{x}}_{a}\|=1+o(1)$ since $\sigma_{k}\gg 1$, $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, and the maximum singular value of $W_{1}$ is $\mathcal{O}(1)$.
 Now note that the steady-state response to the input ${\bf{x}}_{a}$ is given by
@@
 Since $\|{\bf{v}}_{k}\|=1$ and $\sigma_{k}\gg 1$, we have that $\|{\bf{z}}_{a}\|\ll 1$.
-Since $\sigma_{k}\gg 1$, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq.
-(8).
-In summary, low-rank suppression is a generic phenomenon in linearly stable, strongly low-rank networks, regardless of whether connectivity is normal.
-This explains why low-rank suppression was observed in both examples from Figure 1 and Figure 2.
+Therefore, the network strongly suppresses the input ${\bf{x}}_{a}$ from Eq.
+(5) in the sense that $\|{\bf{z}}_{a}\|\ll\|{\bf{x}}_{a}\|$.
 Note from Eq.
@@
 Also note that the response, ${\bf{z}}_{a}$, is perfectly aligned to ${\bf{v}}_{k}$.
-Finally, note that there are $r$ different choices for ${\bf{x}}_{a}$ (one for each $k\in\{1,\ldots,r\}$), all of which are nearly orthogonal to each other.
-Interestingly, the input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 10 in Appendix D for which ${\bf{z}}=W{\bf{x}}$ in the steady-state).
+Hence, strongly low-rank recurrent networks suppress inputs aligned near the dominant left singular vectors (${\bf{u}}_{k}$) and the suppressed response is aligned with the dominant right singular vectors (${\bf{v}}_{k}$).
+Interestingly, these input-response directions along which suppression occurs are reversed from the input-response directions along which amplification occurs in a feedforward network with the same connectivity (for example, the network in Figure 7 in Appendix F for which ${\bf{z}}=W{\bf{x}}$ in the steady-state).
 More specifically, in a feedforward network with connectivity $W$, inputs aligned to the dominant right singular vectors (${\bf{v}}_{k}$) produce amplified responses in the direction of the corresponding left singular vectors (${\bf{u}}_{k}$).
@@
 We call this form of suppression low-rank suppression.
-In later sections, we will explore consequences of low-rank suppression in common low-rank network structures.
-Before discussing the dimensionality of dynamics, we first need to make an additional assumption about the dimensionality of the input.
-Assumption 3: External inputs are high-dimensional and slowly varying.
-To begin discussing the dimensionality of ${\bf{z}}(t)$, we need a model of ${\bf{x}}(t)$ that allows us to quantify dimensionality in a concrete way.
-To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process in which case ${\bf{z}}(t)$, as defined in Eq.
-(6), is also a stationary ergodic process whenever dynamics are stable [31].
-This allows quantify the dimensionality of ${\bf{x}}(t)$ and ${\bf{z}}(t)$ in terms of their principal components or, equivalently, by the eigenvalues of their stationary covariance matrices.
-We assume that ${\bf{x}}(t)$ is high-dimensional in the sense that the singular values of the covariance matrix of ${\bf{x}}(t)$ (i.e., the variance explained by the principle components of ${\bf{x}}(t)$) do not have a small number of dominant terms (as in Figure 1d; see Appendix B for more mathematical details).
-The assumption that ${\bf{x}}(t)$ is high-dimensional means that we are focused on the question of whether the network intrinsically generates low-dimensional dynamics.
-When ${\bf{x}}(t)$ is low-dimensional, low-dimensional dynamics in ${\bf{z}}(t)$ can be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 11 in Appendix D for an example).
-We additionally assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions.
+Low rank suppression occurs regardless of whether connectivity is normal, explaining why low-rank suppression was observed in both Figure 1 and Figure 2.
+Later, we will explore consequences of low-rank suppression in common network structures.
+First, we discuss conditions for the separate but related phenomenon of high-dimensional dynamics.
+IV.4 Assumption 3: External inputs are stationary, high-dimensional, and slowly varying.
+To begin discussing the dimensionality of ${\bf{z}}(t)$, we need to make assumptions that allow us to define precisely what we mean by high- and low-dimensional dynamics.
+To this end, we assume that ${\bf{x}}(t)$ is a stationary ergodic stochastic process.
+In this case ${\bf{z}}(t)$ is also a stationary ergodic process whenever dynamics are stable [42] as in Assumption 1.
+This allows us to quantify the dimensionality of ${\bf{z}}(t)$ in terms of its principal components or, equivalently, by the singular values of its stationary covariance matrix.
+Specifically, the variance explained by the first $k$ principal components represents the amount of variability constrained to a $k$-dimensional linear manifold.
+For this reason, we define ${\bf{z}}(t)$ to have low-dimensional dynamics when there exists an $r\sim\mathcal{O}(1)$ such that $\sigma_{j}/\sigma_{k}\to 0$ as $N\to\infty$ whenever $k\leq r$ and $j>r$.
+This implies that the vast majority of variability lies on a low-dimensional linear manifold.
+Otherwise, we say that ${\bf{z}}(t)$ has high-dimensional dynamics.
+We also assume that ${\bf{x}}(t)$ is high-dimensional.
+If ${\bf{x}}(t)$ were low-dimensional, then low-dimensional dynamics in ${\bf{z}}(t)$ could be inherited from ${\bf{x}}(t)$ instead of being generated intrinsically (see Figure 8 in Appendix F for an example).
+In assuming that ${\bf{x}}(t)$ is high-dimensional, we are focusing on the question of whether low-dimensional dynamics arise from network interactions, not from low-dimensional input alone.
+For simplicity, within this section, we specifically assume that the components, ${\bf{x}}_{j}(t)$, are i.i.d.
+This implies that ${\bf{x}}(t)$ is high-dimensional, is a stronger assumption than necessary, but simplifies the exposition.
+In Appendix D, we consider more general inputs that are not assumed to be i.i.d.
+For now, we also assume that ${\bf{x}}(t)$ varies in time more slowly than the intrinsic timescale, $\tau$, of the network interactions.
 To make this assumption more precise, note that whenever ${\bf{x}}(t)={\bf{x}}$ is constant, solutions, ${\bf{z}}(t)$, converge to the steady-state given by ${\bf{z}}=[I-W]^{-1}{\bf{x}}$.
 We assume that ${\bf{x}}(t)$ varies sufficiently slowly that ${\bf{z}}(t)$ is approximated by the quasi-steady state approximation,
-${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$
-In Appendix B, we give a more mathematically detailed definition of this assumption.
-In Appendix C, we also provide a preliminary generalization of our analysis to the case where ${\bf{x}}(t)$ is not assumed to vary slowly, in which case the dimensionality of transient dynamics can be analyzed in the Laplace domain.
-In simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d.
-Gaussian process with correlation time $\tau_{x}=10$.
-Our assumption of high-dimensionality is satisfied because each ${\bf{x}}_{j}(t)$ is independent.
-Our assumption that ${\bf{x}}(t)$ varies slowly is satisfied because $\tau_{x}$ is much larger than the network’s intrinsic timescale, $\tau=1$.
-Conditions for low-dimensional dynamics.
-In Appendix B, we prove that (under Assumptions 1–3) the dimensionality of dynamics depends on the recurrent alignment matrix [3] of $W_{0}$, defined by
-$P=V^{T}U$
-This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$.
-Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$.
-This matrix is also sometimes referred to as the “overlap matrix” [11].
-Our main result from Appendix B can be stated as follows:
-Claim 2.
-Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
-Here, “low-dimensional dynamics” is interpreted to mean that the covariance matrix of ${\bf{z}}(t)$ has an $\mathcal{O}(1)$ number of dominant eigenvalues, equivalently, ${\bf{z}}(t)$ has a small number of dominant principal component directions.
-Note that this result is equivalent to the following:
-Claim 2.
-Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If all singular values of $P=V^{T}U$ are $\mathcal{O}(1)$ then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
-Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
+${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t).$ (6)
+In Appendix D, we give more details about this assumption. In Appendix E, we extend our analysis and results to the case that ${\bf{x}}(t)$ is not slowly varying.
+IV.5 Conditions for low-dimensional dynamics under Assumptions 1–3
 Figure 3: Conditions for high-dimensional dynamics in a network with rank-two structure.
@@
 See the text for the precise definition of $W_{0}$ in each case.
+Under the quasi-steady state approximation in Eq.
+(6) and the assumption that elements of ${\bf{x}}(t)$ are i.i.d., it is easy to show that the covariance matrix of ${\bf{z}}(t)$ can be written as [42]
+$\overline{C}^{\bf{z}}\propto[I-W]^{-1}[I-W]^{-T}.$
+As a result, the question of whether ${\bf{z}}(t)$ is low-dimensional becomes a question of whether $I-W]^{-1}$ has a small number of dominant singular values whenever $W$ has a small number of dominant singular values.
+More details are given in Appendix D where we prove that the answer to this question depends on the recurrent alignment matrix [4] of $W_{0}$, defined by
+$P=V^{T}U.$
+This $r\times r$ matrix measures the alignment between the left and right singular vectors of $W_{0}$.
+Specifically, $P_{jk}={\bf{v}}_{j}\cdot{\bf{u}}_{k}$ measures the alignment between the $j$th right singular vector and the $k$th left singular vector of $W_{0}$.
+This matrix is also sometimes referred to as the “overlap matrix” [12].
+Our main result from Appendix D can be stated as follows:
+Claim 1.
+Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
+Note that this result is equivalent to the following:
+Claim 1.
+Consider the dynamics defined by Eq. (4) under Assumptions 1–3. If all singular values of $P=V^{T}U$ are $\mathcal{O}(1)$ then the dynamics of ${\bf{z}}(t)$ are high-dimensional.
+Recall that all singular values of $P$ are bounded by $1$, so the condition that all singular values of $P$ are $\mathcal{O}(1)$ simply says that no singular values of $P$ are asymptotically small, hence the two versions of Claim 2 are logical contrapositives.
+We were unable to prove the converse of these claims (that dynamics are low-dimensional whenever $P$ has a small singular value), even though it is consistent with all of our simulations.
+Proving the converse or finding a counterexample could be a fruitful direction for future research.
+Our proof of Claim 1 relied on an assumption that ${\bf{x}}(t)$ varies more slowly than the timescale, $\tau$, of network interactions.
+In Appendix E, we consider the case where ${\bf{x}}(t)$ is not necessarily slow.
+In that case, the dimensionality can be broken down into the contribution from different frequency modes.
+When ${\bf{x}}(t)$ varies at both high and low frequencies, the results in Claim 1 are still valid (see Figure 14 for an example).
 To better understand the conditions on $P$ in Claim 1, we consider simulations of a widely used model from computational neuroscience [35, 43, 3, 44],
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}.$ (9)
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W\tanh({\bf{z}})+{\bf{x}}$ (7)
+with $\tau=1$.
+In these simulations, we take each ${\bf{x}}_{j}(t)$ to be a smooth, i.i.d.
+Gaussian process with correlation time $\tau_{x}=10$.
 We take $W_{0}$ to have rank $r=2$ so $W$ has two dominant singular values (Figure 3a) and we fix
@@
 Instead, it is sufficient that $U$ and $V$ share a column space, $\textrm{col}(U)=\textrm{col}(V)$.
-Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called “equal projector” (EP) matrices [34].
+Low-rank matrices, $W_{0}=U\Sigma V^{T}$, for which $U$ and $V$ share a column space are called EP matrices [45].
 All normal matrices are EP, but an EP matrix is not necessarily normal.
@@
 We now consider the case in which left and right singular vectors are not perfectly aligned, but have some non-vanishing overlap,
-$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$ (10)
+$V$ $=\sqrt{1-c}\left[\begin{array}[]{cc}-{\bf{u}}_{1}&\,{\bf{v}}_{\perp}\end{array}\right]+\sqrt{c}\left[\begin{array}[]{cc}{\bf{-}}{\bf{u}}_{1}&\,-{\bf{u}}_{2}\end{array}\right]$
 where $0<c<1$.
@@
 To our knowledge, this connection between the EP property of matrices and the dimensionality of network dynamics has not been made before.
-The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks with high-dimensional input.
-As indicated above by Claim 1, the eigenvalue spectrum is also not useful for determining whether strongly low-rank networks produce low-rank suppression.
+The examples in Figure 3 also demonstrate that the eigenvalue spectrum is generally not useful for determining whether dynamics are high- or low-dimensional in strongly low-rank networks near a stable equilibrium with high-dimensional input.
+Instead, as Claim 1 shows, the singular values of the recurrent alignment matrix, $P$, are key to determining the dimensionality of dynamics in these networks.
+When low-dimensional dynamics are present, they arise from an effectively feedforward mechanism that takes variability from ${\bf{x}}(t)$ and $W_{1}{\bf{z}}(t)$ aligned with the column space of $V$ and maps it to strong variability along the column space of $U$ (see Appendix D for more details).
+This mechanism is a form of non-normal amplification [26, 27, 28].
+Our results add some new results to the literature on non-normal amplification under the conditions in Assumptions 1–3.
+First, our results show that this form of non-normal amplification requires that $W$ is highly non-EP, not just highly non-normal.
+See, for example, the example in Figure 3e–g, in which $W_{0}$ is highly non-normal, but EP so that amplification does not occur.
+Secondly, our results on low-rank suppression show that the amplified responses in the direction of $U$ associated with non-normal amplification are accompanied by suppressed responses approximately aligned to $V$ (see Figure 2e,f).
+Finally, Claim 1 shows that non-normal amplification is the only way to generate low-dimensional dynamics under Assumptions 1–3.
 It is common in theoretical work to consider models in which the entries of $U$ and $V$ are random and independent with zero mean.
-In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 12 in Appendix D; compare to Figure 2).
+In this case, ${\bf{u}}_{j}$ and ${\bf{v}}_{k}$ are nearly orthogonal so $\sigma_{P}\ll 1$ and the network produces low-dimensional dynamics (Figure 9 in Appendix F; compare to Figure 2).
 However, many networks arising in nature do not have purely random structure.
-In the following sections, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
-Biased weights and modular networks
+Next, we show that many naturally arising network structures satisfy our conditions for high-dimensional dynamics.
+V Biased weights and modular networks
+Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network.
+a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure.
+c) The alignment matrix shows that the network is EP, but not normal.
+d) An input that is constant within each block (purple) is aligned to the low-rank part, but a random input (green) is not.
+e) The norm of $\bf{z}$} is more than 13 times smaller in response to the aligned versus random perturbation.
+f) The excitatory (positive; red) component of mean input (local and external combined) balances with the inhibitory (negative; blue) component to produce a much smaller total input (gray), a widely observed phenomenon in neural circuits.
+Network dynamics obey Eq.
+(7).
 So far, we considered networks with unbiased weights, $E[W_{jk}]=0$, which is common in modeling studies, but many networks in nature have weights with non-zero mean.
@@
 If $m$ and $s$ scale similarly, then the largest singular value of $W$ is near $\sigma_{1}=|m|N$ while the next-largest singular value scales like $\sigma_{2}=\mathcal{O}(s\sqrt{N})$, implying an effective rank-one structure when $N$ is large.
-The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix D; see Appendix B in [35] for related results).
-Figure 4: Low-rank suppression and excitatory-inhibitory balance in a modular network.
-a,b) A modular network with biased blocks modeling excitatory and inhibitory neurons has low-rank structure.
-c) The alignment matrix shows that the network is EP, but not normal.
-d) An input that is constant within each block (red) is aligned to the low-rank part, but a random input (gray) is not.
-e) Response magnitude is suppressed for the aligned input relative to the random input.
-f) The excitatory (positive; red) component of an input balances with the inhibitory (negative; blue) component to produce a much smaller total (gray) component, a widely observed phenomenon in neural circuits.
+The dominant rank-one part has constant entries, so it is normal and the network exhibits low-rank suppression and high-dimensional dynamics (Figure 15 in Appendix F; see Appendix B in [46] for related results).
 More generally, modular structure arises in many natural settings [1, 47].
@@
 $W=\left[\begin{array}[]{ccc}W^{1,1}&\cdots&W^{1,n}\\ W^{2,1}&\cdots&W^{2,n}\\ \vdots&\cdots&\vdots\\ W^{n,1}&\cdots&W^{n,n}\end{array}\right]$
-where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$. In general, each sub-matrix can have a different, non-zero mean and variance,
+where $W^{a,b}$ is a $N_{a}\times N_{b}$ sub-matrix quantifying connections from population $b$ to population $a$.
+In general, each sub-matrix can have a different, non-zero mean and standard deviation,
 $E\left[W^{a,b}_{jk}\right]=m_{ab}\;\textrm{ and }\;\textrm{std}\left(W^{a,b}_{jk}\right)=s_{ab}.$
@@
 $=\{[a\;\cdots\;a\;b\;\cdots\;b]^{T}\,|\,a,b\in{\mathbb{R}}\}$
-where ${\bf{1}}_{e}=\{1,2,\ldots,N_{e}\}$ and ${\bf{1}}_{i}=\{N_{e}+1,N_{e}+2,\ldots,N\}$.
-Therefore, perturbations that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e).
-In other words, perturbations that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to perturbations that are inhomogeneous within one or both populations, consistent with previous work on balanced network models of cortical circuit responses [43, 23, 22, 24].
-As predicted, simulations also show high-dimensional responses to high-dimensional inputs (Figure 16 in Appendix D), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
-Averaging over the excitatory and inhibitory populations represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$.
+where ${\bf{1}}_{e}$ and ${\bf{1}}_{i}$ are indicator vectors for the excitatory and inhibitory populations.
+Therefore, external inputs, $\bf{x}$}, that are uniform within each population are aligned to the low-rank part and suppressed relative to random perturbations (Figure 4d,e).
+In other words, external inputs that stimulate all excitatory neurons equally and all inhibitory neurons equally are suppressed relative to inputs that are inhomogeneous within one or both populations, consistent with previous work on balanced network models [24, 31, 30, 32].
+As predicted, simulations also show high-dimensional responses to high-dimensional stimuli (Figure 16 in Appendix F), consistent with observations that neural responses in monkey visual cortex are high-dimensional when visual stimuli are high-dimensional [16].
+Averaging over the excitatory and inhibitory populations in this network represents a projection onto $\textrm{col}(U)=\textrm{col}(V)$.
 As a result, the cancellation mechanism illustrated in Figure 1h manifests as a tight balance between mean excitatory (positive) and inhibitory (negative) input to neurons (Figure 4f), a phenomenon that is widely observed in neural recordings [51, 52, 53, 54, 55, 56, 57] and widely studied in computational models [29, 39, 32, 4].
-Hence, the widely studied theory of balanced network models can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to a particular modular network structure.
-Amplification of disordered perturbations in networks with spatial structure
+Hence, the widely studied theory of balanced networks can be interpreted as a special case of the low-rank suppression and low-rank cancellation studied here, specific to the particular modular network structure arising from Dale’s law.
 Figure 5: Low-rank suppression in a network with spatial structure.
@@
 b) Singular values of $W$ demonstrate effective low-rank structure.
-c) A spatially smooth perturbation (red) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (gray) is not.
+c) A spatially smooth perturbation (purple) is aligned to the low-rank part of $W$ while a spatially disordered perturbation (green) is not.
 d,e) Therefore, the network response to a smooth perturbation is much weaker than the response to a disordered perturbation (compare vertical tick marks).
+The norm of $\bf{z}$} is more than 7 times smaller in response to the smooth versus disordered perturbation.
+Network dynamics obeyed Eq.
+(7).
+VI Amplification of disordered perturbations in networks with spatial structure
 Many networks in nature exhibit connection strength that depends smoothly on the distance between nodes in physical or other spaces, resulting in an effective low-rank structure [1, 40].
-As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection strength decays like a Guassian function of the distance, $d\theta$, between nodes so that only nearby nodes are strongly connected (Figure 5a).
+As a simple example, we consider a model in which each node is assigned a spatial location, $\theta\in[0,1)$, and connection weights in $W_{0}$ decay like a Guassian function of the distance, $d\theta$, between nodes computed with periodic boundaries.
+We assume strong coupling in the sense that individual weights scale like $1/\sqrt{N}$.
+In this case, only nearby nodes are strongly connected (Figure 5a).
+Note that stability requires that connections in $W_{0}$ are negative since coupling is strong.
 Connectivity is also perturbed by a random component, $W_{1}$, as above.
-This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix D).
-Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix D).
+This connectivity structure is effectively low-rank (Figure 5b) and the low-rank part is normal, with left- and right singular vectors spanned by the low-frequency Fourier basis vectors (Figure 17 in Appendix F).
+Therefore, networks of this form satisfy our conditions for low-rank suppression and high-dimensional dynamics (Figure 18 in Appendix F).
 Perturbations that are smooth in space are aligned to the low-rank part of the connectivity matrix because they are formed by sums of low-frequency spatial Fourier modes.
-Hence, perhaps surprisingly, strongly low-rank networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
-Low-rank suppression and high-dimensional dynamics in an epidemiological network.
+Hence, perhaps surprisingly, strongly connected networks with spatial structure are more sensitive to spatially disordered perturbations than to spatially smooth perturbations (Figure 5c-e; compare tick labels on vertical axes).
+VII Low-rank suppression and high-dimensional dynamics in an epidemiological network.
 We next consider a real epidemiological network, specifically a network of high school social contacts [58], which was used in recent theoretical work on low-rank network dynamics [1].
 In that work, the authors considered a quenched mean-field reduction of the susceptible-infected-susceptible model,
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ W{\bf{z}}$ (11)
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ Q{\bf{z}}$ (8)
 where $g>0$, $\gamma>0$, and $\circ$ denotes element-wise multiplication.
 Each ${\bf{z}}_{j}(t)$ models the probability that an individual is infected.
-Here, $W$ is the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [51].
-This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix D).
-In [1], it was shown that the dynamics generated by Eq.
+Here, $Q$ is the proportional to the proximity matrix of 637 high school students, indicating whether each pair of students were in proximity of each other during a specific school week [58].
+This matrix is effectively low-rank in the sense that it has a small number of dominant singular values (Figure 19a in Appendix F).
+In [1], it was shown that dynamics generated by Eq.
 (8) on this network are effectively low-dimensional.
@@
 a) The response to a perturbation aligned with the dominant low-rank part is weaker than the response to a perturbation of the same magnitude in a random direction.
+The distance of $\bf{z}$} from its fixed point is more than 10 times smaller in response to the aligned versus random perturbation.
 b) In the presence of high-dimensional random perturbations, the variance explained by each principal component of the dynamics is similar to the input, indicating high-dimensional dynamics.
+Network dynamics obeyed Eq.
+(7).
 However, Eq.
 (8) is completely self-contained without any perturbations.
-Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network (for example, parents and siblings outside of the school) and through the natural stochasticity of infection.
-Since the proximity matrix is necessarily symmetric and therefore normal, our analysis predicts that the network produces low-rank suppression and high-dimensional responses to high-dimensional perturbations.
-To test this prediction, we modified the model by adding an external forcing term,
-$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[W{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$
-Simulations of these dynamics indeed demonstrated low-rank suppression (Figure 6a and Figure 19b in Appendix D) and high-dimensional dynamics (Figure 6b) in contrast to the model from [1].
-These results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics.
-Moreover, the results imply that epidemiological networks are, perhaps counterintuitively, more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
-Discussion
-We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices.
-We also derived conditions under which low-rank networks generate high-dimensional dynamics.
+Perturbations arise in epidemiological dynamics through interactions with individuals from outside of the modeled network and through the natural stochasticity of infection.
+Specifically, ${\bf{z}}(t)$ represents a probability of infection.
+Actual infection events will be random, which can be approximated by adding a stochastic forcing term to Eq.
+(8).
+Moreover, students interact with siblings and parents at home (who are not included in the network represented by $Q$), which produces additional external forcing.
+As a simple model of these perturbations, we modified the model by adding an external forcing term to obtain
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+\gamma(1-{\bf{z}})\circ\left[Q{\bf{z}}+{\bf{x}}(t)\circ{\bf{z}}\right].$ (9)
+Eq.
+(9) admits a fixed point when ${\bf{x}}(t)=0$.
+Following the linearization discussed in Section IV.1 and Appendix B, the effective connectivity, $W$, is defined by the Jacobian matrix at the fixed point.
+We find that, like $Q$, the effective connectivity matrix, $W$, has a small number of dominant singular values, so our theory predicts that this network exhibits low-rank suppression.
+Moreover, the recurrent alignment matrix, $P$, defined from $W$ satisfies the conditions for high-dimensional dynamics from Claim 1 (Figure 19a in Appendix F).
+Simulations of Eq.
+(9) confirm that the model exhibits low-rank suppression (Figure 6a and Figure 19b–d in Appendix F) and high-dimensional dynamics (Figure 6b) in contrast to the model without external input from [1].
+Importantly, the “aligned” input direction that we used to demonstrate low-rank suppression in Figure 6a is defined in terms of the dominant right singular vector of the effective connectivity matrix, $W$, i.e., the Jacobian at the fixed point.
+Using the left singular vector of the explicit connectivity, $Q$, in place of the effective connectivity, $W$, does not produce suppression (see Figure 19c,d in Appendix F).
+This observation validates our linearization argument and highlights the necessity of using effective connectivity in place of explicit connectivity in nonlinear networks.
+Taken together, our results highlight the importance of accounting for external perturbations and internal noise when studying the dimensionality of epidemiological dynamics.
+Moreover, the results imply that epidemiological networks can be more sensitive to random perturbations than to perturbations aligned with the network’s low-rank structure.
+For example, if the forcing term is taken to represent interventions like masking or inoculation, then our results predict that it is more effective to apply these interventions randomly to individuals than to apply interventions to individual sub-populations or other structures reflected in the dominant structure of the proximity matrix.
+VIII Discussion
+We presented theory and examples showing that strongly low-rank recurrent networks close to a stable equilibrium suppress inputs or perturbations aligned with the dominant directions of their connectivity matrices.
+We also derived conditions under which these networks generate high-dimensional dynamics.
 We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
@@
 The dimensionality of nonlinear network dynamics away from a stable equilibrium is also an important question.
-However, nonlinear networks can produce any dynamics [29, 30] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks.
-Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then using bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations.
+However, nonlinear networks can produce any dynamics in response to any inputs [37, 38] and effective connectivity can change with network state, so it is difficult to draw broad, general conclusions about the dimensionality of dynamics in nonlinear networks.
+Previous work [1] circumvented this difficulty by parameterizing the model in terms of the vector of local inputs from the network, ${\bf{y}}=W{\bf{z}}$, and then used bounds on the Jacobian matrix of the dynamics with respect to $\bf{y}$}, but they did not consider external inputs or perturbations.
 Combining our approaches by incorporating external inputs into their analytical approach is a promising direction for future work.
+We focused on the response of low-rank networks to external input, whereas much previous work focuses on the intrinsic dynamics of low-rank networks in the absence of external drive.
+Landau and Sompolinsky [4] consider a similar model to ours, but look at the interaction between externally driven responses and intrinsic, chaotic dynamics.
+However, they restrict to low-dimensional external inputs in the column space of the low-rank part of the connectivity matrix.
+Future work could combine these approaches to better understand the interaction between chaotic, intrinsic dynamics and high-dimensional external input.
 It is tempting to ascribe low-rank suppression to the presence of an eigenvalue with large, negative real part (as in Figure 1 and Figure 3b–d).
 A large, negative eigenvalue is well-known to produce a suppressive dynamic.
-However, the results in Figure 2, Figure 12 in Appendix D, and Claim 1 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression.
-Low-rank suppression occurs in strongly low-rank networks regardless of their eigenvalue spectrum.
+However, the results in Figure 2 and Section IV.5 demonstrate that the eigenvalue spectrum is not generally relevant to the presence or absence of low-rank suppression.
+In other words, in recurrent networks driven by external input near a stable equilibrium, low-rank suppression depends only on the presence of large singular values of the connectivity matrix, $W$.
+High-dimensional dynamics, on the other hand, depend additionally on the singular values of the recurrent alignment matrix, $P=V^{T}U$.
 Our results have implications for networks in nature and applications.
@@
 This result can be leveraged to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
-Methods
-All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code.
-All differential equations – except for Figure 6 and Figure 19 in Appendix D – were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
-The simulations in Figure 6 and Figure 19 in Appendix D were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1].
+IX Methods
+All simulations and analyses were performed in Python using a combination of custom written PyTorch and NumPy code.
+All differential equations, except for Figure 6 and Figure 19 in Appendix F, were solved using a simple forward Euler scheme with a step size of $dt=0.01$.
+The simulations in Figure 6 and Figure 19 in Appendix F were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1].
+Networks in Figures 1 and 2 used the linear model from Eq.
+(1).
+Networks in Figures 35 used the nonlinear model from Eq.
+(7).
 Code to produce all figures is available at
 https://github.com/RobertRosenbaum/HighDimLowDimCode
-Acknowledgments
+Acknowledgements.
 We thank Ashok Litwin-Kumar and David G.
 Clark for helpful conversations and comments on drafts of the manuscript.
-This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
-Appendix
+This work was supported by the Air Force Office of Scientific Research (AFOSR) awards numbers FA9550-21-1-0223 and FA9550-26-1-0004, and the National Science Foundation under a Neuronex award number DBI-1707400.
 Appendix A Comparison to previous work on dynamics of low-rank networks
-In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al.
+Here, we focus on the response of networks to high-dimensional external input.
+Some previous work on low-rank networks and low-dimensional dynamics focuses on the dimensionality of intrinsic dynamics generated by the network in the absence of external input or in the presence of low-dimensional external input.
+For example chaotic and other nonlinear dynamics in untrained and trained networks can be low-dimensional (see [22, 21, 2, 23, 1] and Figures 10 and 11 in Appendix F).
+In contrast, we are here primarily interested in the network’s response to high-dimensional external input.
+In our model’s, the network’s activity is driven primarily by the external input (which is filtered by the network’s internal structure) and not by intrinsic dynamics.
+When the network is strongly low-rank, we showed that low-dimensional dynamics can only arise when the recurrent alignment matrix has small singular values (see Claim 1).
+In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, especially work by Ostojic et al.
 [3, 7, 5, 8, 9, 10], work by Thibeault et al.
@@
 Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [3].
-Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
+Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale whenever the spectral radius of $W_{1}$ is $\mathcal{O}(1)$, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
 This difference in scaling is acknowledged by Ostojic et al.
 [3, 7, 5, 8, 9, 10] who refer to their networks as “weakly low-rank.”
+To summarize, when $W_{1}$ is a random matrix with $\mathcal{O}(1)$ spectral radius and when the entries of $W_{0}$ are biased, the primary difference between many strongly and weakly low-rank networks in practice is that the entries of $W_{0}$ scale like $\mathcal{O}(1/N)$ in weakly low rank networks while they are larger (typically $\mathcal{O}(1/\sqrt{N})$) in many strongly low-rank networks, for example in work on balanced networks [6, 4].
+The transition between these two regimes can be visualized by gradually scaling the magnitude of the low-rank part from weak to strong.
+Figure 22 in Appendix F shows that low-rank suppression arises gradually in this case.
 The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent.
@@
 [11] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values.
-While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [2], the absence of a random part means that their connectivity matrices are truly low-rank.
+While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [3], the absence of a random part means that their connectivity matrices are truly low-rank (not just approximately low-rank).
 In that work [11], Mastrogiuseppe et al.
 showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
-For the sake of comparison to the models in [10], let us consider a rank-one network structure like the one in Figure 1.
-Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
+For the sake of comparison to the models in [11], we consider a rank-one network structure like the one in Figure 1. Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
 $W_{0}=c{\bf{u}}{\bf{u}}^{T}.$
@@
 In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$.
-Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 12 in Appendix D which are strongly low-rank without large eigenvalues).
+Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2, Figure 3k–m, and Figure 9 which are strongly low-rank without large eigenvalues).
 We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [11]).
-In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Figures 7 and 8 in Appendix D).
-Figure 7: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
-Figure 8: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
-When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 7).
+In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ when $|c|=\mathcal{O}(1)$ and $\rho\ll 1$.
+Repeating the simulation from Figure 1 demonstrates a weakened form of low-rank suppression with high-dimensional dynamics when $c<0$ and an amplification of aligned inputs with weakly low-dimensional dynamics when $c>0$ (Figures 20 and 21 in Appendix F).
+We also repeated all of the simulations from Figure 3 in networks with $\rho\ll 1$ and $\sigma_{1},\sigma_{2}\sim\mathcal{O}(1)$, so that the network is still low-rank in the sense that the first two singular values dominate, but not strongly low-rank because they are $\mathcal{O}(1)$ (see Figure 23a).
+Under these conditions, the dynamics were high-dimensional regardless of the structure of $W_{0}$ (Figure 23b–f).
+To understand this result, note that low-dimensional dynamics in our models (Figures 2 and 3i,l) arise from strong amplification of a small number of directions by $W$.
+When the low-rank part of $W$ is weaker, these low-dimensional dynamics will be weaker or absent.
+In other words, when the dominant singular values of $W_{0}$ are $\mathcal{O}(1)$ and smaller, the identity in $I-W]^{-1}$ from the quasi-stead state analysis more strongly dominates the $W$ (or, at least, the $W$ does not dominate the $I$).
+More directly, the ${\bf{x}}(t)$ in Eq.
+(1) dominates the $W{\bf{z}}$.
+Therefore, high-dimensional inputs are simply transferred to high-dimensional responses when $W$ is weaker.
+When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Figure 20 in Appendix F).
 Similar examples and results are considered in [11].
 In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
-Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$).
-When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 8), reversing the trend of low-rank suppression.
+When $|c|=\mathcal{O}(1)$, we can take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$).
+When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Figure 21 in Appendix F), reversing the trend of low-rank suppression.
 However, the amplification is weak.
-In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 8d; compare to Figure 2 and Figure 12).
+In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Figure 21d; compare to Figure 2 and Figure 9) because stability imposes a bound on $c$ (specifically, $c<1$) and therefore on $\|W_{0}\|$.
 Landau and Sompolinsky [6, 4] consider strongly low-rank networks.
@@
 are strongly low-rank, in contrast to the weakly low-rank networks considered by Ostojic et al.
-Appendix B Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
-We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d).
-We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–f.
+Appendix B Details for the linearization of dynamics around a stable equilibrium
+Here, we make precise our assumption that dynamics represent a small perturbation around a stable equilibrium, and give a more precise derivation of the linearized dynamics.
+We begin by considering nonlinear dynamics of the form
+$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},\boldsymbol{\eta})$
+where $\boldsymbol{\eta}(t)$ is the external input before any change of coordinates is applied.
+We assume that there exists a constant-in-time, $\boldsymbol{\eta}(t)=\boldsymbol{\eta}_{0}$, that produces a stable steady-state solution, ${\bf{z}}_{0}$.
+In other words, there exist $\boldsymbol{\eta}_{0}$ and ${\bf{z}}_{0}$ such that
+${\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})={\bf{z}}_{0}.$
+We additionally assume that this fixed point is hyperbolically stable. In other words, the Jacobian matrix
+$J=-I+\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$
+has eigenvalues with strictly negative real part.
+Here, $I$ is the identity matrix and $\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$ is the Jacobian matrix of $\bf{F}$} with respect to $\bf{z}$} evaluated at the fixed point.
+We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
+$\boldsymbol{\eta}_{p}(t)=\boldsymbol{\eta}_{0}+\epsilon\boldsymbol{\eta}(t).$
+The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, $\boldsymbol{\eta}_{p}(t)$, can be written to linear order in $\epsilon$ as
+${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
+The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
+$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+{\bf{x}}(t)$ (11)
+where
+$W=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})$
+is interpreted as the effective connectivity matrix and
+${\bf{x}}(t)=\partial_{\boldsymbol{\eta}}{\bf{F}}({\bf{z}}_{0},\boldsymbol{\eta}_{0})\boldsymbol{\eta}(t)$
+is interpreted as the effective input.
+Our assumption that the equilibrium is stable implies that all eigenvalues of $W$ have real part less than $1$.
+We additionally assume that the real part of all eigenvalues of the Jacobian are bounded away from 0, meaning that the real part of all eigenvalues of $W$ are bounded below $1$.
+In other words, there is a $d>0$ with $d\sim\mathcal{O}(1)$ such that all eigenvalues of $W$ satisfy $\textrm{Re}(\lambda_{W})<1-d$.
+In particular, this implies that $\rho<1-d$ where $\rho$ is the spectral radius of the random part, $W_{1}$, of $W$.
+Note that Eq.
+(11) is equivalent to the linear dynamics in Eq.
+(1).
+Hence, the linear network dynamics in Eq.
+(1) are a sort of normal form for the linearized dynamics described here.
+Therefore, we are justified in focusing on the linear dynamics of Eq.
+(1) in the analysis in the Results.
+In many models, the perturbation is purely additive, ${\bf{F}}({\bf{z}},\boldsymbol{\eta})={\bf{H}}({\bf{z}})+\boldsymbol{\eta}$, so that the effective input is equal to the raw input, ${\bf{x}}=\boldsymbol{\eta}$.
+In other models, the raw input can have a different structure and dimensionality to the effective input.
+For illustrative purposes, we consider a contrived example of such a model in Figure 8 in Appendix F.
+If a model is fully linear, i.e., if a model is already written in the form of Eq.
+(11), then we do not need to assume that dynamics remain close to the equilibrium, so inputs and perturbations need not be weak when dynamics are linear.
+Some models are parameterized in terms of a connectivity matrix that is not equal to the Jacobian matrix.
+For example, the system might be written as $d{\bf{z}}/dt={\bf{G}}(W{\bf{z}},{\bf{x}})$ where $W$ is meant to quantify a connectivity matrix, but $\partial_{\bf{z}}{\bf{G}}(W{\bf{z}}_{0},{\bf{x}}_{0})\neq W$.
+In this case, our interpretation of the effective connectivity matrix is not consistent with the original parameterization, but they will often share structural properties such as their approximate dimensionality.
+We consider one such example from epidemiology at the end of the Results.
+In that example, we found that the use of effective connectivity in place of the explicit connectivity matrix is important for our results to hold (see Figure 19 in Appendix F)
+Appendix C Precise definitions of strongly low-rank networks and low-dimensional dynamics.
+As mentioned above, we focus our attention on strongly low-rank networks.
+Intuitively, this means that the connectivity matrix has a small number of large singular values and the rest are not large.
+To make this definition mathematically precise, we need to consider an asymptotic limit of large $N$, where $N$ is the network size.
+Let $\sigma_{k}$ be the $k$th largest singular value of $W$.
+In strongly low-rank networks, there exists an $r\sim\mathcal{O}(1)$ such that $\lim_{N\to\infty}\sigma_{k}=\infty$ for $k=1,\ldots,r$.
+In addition, $\sigma_{k}\leq\mathcal{O}(1)$ for $k>r$.
+We additionally need to specify what we mean by “low-dimensional” and “high-dimensional” dynamics.
+Let $\lambda_{k}$ be the $k$th largest eigenvalue of the covariance matrix of ${\bf{z}}(t)$, equivalently the $k$th largest singular value since covariance matrices are symmetric and positive definite.
+Note that $\lambda_{k}$ is a non-negative real number representing the amount of variance explained by the $k$th principal component of ${\bf{z}}(t)$.
+Intuitively, ${\bf{z}}(t)$ is low-dimensional whenever a large portion of the variance is explained by a small number of singular values.
+To make this precise, we say that the dynamics of ${\bf{z}}(t)$ are low-dimensional whenever there is an $r\sim\mathcal{O}(1)$ such that $\lim_{N\to\infty}\lambda_{j}/\lambda_{k}=0$ whenever $j>r$ and $k\leq r$.
+Appendix D Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
+We now consider conditions for the emergence of high-dimensional dynamics.
+We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process.
 Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
@@
 where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance.
-When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$.
+When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=Ic_{x}(f)$ is a multiple of the identity matrix where $c_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$.
 The assumption of i.i.d.
-elements is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
-The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics Eq. (6) to get [31, 53, 41, 42, 54]
-$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}$ (13)
+elements is true for all of the examples we consider, but we will not apply this simplification until later in our calculation.
+The the cross-spectral density of the network response, ${\bf{z}}(t)$, is defined analogously and it can be derived under the dynamics in Eq. (11) to get [42, 59, 40, 41, 60]
+$\widetilde{C}^{{\bf{z}}}(f)$ $=[r(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[r(f)I-W]^{-*}$ (12)
 where
-$\widetilde{r}(f)=1-2\pi\tau fi$
+$r(f)=1-2\pi\tau fi$
 is scalar.
 Eq. (12) quantifies the covariance structure of ${\bf{z}}(t)$ at any given frequency mode, $f$, but we are often specifically interested in the zero-lag temporal covariance,
-$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$
+$\overline{C}^{{\bf{z}}}_{jk}=\text{cov}({\bf{z}}_{j}(t),{\bf{z}}_{k}(t))=\int_{-\infty}^{\infty}\widetilde{C}^{{\bf{z}}}_{jk}(f)df.$ (13)
 From Eq. (12), we therefore have
-$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}df$ (14)
-If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$.
+$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[r(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[r(f)I-W]^{-*}df$ (14)
+If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ at frequencies much higher than $1/(2\pi\tau)$ (see Figure 13 in Appendix F) .
 Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$.
-Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero.
+Since $r(f)\approx 1$ whenever $f\ll 1/(2\pi\tau)$, we therefore have that $r(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero (see Figure 13 in Appendix F).
 As a result, the only parts of the integrand that contribute to the integral in Eq.
 (14) are the low frequency components, $f\approx 0$.
-In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
+In this case, we can replace $r(f)$ with $1$ to obtain
 $\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}\overline{C}^{{\bf{x}}}[I-W]^{-T}.$ (15)
@@
 (15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$.
-Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate, but our approach here is more mathematically precise.
+Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>1/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate.
+Either approach can be used to derive Eq.
+(15).
+In Appendix E below, we provide more details about the frequency dependence of variability and how it relates to the speed at which ${\bf{x}}(t)$ varies.
 When principal component analysis is applied to ${\bf{z}}(t)$, the variance explained by each principal component is given by the ordered list of eigenvalues of the covariance matrix, $\overline{C}^{{\bf{z}}}$.
@@
 $R=[I-W]^{-1}\sqrt{v}$
-and therefore
+and
 $\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}[I-W]^{-T}v$
 where recall that $v>0$ is a scalar.
-In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank.
-However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Figure 11), and would not be related to the intrinsic dynamics of interactions within the network.
+In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank (see Figure 8 in Appendix F).
+However, in this situation, low-dimensional dynamics are inherited from the input, not generated by interactions within the network.
+By restricting to high-dimensional ${\bf{x}}(t)$, we are focused on the question of whether low-dimensional dynamics arise from network interactions.
 In conclusion, when ${\bf{x}}(t)$ is high-dimensional, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
 $A=[I-W]^{-1}.$
-More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is given by the square of the singular values of $A$,
+More specifically, the variance explained by each principal component of ${\bf{z}}(t)$ is proportional to the square of the singular values of $A$,
 $\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{A,k}\,v$
@@
 where $\sigma_{Q,N-k}$ is the (Nk)(N-k)th singular value of $Q$, i.e., the $k$th from the last singular value.
-Figure 9: Comparison between theory and simulations for shorter and longer simulations.
-a) Same as Figure 1d except we added the theoretical values form Eq.
-(16) (red dots) and the values obtained from Eq.
-(17) (green dots).
-The network was simulated for $T=5\times 10^{3}$ time units.
-b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.
 To demonstrate these analytical results, we repeated the simulation from Figure 1d and added the predictions from Eq.
-(16) as red dots (Figure 9a).
+(16) as red dots (Figure 24a in Appendix F).
 Surprisingly, the theory did not closely match the simulations.
-We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=5$).
+We suspected that this was due to finite sampling: The simulation was performed over the time interval $t\in[0,T]$ where $T=5\times 10^{3}$ (for comparison, $\tau=1$ and the correlation timescale of ${\bf{x}}(t)$ was $\tau_{x}=10$).
 We suspected that Eq.
@@
 (16) would be replaced by
-$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{U,N-k}}$
+$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{1}{\sigma^{2}_{\hat{B},N-k}}$
 where
-$U=[I-W]\sqrt{{\overline{C}}^{\bf{x}}}^{-1}$
-and the empirical covariance matrix is used for ${\overline{C}}^{\bf{x}}$. Or, equivalently and more simply,
+$\hat{B}=[I-W]\sqrt{{\hat{C}}^{\bf{x}}}^{-1}$
+and ${\hat{C}}^{\bf{x}}$ is the the empirical covariance matrix. Or, equivalently and more simply,
 $\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{\hat{R},k}$ (17)
 where
-$\hat{R}=[I-W]^{-1}\sqrt{{\overline{C}}^{\bf{x}}}$
-is the sampled value of $R$ and the empirical value of ${\overline{C}}^{\bf{x}}$ is again used.
+$\hat{R}=[I-W]^{-1}\sqrt{{\hat{C}}^{\bf{x}}}$
+is the sampled value of $R$.
 Using Eq.
-(17) gives a much more accurate prediction (Figure 9a, green dots).
+(17) gives a more accurate prediction (Figure 24a in Appendix F, green dots).
 This confirms that the error in the red dots from Figure 24a is due largely to under-sampling of ${\bf{x}}(t)$.
@@
 In this case, the original Eq.
-(16) was more accurate (Figure 9b, red dots), further confirming that the errors in Figure 9a are due largely to sampling error.
+(16) was more accurate (Figure 24b in Appendix F, red dots), further confirming that the errors in Figure 24a are due largely to sampling error.
 As concluded previously, ${\bf{z}}(t)$ exhibits low-rank suppression (as demonstrated by the last blue dot in Figure 1d) whenever $Q$ has a small number of asymptotically large singular values (equivalently, whenever $A$ has a small number of asymptotically small singular values).
@@
 Conversely, high-dimensional dynamics can only occur whenever $Q$ does not have any asymptotically small singular values (equivalently, whenever $A$ does not have any asymptotically large singular values).
-Combining these conclusions with those in Claim 1, we can summarize as follows
+We can summarize as follows
 Low-rank suppression (as in Figure 1g) occurs whenever Q=IWQ=I-W has at least one asymptotically large singular value.
 High-dimensional dynamics (as in Figure 1d) can only occur whenever Q=IWQ=I-W lacks any asymptotically small singular values.
-Previously, in proving Claim 1, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is effectively low-rank.
+Previously, in Section IV.5, we effectively showed that Q=IWQ=I-W has large singular values (and therefore low-rank suppression occurs) whenever $W$ is strongly low-rank.
 Conditions under which Q=IWQ=I-W lacks small singular values (and therefore high-dimensional dynamics occur) are not so simple.
@@
 If the dynamics of ${\bf{z}}(t)$ are low-dimensional then $P=V^{T}U$ has at least one asymptotically small singular value.
-The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2 and Figure 3g (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d).
+The dynamics of ${\bf{z}}(t)$ are low dimensional whenever ${\bf{z}}(t)$ has a small number of dominant principal components as in Figure 2d (conversely, ${\bf{z}}(t)$ is high-dimensional whenever there is no such dominant principal component, as in Figure 1d).
 From the discussion above, we know that ${\bf{z}}(t)$ is low dimensional whenever $A=[I-W]^{-1}$ has at least one asymptotically large singular value or, equivalently, whenever Q=IWQ=I-W has an asymptotically small singular value.
@@
 Since $\bf{z}$} is approximately aligned to the $U$ in the proof, we may conclude that low-dimensional dynamics are caused by excess variability along the column space of $U$, i.e., the column space of $W_{0}$.
-The conclusions above were reached by assuming that ${\bf{x}}(t)$ varies more slowly than $\tau$.
-However, note that Eq.
-(13) shows that the same conclusions can be reached in the context of variability at any frequency mode, $f$, regardless of how quickly ${\bf{x}}(t)$ varies.
-Specifically, variability in ${\bf{z}}(t)$ at a particular frequency mode (as quantified by $C^{{\bf{z}}}(f)$) is defined in terms of a regularized inverse of $W$, just like the stationary variance in Eq.
+This explains the observations in Figure 2e.
+More specifically, low-dimensional dynamics arise when external input and the random part of internal input (i.e., $W_{1}{\bf{z}}$) have variability that overlap with the column space of $V$.
+Because these right singular vectors are associated with large singular values of $W$, this variability gets mapped to strong variability in the column space of $U$.
+Since $P$ has small singular values, this variability in the column space of $U$ does not overlap with the column space of $V$, so it cannot be cancelled by recurrent interactions.
+This creates an effectively feedforward mapping that amplifies variability along the directions ${\bf{v}}_{k}$ and maps it to variability in the directions of ${\bf{u}}_{k}$.
+This mechanism is a form of non-normal amplification [26, 27, 28] (see the discussion in Section IV.5 for a more in-depth comparison).
+These results can also help us understand some subtleties of low-rank suppression.
+Our theory from Section IV.3 shows that the input ${\bf{x}}_{a}={\bf{u}}_{k}+(W_{1}{\bf{v}}_{k}+{\bf{v}}_{k})/\sigma_{k}$ is suppressed.
+Since $\sigma_{k}$ is large, this input is close to ${\bf{u}}_{k}$, but has a correction term ${\bf{x}}_{c}=(W_{1}{\bf{v}}_{k}+{\bf{v}}_{k})/\sigma_{k}$.
+In Figure 1g, the “aligned” input that produced a suppressed response was chosen simply as ${\bf{x}}={\bf{u}}$, without the correction term (using ${\bf{x}}_{a}$ with the correction term also produced a suppressed response; not pictured).
+For the example in Figure 2, we used ${\bf{x}}={\bf{x}}_{a}$ with the correction term because using ${\bf{x}}={\bf{u}}$ without a correction term did not produce a suppressed response.
+To understand why this happens, note that networks like the one in Figure 2 (for which $P=V^{T}U$ is singular or has a small singular value) take inputs from the directions ${\bf{v}}_{k}$ and amplify them in proportion to $\sigma_{k}$ (according to our proof of Claim 1).
+Hence, the ${\bf{v}}_{k}/\sigma_{k}$ term in ${\bf{x}}_{a}$ is small by a factor of $\sigma_{k}$, but the response is also amplified by a factor of $\sigma_{k}$, so the correction term produces a $\mathcal{O}(1)$ response and cannot be ignored.
+In summary, the correction term, ${\bf{x}}_{c}$, is necessary to include when $P$ is singular or has small singular values (as in Figure 2).
+Appendix E Dimensionality of dynamics when inputs are not slowly varying.
+In Section IV.5 and Appendix D, we assumed that ${\bf{x}}(t)$ varied on a slower timescale than the timescale, $\tau$, of the network dynamics, justifying the quasi-steady state approximation in Eq.
 (6).
-The only salient difference is that the regularizer is a scalar multiple of the identity, $\widetilde{r}(f)I$, in Eq.
-(13) instead of the identity itself, as in Eq.
-(15).
-This difference could be important, for example, if $\widetilde{r}(f)=0$ at some frequency, $f$, for which $C^{{\bf{x}}}(f)$ is not close to zero (corresponding to a situation in which variability in ${\bf{x}}(t)$ is faster than $\tau$).
-Next, in Appendix C, we generalize away from the assumption of slow inputs using a Laplace transform.
-Appendix C The analysis of fast and transient dynamics.
-Above, we considered slow or static perturbations, ${\bf{x}}(t)$.
-Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models).
-This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics.
-We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales.
-We can no longer rely on the quasi-steady state approximation from Eq.
-(2) in this case.
-Instead, we take the Laplace transform in Eq.
-(6) to obtain
-$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (18)
-where $\hat{\psi}(s)=\int\psi(t)e^{-st}dt$ is the Laplace transform of $\psi(t)$ and
-$\hat{H}(s)$ $=(1+\tau s)^{-1}W$ (19)
-$\hat{B}(s)$ $=(1+\tau s)^{-1}.$
-Solving the implicit Eq. (18) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
-$\hat{{\bf{z}}}=[I-\hat{H}]^{-1}\hat{B}\hat{{\bf{x}}}$. (20)
-The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$ because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$.
-At mode $s=0$ (the static or “DC” mode), we recover the connectivity from the steady-state or quasi-steady state analysis, $\hat{H}(0)=W$.
-Similarly, $\hat{B}(0)=1$.
-Hence, the quasi-steady state analysis from above evaluated how the DC component of the input, $\hat{{\bf{x}}}(0)$, affects the DC component of the dynamics, $\hat{{\bf{z}}}(0)$, even when we do not assume that ${\bf{x}}(t)$ changes slowly.
-The analysis here is a generalization of the quasi-steady state analysis to non-static modes, $s\neq 0$.
-Note also that Eq.
-(13) from Appendix B can be derived from Eqs.
-(20) and (19) above by first taking $s=-2\pi if$ in Eq.
-(20) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$.
-Hence, the approach in this section is a generalization of the approach in Appendix B.
-To interpret Eq.
-(20), let us consider two separate limits.
-At modes that are much slower than the intrinsic timescale of the dynamics, $|s|\ll 1/\tau$, we have that $\hat{H}(s)\approx W$ and $\hat{B}(s)\approx 1$, so
-$\hat{{\bf{z}}}(s)\approx[I-W]^{-1}\hat{{\bf{x}}}(s)$
-and we recover the results from our quasi-steady state analysis at these modes.
-In particular, Claim 1 holds and we expect low-rank suppression at mode $s$ whenever $W$ is strongly low rank.
-While this may at first seem like a repetition of our results from the quasi-steady state analysis, note that we do not need to assume that ${\bf{x}}(t)$ varies slowly here.
-The Laplace analysis at slow modes ($|s|\ll 1/\tau$) applies even when ${\bf{x}}(t)$ has power at slow and fast modes.
-At fast modes, $|s|\gg\tau$, the identity term in Eq. (20) dominates the $\hat{H}$ and $\hat{B}$ terms so that
-$\hat{{\bf{z}}}(s)\approx\hat{{\bf{x}}}(s)$
-at fast modes.
-Therefore, fast modes (representing transient responses to transient input dynamics) do not exhibit low-rank suppression, but they do produce high-dimensional responses to high-dimensional stimuli.
-Modes between these two extremes interpolate between them.
-In summary, when our previous assumption that ${\bf{x}}(t)$ varies more slowly than $\tau$ is violated, we expect to see low-rank suppression at slow temporal modes, but not at fast temporal modes.
-Additionally, working in the Laplace domain allows us to expand our class of models.
-In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq.
-(4), which we later linearized.
-The dynamics in Eq.
-(4) capture a large class of models, but are not fully general.
-Specifically, in Eq.
-(4), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$.
-However, some systems like integral equations, integro-differential equations, operator equations, and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq.
-(4).
-In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
-${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (21)
-In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivariant, causal function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$.
-Eq.
-(21) encompasses an extremely broad class of systems, including those described by Eq.
-(4).
-We again begin by describing how to define a network of interactions, $W$, from the general definition of network dynamics in Eq.
-(21).
-As above, we consider a linearization around the response, ${\bf{z}}_{0}$, to a baseline input, ${\bf{x}}_{0}$.
-However, in this case, the baseline input and response can depend on time, ${\bf{x}}_{0}(t)$ and ${\bf{z}}_{0}(t)$.
-We again consider a weak perturbation away from the baseline, ${\bf{x}}_{p}(t)={\bf{x}}_{0}(t)+\epsilon{\bf{x}}(t),$ which produces a corresponding deviation
-${\bf{z}}_{p}(t)=\mathcal{F}[{\bf{z}}_{p},{\bf{x}}_{p}]={\bf{z}}_{0}(t)+\epsilon{\bf{z}}(t)+\mathcal{O}(\epsilon^{2}).$
-A first order Volterra expansion gives Eq.
-(18) as an implicit equation for ${\bf{z}}(t)$, but the matrices $\hat{H}$ and $\hat{B}$ are no longer defined as above.
-Instead, the matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline.
-Under the dynamics in Eq.
-(6), we recover the $\hat{H}$ and $\hat{B}$ defined in Eq.
-(19).
-Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq.
-(21).
-Appendix D Additional Simulations and Data
-This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text.
+While this assumption simplifies our analysis and conclusions, we can understand the dimensionality of dynamics without it.
+Here, we analyze the dimensionality of ${\bf{z}}(t)$ under all of the conditions from Assumptions 1–3 except the assumption that ${\bf{x}}(t)$ is slowly varying.
+Beginning with Eq.
+(14), we see that the dimensionality of ${\bf{z}}(t)$ in general depends on the interaction between $\widetilde{C}^{{\bf{x}}}(f)$, $W$ and $r(f)$ across different frequencies, $f$.
+Here, we again focus on the special case that each ${\bf{x}}_{j}(t)$ is i.i.d.
+In this case, $\widetilde{C}^{{\bf{x}}}(f)=c_{x}(f)I$ where $c_{x}(f)$ is the scalar power spectral density of each ${\bf{x}}_{j}(t)$.
+Therefore, the covariance matrix of ${\bf{z}}(t)$ can be written as [42]
+$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[r(f)I-W]^{-1}[r(f)I-W]^{-*}c_{x}(f)df$ (18)
+where $\cdot^{-*}$ is the inverse conjugate transpose and $r(f)=1-2\pi\tau fi$.
+First consider $\widetilde{C}^{{\bf{z}}}(f)$ at frequencies lower than the timescale of the dynamics, $f\ll 1/(2\pi\tau)$. At these frequencies, we have that $r(f)\approx 1$ so that
+$\widetilde{C}^{{\bf{z}}}(f)\approx[I-W]^{-1}[I-W]^{-T}c_{x}(f)$
+at sufficiently low frequencies. Since $c_{x}(f)$ is a scalar, we have that
+$\overline{C}^{{\bf{z}}}\propto[I-W]^{-1}[I-W]^{-T}$
+for $f\ll 1/(2\pi\tau)$.
+Therefore, the arguments and conclusions from Section IV.5, Appendix D, and Claim 1 apply in general to the dimensionality of dynamics at low frequencies.
+Specifically, the dimensionality of the dynamics at low frequencies (i.e., the slow dynamics) can only be low-dimensional when $P=V^{T}U$ has a small singular value.
+This is true even when ${\bf{x}}(t)$ has power at high frequencies.
+However, the stationary covariance matrix of ${\bf{z}}(t)$ accounts for variability at all frequencies (see Eqs.
+(18)).
+When ${\bf{x}}(t)$ is slowly varying, $c_{x}(f)\approx 0$ at higher frequencies (see Figure 13), so the low-frequency analysis above is sufficient (consistent with the results obtained in Appendix D and Section IV.5).
+When ${\bf{x}}(t)$ has power at high frequencies, these higher frequencies need to be accounted for.
+At high frequencies satisfying $f\gg\|W\|/(2\pi\tau)$, the $r(f)I$ terms in Eq. (18) dominates the $W$ terms so that
+$\widetilde{C}^{{\bf{z}}}(f)\approx\frac{c_{x}(f)}{1+4\pi^{2}f^{2}\tau^{2}}I$
+which is a multiple of the identity and therefore full-rank.
+Hence, when ${\bf{x}}(t)$ has variability at these high frequencies (i.e., when $c_{x}(f)$ is not close to zero for $f\gg\|W\|/(2\pi\tau)$; see Figure 14a), these high frequency components can only add high-dimensional variability to ${\bf{z}}(t)$ because ${\bf{z}}(t)$ tracks ${\bf{x}}(t)$ at high frequencies.
+If ${\bf{x}}(t)$ has variability only at these high frequencies ($f\gg\|W\|/(2\pi\tau)$), then only these frequencies would contribute to the integral in Eq.
+(18) and therefore ${\bf{z}}(t)$ would be high-dimensional regardless of the structure of $W$.
+However, it is somewhat rare that natural stochastic processes have power only at higher frequencies.
+More commonly, when ${\bf{x}}(t)$ varies quickly, it has power at both low and high frequencies.
+In other words, $c_{x}(f)$ is far from zero for both small and large $f$.
+For example, white noise has equal power at all frequencies and Orenstein Uhlenbeck processes have more power at low than high frequencies even when they have fast timescales.
+When ${\bf{x}}(t)$ has power at low and high frequencies, the integral in Eq.
+(18) has contributions from both low and high frequencies.
+Low-dimensional variability can only be introduced at low frequencies and, by Claim 1 and our arguments above, can only be introduced when $P=V^{T}U$ has a small singular value.
+Since these low-dimensional dynamics represent an asymptotically large singular value, this combination of low-dimensional variability at low frequencies and high-dimensional variability at high frequencies results in low-dimensional variability overall since the low-dimensional variability is strong while the high-dimensional variability is $\mathcal{O}(1)$.
+As a result, when ${\bf{x}}(t)$ has power at both low and high frequencies, the conclusions of Claim 1 apply.
+Specifically, low-dimensional variability arises only when $P$ has small singular values.
+To test this conclusion, we repeated the simulations from Figure 3 with timescales swapped.
+Specifically, we set $\tau=10$ and $\tau_{x}=1$ so that $c_{x}(f)$ is far from zero at both low frequencies ($f\ll 1/(2\pi\tau)$) and high frequencies ($f\gg\|W\|/(2\pi\tau)$; see Figure 14a in Appendix F).
+As predicted, low-dimensional variability only arises in the case that $P=V^{T}U$ has a singular value at zero (Figure 14b–f), identical to the case when ${\bf{x}}(t)$ varies slowly (compare Figure 14 to Figure 3).
+Appendix F Additional Simulations and Data
+This appendix contains results from additional simulations and data. The simulations themselves and data are described in the respective captions and contextualized within the main text and Appendix.
 Figure 7: Response properties of a feedforward network with rank-one structure.
@@
 Figure 8: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure.
-Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$.
-The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a).
+Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\boldsymbol{\eta}(t)$ where $\boldsymbol{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$.
+The “effective input”, ${\bf{x}}(t)=W_{x}\boldsymbol{\eta}(t)$, is therefore low-dimensional (see green dots in panel a).
 The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
-Figure 12: Response properties of a network with independent singular left- and right-singular vectors.
+Figure 9: Response properties of a network with independent left- and right-singular vectors.
 Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors.
@@
 The fixed point at ${\bf{z}}=0$ in unstable because $c>1$.
+Figure 12: Dynamics of a stable network with rank-one structure close to criticality.
+Same as Figure 1 except $\rho=0.95$.
+The fixed point at ${\bf{z}}=0$ is stable, but close to instability because the maximum real part of the eigenvalues of the Jacobian is close to 0.
+Figure 13: Power spectral density of ${\bf{x}}(t)$ used in simulations.
+Average power spectral density of ${\bf{x}}_{j}(t)$ (green) compared to $|r(f)|^{2}=1/(1+4\pi^{2}f^{2}\tau^{2})$ (gray) when $\tau_{x}=10$ and $\tau=1$, as in all figures except for Figure 14.
+Vertical red line shows $f=1/(2\pi\tau)$.
+Figure 14: High- and low-dimensional dynamics when inputs are not slow.
+Simulations from Figure 3 were repeated with $\tau_{x}=1$ and $\tau=10$.
+a) Average power spectral density of ${\bf{x}}_{j}(t)$ (green) compared to $|r(f)|^{2}=1/(1+4\pi^{2}f^{2}\tau^{2})$ (gray).
+Vertical red line shows $f=1/(2\pi\tau)$.
+Since ${\bf{x}}(t)$ is faster than intrinsic network dynamics ($\tau_{x}<\tau$), $c_{x}(f)$ has power at higher frequencies where $r(f)\approx 0$.
+b-e) Same as Figure 3 panels c,f,i,l,o respectively except that we set $\tau_{x}=1$ and $\tau=10$.
+Variability in ${\bf{z}}(t)$ is high- and low-dimensional under the same conditions.
 Figure 15: Response properties of a network with biased weights.
@@
 Figure 18: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (7) and the spatial network from Figure 5.
-Figure 19: Singular values and dynamics in an epidemiological network. a) Singular values of $W$ and b) projections of $\bf{z}$} for the model in Figure 6.
+Figure 19: Singular values and dynamics in an epidemiological network.
+a) Singular values of $Q$ (red) and $W$ (blue).
+b) Projections of $\bf{z}$} for the model in Figure 6.
+c) Same as Figure 6a except the dominant singular vector of $Q$ was used to define the aligned direction instead of the dominant singular value of $W$.
+The lack of suppression demonstrates the necessity of using the effective connectivity matrix, $W$, in place of $Q$, which is the connectivity matrix used to define the dynamics.
+Figure 20: Response properties of a weakly low-rank network with $\rho\ll 1$ and $c<0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
+Figure 21: Response properties of a weakly low-rank network with $\rho\ll 1$ and $c>0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$. Note that stability requires that $c<1$ when $c>0$.
+Figure 22: Quantifying low-rank suppression as the maximum singular value is scaled.
+We repeated the simulation from Figure 1 while changing the value of $c<0$ to change the value of the maximum singular value of $W$, $\sigma_{\textrm{max}}=|c|$.
+This demonstrates a smooth transition between weakly and strongly low-rank structure.
+For each value of $c$, we computed the ratio of $\|{\bf{z}}\|$ in response to an aligned input versus a random input, i.e., the height of the two bumps in Figure 1g as a measure of low-rank suppression.
+Low-rank suppression occurs when this ratio is small.
+Figure 23: High-dimensional dynamics in weakly low-rank networks.
+a) Singular values for a weakly low-rank network in which $\rho=0.01$, $\sigma_{1}=0.5$, and $\sigma_{2}=0.25$.
+b-f) Same as Figure 3c,f,i,l,o except that $\rho=0.01$, , $\sigma_{1}=0.5$, and $\sigma_{2}=0.25$.
+The network response is high-dimensional in every case.
+Figure 24: Comparison between theory and simulations for shorter and longer simulations.
+a) Same as Figure 1d except we added the theoretical values form Eq.
+(16) (red dots) and the values obtained from Eq.
+(17) (green dots).
+The network was simulated for $T=5\times 10^{3}$ time units.
+b) Same as b except we increased the simulation time to $T=5\times 10^{4}$.